### v2

Abstract
This paper addresses information design with linear best responses of agents, quadratic objective functions of an information designer, and a payoff state distributed according to a Gaussian distribution. We formulate the problem as semidefinite programming (SDP) and use the duality principle to characterize an optimal information structure. There exists a Gaussian information structure that is optimal among all information structures. A necessary and sufficient condition for optimality is that the realizations of the induced action profile and a state satisfy linear constraints derived from the primal and dual SDP. As a result, an observed action profile typically reveals the true state even if individual agents have only partial knowledge. In symmetric network games, an optimal information structure inherits this symmetry, which facilitates the computation of an optimal information structure.
JEL classification: C72, D82
Keywords: linear-quadratic Gaussian, information design, Bayesian persuasion, network games, semidefinite programming, strong duality
1 Introduction
We study games among agents who strategically interact with one another under incomplete information. The optimal action of each agent—e.g., a level of investment—is dependent on other agents’ action choices through payoff externalities, as well as on exogenous payoff-relevant states, such as random returns on investment. An extra player of the game, called an information designer, chooses an information structure—i.e., a joint distribution of payoff states and agents’ private signals—to manipulate agents’ incentives. Our aim is to understand how the information designer can optimally design such information structures for her objective maximization.
To bring insights into the relationship between optimal information structures and economic environments, this paper addresses a tractable class of information design problems that are neither too specific nor too general. To this end, we focus on linear-quadratic Gaussian (LQG) games, where agents’ payoff functions and designer’s objective function are quadratic in continuous actions and payoff states, which follow a multivariate normal, or Gaussian, distribution. Since payoff functions are arbitrary as long as they are quadratic, the class can encapsulate a wide array of strategic situations. [^3] The origin of LQG games dates back to the seminal work of Radner (1962). Several instances of LQG games include Cournot and Bertrand oligopoly (Vives, 1984, 1999), financial markets and beauty contests (Morris and Shin, 2002, Angeletos and Pavan, 2007), public goods provision (Bramoullé and Kranton, 2007), and network games (Ballester et al., 2006, Bramoullé et al., 2014) among other applications. In addition, when payoffs are quadratic, an objective can typically be quadratic, as when a utilitarian planner aims to maximize the sum of agents’ payoffs. The information designer has the ability to select and commit to an information structure from a feasible set, fixing the payoff functions and the prior distribution of states. LQG information design is a problem of finding an optimal information structure that maximizes the expected value of the objective function.
No a priori restrictions are imposed on the distribution of private signals. However, our foundational observation is that among all information structures, there always exists an optimal Gaussian information structure, under which states and signals are jointly normally distributed. This arises from the nature of quadratic objective and payoff functions: In expectation, the designer’s objective is determined solely by the covariance matrix of equilibrium actions and states. In Lemma 1, we then show that the class of Gaussian information structures is versatile enough to implement any covariance matrix that could emerge under an arbitrary information structure. Consequently, it is without loss of optimality for the information designer to restrict attention to Gaussian information structures, despite many other alternatives.
A Gaussian information structure is determined solely by its first and second moments, whereas the first must be constant across all information structures by Lemma 1. Hence, building upon the optimality of Gaussian information structures, we can formulate LQG information design as the problem of choosing an optimal covariance matrix. This formulation entails maximizing a linear objective function of a matrix subject to two fundamental constraints: a system of linear constraints regulating each agent’s incentives, and the positive semidefiniteness constraint to ensure that the matrix serves as a statistically valid covariance matrix. This problem structure—maximizing a linear function of a positive semidefinite matrix subject to linear constraints—mirrors that of semidefinite programming (SDP), and for this reason, we can employ several methodological tools developed in convex optimization. [^4] SDP is known as a generalization of linear programming; see, e.g., Vandenberghe and Boyd (1996), Boyd and Vandenberghe (2004), or Luenberger and Ye (2021) for further details.
Departing from the information designer’s maximization problem, we derive the associated dual problem, formulated as the minimization of a function of Lagrangian multipliers corresponding to each agent’s incentive constraint. Proposition 2 then asserts that the dual problem provides a tight upper bound on the information designer’s optimal value; in other words, strong duality holds. This strong duality leads to the characterization of an optimal information structure based on its “certifiability” by Lagrangian multipliers. Our main technical result, Proposition 3, states that an action-state covariance matrix is optimal in LQG information design if and only if it can satisfy the system of matrix equations together with an appropriate choice of Lagrangian multipliers. By exploiting the statistical interpretation of these equations, in Corollary 6, we show that the optimal information structure exhibits a linear relationship that holds between the equilibrium action profile and states.
Our characterization is useful particularly when one’s interest lies in checking the optimality of a given information structure. By using it, for example, Corollary 7 obtains necessary and sufficient conditions for two extreme forms of information disclosure—no disclosure (i.e., nobody receives an informative signal) and full disclosure (i.e., everyone is fully informed of the true state)—in terms of payoff parameters. It turns out that the optimality of these cases critically hinges on the dimension of the payoff state. Specifically, when the state is multidimensional, each of them is suboptimal for generic payoff structures because the corresponding optimality condition imposes a greater number of linear equations than the number of Lagrangian multipliers that can be chosen as certifiers. In contrast, in the one-dimensional case, we obtain a “bang-bang” result, Corollary 10, which asserts that an optimal information structure is either no disclosure or “aggregative” full disclosure, meaning that each individual may remain partially informed, yet the true state is fully revealed to the entire population.
In general, it is not an easy task to analytically derive optimal information structures since our characterization is based on a high-dimensional system of equations. However, there are several known optimization tools that calculate numerical solutions to SDP problems. Moreover, by assuming a certain degree of symmetry across agents’ payoffs, we demonstrate that the dimension can be effectively reduced so that an optimal information structure would be available not only numerically but also analytically.
Lastly, we explore the case where the information designer is limited to public information structures—under which all agents receive identical signals—and study an optimal information structure within this class. This problem holds practical importance when individualized information disclosure is unfeasible due to technological constraints, transparency principles, etc. Focusing on public information structures gains analytical tractability, as each agent no longer encounters strategic uncertainty regarding others’ actions. Incidentally, public LQG information design reduces to a Bayesian persuasion problem with a single representative agent, which is analogous to the problem studied by Tamura (2018). Leveraging his solution method, Proposition 15 provides a closed-form expression for the optimal public information structure. It is noteworthy that the optimal public information structure is generically suboptimal within the class of all information structures, including non-public ones, unless the state is one-dimensional. This indicates that the information designer typically benefits from manipulating each agent’s higher-order belief regarding others’ private information. [^5] Arieli and Babichenko (2019) and Hoshino (2022) also highlight the importance of leveraging strategic uncertainty in games with finite actions and binary states, while their settings do not exhibit the generic gap between private and public information structures as in our result.
This paper contributes to the growing literature on information design, which has yielded a wide array of economic applications in various contexts. [^6] See survey papers by Bergemann and Morris (2019) and Kamenica (2019). A single-receiver version of information design is also called Bayesian persuasion (Kamenica and Gentzkow, 2011). Applications of information design with multiple receivers include matching markets (Ostrovsky and Schwarz, 2010), voting games (Alonso and Câmara, 2016), congestion games (Das et al., 2017), auctions (Bergemann et al., 2017a), contests (Zhang and Zhou, 2016), and stress testing (Inostroza and Pavan, 2023). While the literature has investigated several applications and general approaches, research on information design with multiple interacting receivers remains relatively sparse due to technical challenges arising from players’ intricately interlinked incentives in the presence of fundamental and strategic uncertainties. The main contribution of this paper lies in formulating the information designer’s problem in the SDP form and characterizing optimal information structures through relevant optimality conditions.
Our approach to information design is built on an outcome-based approach, which involves the following decomposition of the problem into two steps: [^7] See Bergemann and Morris (2013, 2016a, 2016b, 2019), Bergemann et al. (2015, 2017b), and Taneva (2019). Mathevet et al. (2020) develop a different belief-based approach, which is a natural analogue of the concavification method of Kamenica and Gentzkow (2011) for Bayesian persuasion problems. First, we identify the feasible set of outcomes—i.e., action-state joint distributions—that could be induced by the information designer. Second, we determine which of these outcomes would maximize the information designer’s objective. In symmetric LQG games with agents sharing identical payoff functions, Bergemann and Morris (2013) characterize the feasible outcomes in terms of linear constraints on the first and second moments of outcomes by restricting attention to a stylized class of conditionally i.i.d. symmetric information structures. [^8] Bergemann et al. (2017b) and Miyashita (2024) extend these constraints to asymmetric LQG games with finite/continuum populations, respectively. We further show that the constraints are necessarily satisfied even when signals are not Gaussian; this result serves as a building block to establish the optimality of Gaussian information structures. However, it has remained unclear (i) whether there exists an asymmetric or non-Gaussian information structure that surpasses the optimal conditional i.i.d. information structure, and (ii) how to solve the second step optimization when agents exhibit asymmetry in terms of payoff externalities and/or private information. We advance the analysis of Bergemann and Morris (2013), as well as subsequent studies, by broadening the class of environments considered and by recasting the problem in the SDP form. This comprehensive extension enables us to establish the “global” optimality of Gaussian information structures and characterize it via duality.
The duality methodology is widely used in optimization problems, including information design. [^9] For example, Kolotilin (2018), Dworczak and Martini (2019), and Kolotilin et al. (forthcoming) employ the duality methodology to study persuasion problems with a single receiver. Among the closest to our work is Smolin and Yamashita (2023), which study information design in games with strictly concave utility functions. In their framework, the information design problem is formulated as infinite-dimensional linear programming, with the objective being linear in a probability measure over actions and states, subject to linear constraints. They derive the associated dual problem and establish weak duality. In contrast, under quadratic specifications, we formulate the problem as SDP with respect to an action-state covariance matrix, involving only finitely many variables. This reduction in dimensionality enhances tractability, allowing us to establish strong duality and complementary slackness conditions. Furthermore, all economic applications provided by Smolin and Yamashita (2023) are nested within our framework, and we demonstrate that their insights regarding the optimality of Gaussian information structures extend generally to all LQG environments.
This paper also relates to the literature on network interventions, an active topic studied across many disciplines. [^10] See, for example, Valente (2012) for a general introduction to the subject. In economics, the seminal work by Ballester et al. (2006) formulate the intervention problem as a network game and quantify the structural importance of each individual by comparing equilibrium group behavior with and without the individual in the network. Galeotti et al. (2020) consider the planner’s intervention of changing targeted individuals’ base action levels. Within our model, these interventions can be precisely captured as exogenous changes in payoff structures. In contrast, our focus lies in altering information structures by considering the effect of information disclosure on equilibrium behavior. [^11] This type of informational intervention has garnered significant attention in recent empirical literature. For example, Banerjee et al. (2013) examine the effect of information transmission on microfinance participation using demographic and social network data. Thus, our study provides complementary insights for planners aiming to influence group behavior by manipulating individuals’ incentives through direct interventions and indirect strategic spillovers to their neighbors.
The rest of the paper is organized as follows. Section 2 presents a leading example to motivate our general analysis. Section 3 introduces the model of LQG information design and characterizes a Bayes Nash equilibrium through moment restrictions. Section 4 formulates LQG information design as an SDP problem, sets up the dual problem, and offers our main results. Sections 5 and 6 focus on special cases: symmetric LQG games with symmetric information structures and asymmetric LQG games with public information structures, respectively. All omitted proofs can be found in the appendix.
2 Leading example
Consider a Cournot duopoly game with a homogeneous product. Firm $i\in\{1,2\}$ produces $a_{i}$ units of the product with a marginal cost $\theta_{i}$. The inverse demand function is specified as $p(a_{1},a_{2})=1-(a_{1}+a_{2})$ so that firm $i$’s profit function is
$\quantity(1-(a_{i}+a_{j}))a_{i}-\theta_{i}a_{i}\quad{\rm where}\quad j\neq i.$
Assume that $\theta=(\theta_{1},\theta_{2})$ follows a joint normal distribution with $\operatorname{E}\quantity[\theta_{1}]=\operatorname{E}\quantity[\theta_{2}]=\bar{\theta}$ and $\operatorname{Var}\quantity[\theta_{1}]=\operatorname{Var}\quantity[\theta_{2}]=1$. Regarding the covariance, we consider two cases: In the common-value case, we let $\operatorname{Cov}\quantity[\theta_{1},\theta_{2}]=1$, wherein two firms are exposed to the same marginal cost $\theta_{1}=\theta_{2}$. In the private-value case, we let $\operatorname{Cov}\quantity[\theta_{1},\theta_{2}]=0$ so that $\theta_{1}$ and $\theta_{2}$ are independently distributed.
Each firm does not know the realization of $\theta$ when deciding on its production level. Instead, there is an omniscient information designer who provides private information $t_{i}$ about $\theta$ to firm $i$ according to a pre-announced information structure $\pi$, which is a joint probability distribution over $(t,\theta)=(t_{1},t_{2},\theta_{1},\theta_{2})$. Firm $i$’s strategy $\sigma_{i}$ is a function that assigns an output level $\sigma_{i}(t_{i})\in\mathbb{R}$ to each realization of a private signal $t_{i}$. Regardless of the information structures, it is known that there exists a unique BNE, denoted by $\sigma^{\pi}=(\sigma_{1}^{\pi},\sigma_{2}^{\pi})$, and the expected value of an action $\operatorname{E}[\sigma_{i}^{\pi}]$ does not depend upon $\pi$. In addition, the variance and covariance of actions necessarily maintain the following linear constraints:
$2\operatorname{Var}\quantity[\sigma_{i}^{\pi}]+\operatorname{Cov}\bigl{[}\sigma_{i}^{\pi},\sigma_{j}^{\pi}\bigr{]}+\operatorname{Cov}\quantity[\sigma_{i}^{\pi},\theta_{i}]=0,\quad\forall i\neq j.$ (1)
We postulate that the information designer chooses $\pi$ to maximize a convex combination of producer surplus (PS) and consumer surplus (CS) by anticipating that the firms follow a unique BNE $\sigma^{\pi}$ under the chosen information structure. PS is defined as the total profit of the firms (since there is no fixed cost), and CS is calculated as the integral of consumer’s demand net of the price (which equals the half of the squared aggregated production in the linear demand model). Hence, the designer’s ex-post objective function is
$(1-\lambda)\cdot\underbrace{\quantity[p(a_{1},a_{2})(a_{1}+a_{2})-\theta_{1}a_{1}-\theta_{2}a_{2}]}_{\rm(PS)}+\lambda\cdot\underbrace{\quantity[\frac{\quantity(a_{1}+a_{2})^{2}}{2}]}_{\rm(CS)},$
where $\lambda\in[0,1]$ is a weight on CS. Taking the expectation when $(a_{1},a_{2})=\sigma^{\pi}$, we obtain the value of $\pi$, which is shown to satisfy
$\operatorname{E}\quantity[(1-\lambda)\cdot{\rm PS}+\lambda\cdot{\rm CS}]\propto\operatorname{Var}\quantity[\sigma_{1}^{\pi}]+2c\cdot\operatorname{Cov}\quantity[\sigma^{\pi}_{1},\sigma^{\pi}_{2}]+\operatorname{Var}\quantity[\sigma_{2}^{\pi}],$ (2)
where $c\equiv\lambda/(2-\lambda)$. There is no term involving $\operatorname{Cov}[\sigma_{i}^{\pi},\theta_{i}]$ since we can remove it by using (1); see Appendix A for detailed calculations. Note that the designer’s objective is proportional to PS (resp. CS) if $c=0$ (resp. $c=1$). In addition, it is proportional to total surplus (TS)—i.e., the sum of PS and CS—if $c=1/3$ because this implies $\lambda=1/2$.
The designer’s objective (2) is dependent on the choice of $\pi$ only through the covariance matrix of the induced action-state joint distribution. Specifically, if we denote by $x_{ij}=\operatorname{Cov}[\sigma^{\pi}_{i},\sigma^{\pi}_{j}]$ and $y_{ij}=\operatorname{Cov}[\sigma^{\pi}_{i},\theta_{j}]$, then (2) can be written as
$x_{11}+x_{22}+cx_{12}+cx_{21},$ (3)
where the choice is subject to (1), i.e.,
$02x_{11}+x_{12}+y_{11}=0\quad{\rm and}\quad 2x_{22}+x_{21}+y_{22}=$ 0 . (4)
In addition, $2\times 2$ matrices $X=[x_{ij}]_{2\times 2}$ and $Y=[y_{ij}]_{2\times 2}$ necessarily satisfy that
$\matrixquantity[X&Y\\ Y^{\top}&\operatorname{Var}\quantity[\theta]]\text{ is positive semidefinite}$ (5)
since (5) is a covariance matrix. This implies that the designer’s optimal value is at most the maximum of (3) subject to (4) and (5). In fact, this upper bound is tight since the constraints are not only necessary but also sufficient for any targeted action-state covariance matrix to be induced in a BNE under some Gaussian information structure. [^12] This result obtains as a consequence of Lemma 1 in Section 3.2. In Miyashita and Ui (2024), we also provide a similar result in a continuum-population setting for the common-value case. The lemma has another important implication that we can focus on Gaussian information structures when searching for optimal information structures, including non-Gaussian ones. Although this point is implicitly suggested by Bergemann and Morris (2013), it is not explicitly mentioned in their work.
Therefore, we can reformulate the information designer’s problem as the maximization of a linear objective subject to linear and positive semidefiniteness constraints, which are the defining features of optimization problems known as semidefinite programming (SDP). There are several SDP optimization tools that are readily available, and by using one of them, we can solve the current problem numerically for both common-value and private-value cases.
Table 1: The common-value case
PS-optimal: $c=0$ TS-optimal: $c=1/3$ CS-optimal: $c=1$ Full-disclosure
[…]
$\operatorname{Var}\quantity[\theta_{1}\mid\sigma^{\pi}]$ $0.000$ $0.000$ $0.000$ $0.000$
Table 1 summarizes the numerical results for the common-value case. The first row lists the covariance matrices of actions $\operatorname{Var}[\sigma^{\pi}]$ that maximize PS, CS, and TS in that order, along with that of full disclosure (i.e., $t_{1}=t_{2}=\theta$) for the sake of comparison. The second row lists the corresponding conditional covariance matrices of actions $\operatorname{Var}[\sigma^{\pi}\mid\theta]$ given $\theta$. Full disclosure is both TS-optimal and CS-optimal, as the optimal covariance matrices are identical to that of full disclosure, under which $\sigma^{\pi}_{1}$ and $\sigma^{\pi}_{2}$ are perfectly correlated with each other and with $\theta$, resulting in identical entries of $\operatorname{Var}[\sigma^{\pi}]$. In contrast, full disclosure is not PS-optimal, as $\operatorname{Var}[\sigma^{\pi}]$ is a diagonal matrix, meaning that $\sigma^{\pi}_{1}$ and $\sigma^{\pi}_{2}$ are independently distributed. While $\theta$ does not determine each output, it precisely determines the total output, as one can see that
$\operatorname{Var}\quantity[\sigma_{1}^{\pi}+\sigma_{2}^{\pi}\mid\theta]=\matrixquantity[1\\ 1]^{\top}\operatorname{Var}\quantity[\sigma^{\pi}\mid\theta]\matrixquantity[1\\ 1]=\matrixquantity[1\\ 1]^{\top}\matrixquantity[0.063&-0.063\\ -0.063&0.063]\matrixquantity[1\\ 1]=0.000.$
This suggests that the true value of $\theta$ can be precisely inferred from the observed action profile, while each firm is kept partially informed. In fact, $\theta$ has zero variance conditional on the equilibrium action profile, as can be seen from the third row of Table 1.
Table 2: The private-value case
PS-optimal: $c=0$ TS-optimal: $c=1/3$ CS-optimal: $c=1$ Full-disclosure
[…]
Table 2 summarizes the numerical results for the private-value case. While full disclosure is neither PS-, TS-, nor CS-optimal, $\theta$ uniquely determines each firm’s output under any optimal information structure, as the conditional covariance matrices are zero in the second row. This also suggests that the observed action profile fully reveals $\theta$, and this is confirmed by the zero covariance matrix conditional on the equilibrium action profile in the third row. For example, $\sigma^{\pi}_{1}$ and $\sigma^{\pi}_{2}$ are independently distributed under the CS-optimal information structure since the covariance matrix is diagonal. As can be verified in this case, firm $i$ knows its own marginal cost $\theta_{i}$ precisely but lacks information on the marginal cost of the opponent $\theta_{j}$.
These numerical results present a sharp contrast with respect to the optimality of full disclosure. Nevertheless, in both cases, the true value of $\theta$ can be precisely inferred from the observed action profile $(a_{1},a_{2})$ under optimal information structures; that is, the information designer eventually reveals all the information about $\theta$ to the population. In addition, all optimal covariance matrices turn out to be invariant up to permuting firms’ indeces, which may be reflecting the symmetry in firms’ profit functions. As these observations are specific to the above numerical solutions, we proceed with an analytical study of a more general class of information design problems.
3 Model
[…]
3.1 Basic games and information structures
There are $n$ agents indexed by $i=1,\ldots,n$, each of whom chooses an action $a_{i}\in\mathbb{R}$ as a real number. Agent $i$’s payoff is quadratically dependent on an action profile $a\equiv(a_{1},\ldots,a_{n})\in\mathbb{R}^{n}$ and a payoff state $\theta\equiv(\theta_{1},\cdots,\theta_{m})\in\mathbb{R}^{m}$, where $m\geq 1$ is an integer. Specifically, let the payoff function be given as
$u_{i}(a,\theta)\equiv-\underbrace{\frac{q_{ii}}{2}\cdot a_{i}^{2}}_{\begin{subarray}{c}\text{private cost}\\ \text{of own action}\end{subarray}}+\ \underbrace{a_{i}\cdot\quantity(\sum_{k\in[m]}r_{ik}\theta_{k}-\sum_{j\in[n]\setminus\{i\}}q_{ij}a_{j})}_{\text{returns from own action}}+\underbrace{h_{i}(a_{-i},\theta)}_{\text{pure externalities}},$ (6)
where $q_{ij}$ and $r_{ik}$ are constant, and $h_{i}(a_{-i},\theta)$ is an arbitrary function of the opponents’ actions $a_{-i}\equiv(a_{j})_{j\neq i}$ and a payoff state $\theta$. Let $Q\equiv[q_{ij}]_{n\times n}$ and $R\equiv[r_{ik}]_{n\times m}$ be the matrices consisting of the coefficients of the quadratic terms in (6).
Throughout the paper, we employ the following regularity conditions.
Assumption 1.
The matrix $Q+Q^{\top}$ is positive definite. A payoff state $\theta$ follows a multivariate normal distribution with a mean vector $\bar{\theta}$ and a covariance matrix $Z\equiv\operatorname{Var}[\theta]\neq O$.
The positive definiteness of $Q$ guarantees the unique existence of an equilibrium under an arbitrary information structure. That $Z\neq O$ is a minimal assumption to make our information design problem non-trivial. In each agent’s payoff function (6), $h_{i}$ has no influence on agent’s decisions so that the strategic environment can be summarized by the matrices $Q$ and $R$ alone. Furthermore, since $\theta$ is normally distributed, its distribution can be summarized by the mean vector $\bar{\theta}$ and the covariance matrix $Z$, while $\bar{\theta}$ will be of no importance in our problem. In light of these, we refer to the profile $(Q,R,Z)$ as a basic game.
Agent $i$ receives a private signal which is a random variable $t_{i}$. Let $\pi(t,\theta)$ be the joint distribution of a signal profile $t=(t_{1},\ldots,t_{n})$ and a state $\theta$, with the marginal distribution over $\theta$ being normal with mean $\bar{\theta}$ and variance $Z$. We refer to $\pi$ as an information structure. If $t$ and $\theta$ are jointly normally distributed, we refer to $\pi$ as a Gaussian information structure. Let $\Pi^{*}$ denote the set of all information structures, and let $\Pi^{\rm g}\subseteq\Pi^{*}$ be its subset consisting of all Gaussian ones.
Together with a a basic game $(Q,R,Z)$, an information structure $\pi$ defines a Bayesian game. In particular, we say that the game is linear-quadratic-Gaussian (LQG) when $\pi$ is given as a Gaussian information structure. To illustrate the flexibility of the present model, some examples are listed below.
Example 1.
An LQG game with identical payoff functions is known as an LQG team, originated from the seminal study of Radner (1962). Specifically, his setup assumes an identical payoff function across agents,
$u_{i}(a,\theta)=-a^{\top}Qa+2a^{\top}R\theta,$ (7)
where $Q\in\mathcal{S}^{n}$ and $R\in\mathcal{M}^{n,m}$. Any LQG game with symmetric $Q$ has the same system of best responses as that of the LQG team with the same coefficient matrices $Q$ and $R$, and for this reason, Radner’s characterization of an equilibrium is useful even without the assumption of identical payoff functions. [^13] In other words, an LQG game with a symmetric coefficient matrix is a potential game (Monderer and Shapley, 1996) with a potential function (7); see Ui (2009) for more details. The next two examples are special cases.
Example 2.
The Cournot duopoly in Section 2 can be extended to oligopoly with product differentiation (Vives, 1999). Firm $i$ produces $a_{i}$ units of good $i$ with a marginal cost $\theta_{i}$. Assuming a linear inverse demand function, firm $i$’s profit is
$\underbrace{\quantity(\alpha-\beta_{i}a_{i}-\sum_{j\neq i}\gamma_{ij}a_{j})}_{\text{price}}a_{i}-\underbrace{\theta_{i}a_{i}}_{\text{cost}}.$
A Bertrand game is another example, where $a_{i}$ is the price of good $i$ (Vives, 1999). Firm $i$’s profit is
$\underbrace{\quantity(\alpha-\beta_{i}a_{i}+\sum_{j\neq i}\gamma_{ij}a_{j})}_{\text{demand}}\cdot\underbrace{(a_{i}-\theta_{i})}_{\text{net price}}.$
Example 3.
Agents are embedded in a network, consisting of a set of vertices $[n]$ and a set of edges $G\subseteq[n]^{2}$. We assume there are no self-loops, so $(i,i)\notin G$ for every $i\in[n]$. The payoff to agent $i$ is given by
$-\frac{a_{i}^{2}}{2}+a_{i}\cdot\underbrace{\quantity(q\sum_{j:(i,j)\in G}a_{j}+\sum_{k=1}^{m}r_{ik}\theta_{k})}_{\text{marginal return from own action}}.$
In this payoff function, the marginal return from an agent’s own action depends on the actions of neighbors—those agents $j$ such that $(i,j)\in G$—weighted by the degree $q\in\mathbb{R}$ of payoff externality. We can identify the network structure $G$ as the adjacency matrix $G=[g_{ij}]_{n\times n}$, where $g_{ij}=1$ if $(i,j)\in G$ and $g_{ij}=0$ otherwise. Consequently, the payoff coefficient matrix is given by $Q=I+qG$. This setting corresponds to network games studied by Ballester et al. (2006) and Bramoullé et al. (2014) when $\theta$ is non-random. Calvó-Armengol et al. (2015) consider normally distributed $\theta$ to analyze endogenous communication in organizations. Galeotti et al. (2020) consider weighted graphs like our general model to study network intervention problems.
3.2 Bayes Nash equilibrium
Given any information structure $\pi$, agent $i$’s strategy is given as a measurable function $\sigma_{i}$ of $t_{i}$ with $\operatorname{E}|\sigma_{i}(t_{i})|^{2}<\infty$. A strategy profile $\sigma=(\sigma_{1},\ldots,\sigma_{n})$ is a Bayes Nash equilibrium (BNE) under $\pi$ if each agent’s strategy maximizes her conditional expected payoff given the opponents’ strategies; that is,
$\sigma_{i}(t_{i})\in\mathop{\rm arg~{}max}\limits_{a_{i}\in\mathbb{R}}\operatorname{E}_{\pi}\quantity[u_{i}((a_{i},\sigma_{-i}(t_{-i})),\theta)\mid t_{i}],\quad\forall i\in[n]$
hods almost surely (a.s.), where $\sigma_{-i}(t_{-i})=(\sigma_{j}(t_{j}))_{j\neq i}$ and $\operatorname{E}_{\pi}[\cdot\mid t_{i}]$ is the conditional expectation operator given $t_{i}$ with respect to $\pi$. We simply write $\operatorname{E}\quantity[\cdot]$ instead of $\operatorname{E}_{\pi}\quantity[\cdot]$ as long as there is no risk of confusion. It is known that under Assumption 1, a unique BNE exists for any information structure. [^14] See Radner (1962, Theorem 3) for the case of symmetric $Q$ and Ui (2016, Proposition 4) for the case of asymmetric $Q$. By means of this fact, we denote by $\sigma^{\pi}$ the unique BNE under each $\pi$.
Fact 1.
If $Q+Q^{\top}$ is positive definite, then a unique BNE $\sigma^{\pi}$ exists for each $\pi\in\Pi^{*}$. In particular, if $\pi\in\Pi^{\rm g}$, then $(\sigma^{\pi},\theta)$ is jointly normally distributed.
Since $u_{i}$ is quadratic and concave, as specified in (6), we can see that a strategy profile $\sigma$ is a BNE if and only if
$\sum_{j=1}^{n}q_{ij}\operatorname{E}\quantity[\sigma_{j}\mid t_{i}]=\sum_{k=1}^{m}r_{ik}\operatorname{E}\quantity[\theta_{k}\mid t_{i}],\quad\forall i\in[n],$ (8)
where $\sigma_{j}=\sigma_{j}(t_{j})$. This condition displays linear constraints that the conditional expectations of the state and equilibrium actions have to obey. Moreover, applying the law of iterated expectations, we can show that the unconditional counterpart of (8) holds for expectations and covariances. Specifically, the unique equilibrium $\sigma^{\pi}$ under any (possibly, non-Gaussian) information structure necessary satisfies the following moment restrictions:
$\sum_{j=1}^{n}q_{ij}\operatorname{E}[\sigma_{j}^{\pi}]=\sum_{k=1}^{m}r_{ik}\operatorname{E}[\theta_{k}],\quad\forall i\in[n],$ (9)
$\sum_{j=1}^{n}q_{ij}\operatorname{Cov}[\sigma_{i}^{\pi},\sigma_{j}^{\pi}]=\sum_{k=1}^{m}r_{ik}\operatorname{Cov}[\sigma_{i}^{\pi},\theta_{k}],\quad\forall i\in[n].$ (10)
In matrix form, these restrictions are rewritten as
$Q\operatorname{E}\quantity[\sigma^{\pi}]=R\bar{\theta}\quad\text{and}\quad\operatorname{diag}(Q\operatorname{Var}\quantity[\sigma])=\operatorname{diag}\quantity(R\operatorname{Cov}\quantity[\theta,\sigma]).$
Since $Q$ is invertible under Assumption 1, the first equation uniquely pins down the expectation of an action profile as $\bar{a}\equiv\operatorname{E}\quantity[\sigma^{\pi}]=Q^{-1}R\bar{\theta}$ across all information structures. From the information designer’s perspective, this indicates that there is no room for changing the expected action levels by manipulating information structures.
The second equation places some restrictions on the variance matrix of an equilibrium action-state joint distribution, but in a less stringent manner than the first equation does on the expectation. Specifically, let $\mathbf{M}^{\pi}$ denote the variance matrix that arises in the unique equilibrium under an information structure $\pi$; that is,
Since $\mathbf{M}^{\pi}$ is a symmetric matrix whose lower-right block is fixed by the distribution of $\theta$, it contains ${n(n+m+1)}/{2}$ free variables, which satisfy $n$ linear equations in (10).
In the next lemma, we establish that these restrictions are not only necessary but also sufficient for a given $(n+m)\times(n+m)$ positive semidefinite matrix to be inducible as the covariance matrix of an equilibrium action-state joint distribution under some Gaussian information structure. This extends the result of Bergemann and Morris (2013) to the present setting with heterogeneous payoff functions and a multi-dimensional state with arbitrary information structures. [^15] Bergemann and Morris (2013) focus on symmetric Gaussian information structures and symmetric basic games with $m=1$ and show that the constraint (11) is necessary and sufficient for the equilibrium covariance matrices under Gaussian information structures. Proposition 5 of Miyashita and Ui (2024) provides a similar sufficiency result in an asymmetric setting with $m=1$ and a continuum of agents.
Lemma 1.
For any information structure $\pi\in\Pi^{*}$, if we write $X=\operatorname{Var}[\sigma^{\pi}]$ and $Y=\operatorname{Cov}[\sigma^{\pi},\theta]$, then it holds that
$\operatorname{diag}(XQ^{\top})=\operatorname{diag}(YR^{\top})\quad\text{and}\quad\mathbf{M}_{X,Y}\equiv\matrixquantity[X&Y\\ Y^{\top}&Z]\in\mathcal{S}_{+}^{n+m}.$ (11)
Conversely, if $X\in\mathcal{S}^{n}$ and $Y\in\mathcal{M}^{n\times m}$ satisfy (11), then there exists a Gaussian information structure $\pi\in\Pi^{\rm g}$ such that $\mathbf{M}^{\pi}=\mathbf{M}_{X,Y}$.
Definition 1.
A pair of matrices $(X,Y)\in\mathcal{S}^{n}\times\mathcal{M}^{n,m}$ is said to be (primal) feasible under $(Q,R)$ if it satisfies (11).
We simply say that $(X,Y)$ is feasible when $(Q,R)$ is clear from the context.
Lemma 1 implies that the set of the covariance matrices under BNE over all information structures coincides with that over all Gaussian ones, since
$\quantity{\mathbf{M}^{\pi}:\pi\in\Pi^{*}}=\quantity{\mathbf{M}_{X,Y}:\text{$(X,Y)$ is primal feasible}}=\quantity{\mathbf{M}^{\pi}:\pi\in\Pi^{\rm g}}.$ (12)
This observation will guarantee the optimality of Gaussian information structures in our information design problem, where the designer’s objective is given as a quadratic function in the state and actions.
3.3 LQG information design
Given a basic game $(Q,R,Z)$, we now introduce an (information) designer, who chooses an information structure $\pi$ from a set of feasible information structures $\Pi\subseteq\Pi^{*}$. The designer can make a commitment to provide information according to the following timeline, which is a standard assumption in the literature: [^16] See Mathevet et al. (2020) for detailed discussions on information design with multiple equilibria. First, the designer chooses $\pi\in\Pi$ and informs all agents of $\pi$. Second, the realization of $(t,\theta)$ is drawn according to $\pi$, and each agent $i$ observes $t_{i}$. Third, agents follow the unique BNE $\sigma^{\pi}$ under the chosen information structure $\pi$.
The designer aims at maximizing the expected value of a given objective function $v(a,\theta)$ by anticipating that the unique BNE is played under the chosen information structure. We assume that $v$ is quadratically dependent on the state $\theta$ and an action profile $a$ as follows:
and $\bar{v}$ is a linear function of $(a,\theta)$. Note that $\mathbf{V}$ is a symmetric matrix consisting of the coefficients of quadratic terms in $v(a,\theta)$.
LQG information design is the problem of finding an information structure that maximizes the expected value of the quadratic objective function over the set of feasible information structures $\Pi$:
$v^{\rm p}(\Pi)=\max_{\pi\in\Pi}\operatorname{E}\quantity[v(\sigma^{\pi},\theta)].$ (13)
An information structure $\pi\in\Pi$ is said to be optimal in $\Pi$ if it is a solution to this problem. When a feasible set is unrestricted, i.e., $\Pi=\Pi^{*}$, we simply say that such $\pi$ is optimal and write $v^{\rm p}$ instead of $v^{\rm p}(\Pi^{*})$.
The problem (13) can be reformulated in a more tractable expression, since the objective function in (13) is reduced to a linear function of the variances and covariances of actions and states. Specifically, for any $\pi\in\Pi$, it follows from the property of trace that
[…]
Therefore, by the linearity of $\hat{v}$, the objective function in (13) can be rewritten as $\mathbf{V}\bullet\mathbf{M}^{\pi}+\bar{v}(\bar{a},\bar{\theta})$, wherein the second term is not controllable since $\bar{\theta}$ is exogenously fixed and $\bar{a}$ is constant across all information structures. Therefore, it is without loss of generality to normalize $v$ by making the following assumption.
Assumption 2.
An objective function $v$ is given by (3.3) with $\bar{v}(\bar{a},\bar{\theta})=0$.
Under this assumption, the objective function in (13) can be expressed as
$\operatorname{E}\quantity[v(\sigma^{\pi},\theta)]$ $=\mathbf{V}\bullet\mathbf{M}^{\pi}$
$=V\bullet\operatorname{Var}\quantity[\sigma^{\pi}]+W\bullet\operatorname{Cov}\quantity[\sigma^{\pi},\theta]$
$=\sum_{i=1}^{n}\sum_{j=1}^{n}v_{ij}\operatorname{Cov}\quantity[\sigma^{\pi}_{i},\sigma^{\pi}_{j}]+\sum_{i=1}^{n}\sum_{k=1}^{m}w_{ik}\operatorname{Cov}\quantity[\sigma^{\pi}_{i},\theta_{k}],$ (14)
where $v_{ij}$ is the $(i,j)$-entry of $V$, which measures the marginal contribution of the covariance of the actions of agents $i$ and $j$ to the objective function. Similarly, $w_{ik}$ is the $(i,k)$-entry of $W$, which measures the marginal contribution of the covariance of agent $i$’s action and the $k$-th state.
We call $\mathbf{V}\bullet\mathbf{M}^{\pi}$ the (designer’s) value of $\pi$. For example, consider no disclosure, denoted by $\underline{\pi}$, where each agent is informed of nothing. Under $\underline{\pi}$, agents follow the constant action profile $\bar{a}$ in equilibrium, and thus the value of $\underline{\pi}$ equals $\mathbf{V}\bullet\mathbf{M}^{\underline{\pi}}=0$. Alternatively, consider full disclosure, denoted by $\overline{\pi}$, where each agent is fully informed of the state realization (and nothing else). Since $\theta$ is common knowledge under $\overline{\pi}$, agents follow the action profile $Q^{-1}R\theta$ in each state $\theta$, resulting in $\operatorname{Var}[\sigma^{\overline{\pi}}]=Q^{-1}RZ(Q^{-1}R)^{\top}$ and $\operatorname{Cov}[\sigma^{\overline{\pi}},\theta]=Q^{-1}RZ$. Consequently, the value of $\overline{\pi}$ equals $\mathbf{V}\bullet\mathbf{M}^{\overline{\pi}}=\overline{C}\bullet Z$, where
$\overline{C}\equiv(Q^{-1}R)^{\top}VQ^{-1}R+\frac{(Q^{-1}R)^{\top}W+W^{\top}Q^{-1}R}{2}\in\mathcal{S}^{m}.$ (15)
These information structures are two extremes in terms of the informativeness of private signals about the state. We refer to all remaining cases as partial disclosure. Observe that the value of $\underline{\pi}$ and that of $\overline{\pi}$ are expressed as a Frobenius inner product $C\bullet Z$, where $C=O$ for $\underline{\pi}$ and $C=\overline{C}$ for $\overline{\pi}$. As we will see in Corollary 5, this is a general property of optimal information structures.
The optimal choice of $\pi$ depends on the agents’ payoff parameters, $Q$ and $R$, and the information designer’s payoff parameters, $V$ and $W$. We collect all these coefficients of quadratic terms in a profile $\mathscr{P}\equiv(Q,R,V,W)$ and refer to it as the payoff structure of the LQG information design.
4 Main results
In this section, we seek for an optimal information structure by assuming that all information structures are feasible, i.e., $\Pi=\Pi^{*}$. In doing so, Lemma 1 brings us crucial implications. As indicated by (12), the set of equilibrium covariance matrices over all information structures coincides with that over all Gaussian ones, which is a closed convex set. Thus, the information designer can focus exclusively on Gaussian information structures without losing optimality since the objective function of LQG information design is given as a function of an equilibrium covariance matrix.
Proposition 1.
There exists a Gaussian information structure that attains $v^{\rm p}$.
This insight is novel in the context of LQG information design, although some have predicted similar results to hold in more specific contexts, e.g. Bergemann and Morris (2013), Bergemann et al. (2015), Tamura (2018), and Smolin and Yamashita (2023).
4.1 Characterization via duality
Lemma 1 further establishes an equivalent representation of LQG information design that allows applications of mathematical tools from convex optimization. Recall that a matrix $\mathbf{M}$ is an equilibrium covariance matrix under some Gaussian information structure $\pi\in\Pi^{\rm g}$ if and only if $\mathbf{M}=\mathbf{M}_{X,Y}$ for some feasible pair of matrices $(X,Y)$. By means of this observation, we can equivalently state the problem (13) as follows:
$v^{\rm p}=\max_{X\in\mathcal{S}^{n},\,Y\in\mathcal{M}^{n,m}}\mathbf{V}\bullet\mathbf{M}_{X,Y}\quad\text{subject to}\quad\operatorname{diag}\quantity(XQ^{\top})=\operatorname{diag}\quantity(YR^{\top}),\quad\mathbf{M}_{X,Y}\succeq O,$ (P)
where the objective function is linear in a feasible pair of matrices $(X,Y)$.
This formulation exhibits commonality with convex optimization known as semidefinite programming (SDP). SDP is an extension of classical linear programming, characterized as the maximization of a linear function of a positive semidefinite matrix subject to linear constraints. The associated dual problem of (P) is given by the following, where the optimization variables are a diagonal matrix $\Lambda=\operatorname{Diag}\quantity(\lambda_{1},\ldots,\lambda_{n})$ and a symmetric matrix $\Gamma=\quantity[\gamma_{kk^{\prime}}]_{m\times m}$:
Definition 2.
A pair of matrices $(\Lambda,\Gamma)\in\mathcal{S}^{n}_{\rm diag}\times\mathcal{S}^{m}$ is said to be (dual) feasible if it satisfies the constraints of (4.1).
The interpretation is that each $\gamma_{kk^{\prime}}$ is the shadow price of $\operatorname{Cov}\quantity[\theta_{k},\theta_{k^{\prime}}]$, and $\lambda_{i}$ is the Lagrangian multiplier capturing the value of relaxing agent $i$’s incentive constraint. The problem then takes the form of minimizing the total expenditure, $Z\bullet\Gamma$ subject to the dual feasibility constraint. As such, the dual feasibility constraint essentially says that the total expenditure is no less than the designer’s value, $\mathbf{V}\bullet\mathbf{M}_{X,Y}$, for all primal feasible $(X,Y)$. Furthermore, the next proposition establishes strong duality, i.e., there is no gap between $v^{\rm p}$ and $v^{\rm d}$.
Proposition 2.
Both primal problem (P) and dual problem (4.1) admit solutions. For any primal feasible $(X,Y)$ and any dual feasible $(\Lambda,\Gamma)$, it holds that $Z\bullet\Gamma\geq\mathbf{V}\bullet\mathbf{M}_{X,Y}$. Moreover, strong duality holds, i.e., $v^{\rm p}=v^{\rm d}$.
We can simplify the dual problem (4.1) using the following augmented incentive matrices:
$A_{\Lambda}\equiv\frac{\Lambda Q+Q^{\top}\Lambda}{2}-V\quad{\rm and}\quad B_{\Lambda}\equiv\frac{\Lambda R+W}{2},$ (16)
by which (4.1) is written as
Additionally, let $C_{\Lambda}$ be defined by
$C_{\Lambda}\equiv B_{\Lambda}A_{\Lambda}^{+}B_{\Lambda},$
where $A_{\Lambda}^{+}$ denotes the Moore–Penrose pseudoinverse of $A_{\Lambda}$. [^17] The Moore–Penrose inverse of $M\in\mathcal{M}^{m,n}$ is the unique matrix $M^{+}\in\mathcal{M}^{n,m}$ satisfying the following conditions: $MM^{+}M=M$, $(MM^{+})^{\top}=MM^{+}$, $M^{+}MM^{+}=M^{+}$, and $(M^{+}M)^{\top}=M^{+}M$. See, e.g., Chapter 8 of Bernstein (2018) for more details. Then, the block structure in (4.1) implies $(\Lambda,\Gamma)$ is dual feasible if and only if (i) $A_{\Lambda}$is positive semidefinite, (ii) the generalized Schur complement $\Gamma-C_{\Lambda}$ is positive semidefinite, and (iii) $B_{\Lambda}=A_{\Lambda}A_{\Lambda}^{+}B_{\Lambda}$holds true. [^18] See Proposition 10.25 of Bernstein (2018). The last condition, $B_{\Lambda}=A_{\Lambda}A_{\Lambda}^{+}B_{\Lambda}$, essentially says that each column vector of $B_{\Lambda}$ belongs to the range of $A_{\Lambda}$, which is trivially satisfied when $A_{\Lambda}$ has full rank. Therefore, since $Z\succeq O$, the optimal value of (4.1) shall take the the form of $Z\bullet C_{\Lambda}$ that can be achieved when $\Gamma$ equals the “minimum” matrix $C_{\Lambda}$. In light of this observation, we can reduce (4.1) as
We simply say that $\Lambda\in\mathcal{S}^{n}_{\rm diag}$ is (dual) feasible when $(\Lambda,C_{\Lambda})$ is dual feasible (i.e., $\mathbf{M}_{\Lambda}\succeq O$) and it is (dual) optimal if it is feasible and $v^{\rm d}=Z\bullet C_{\Lambda}$.
Now, we are ready to present the main result of this paper, which identifies optimality conditions in terms of Lagrange multipliers $\Lambda$. The conditions enable us to determine the optimality of a given information structure through the notion of “certifiability” by $\Lambda$.
Proposition 3.
A Gaussian information structure $\pi\in\Pi^{\rm g}$ is optimal if and only if there exists $\Lambda\in\mathcal{S}^{n}_{\rm diag}$ that is dual feasible and satisfies
$A_{\Lambda}\operatorname{Var}\quantity[\sigma^{\pi}]=B_{\Lambda}\operatorname{Cov}\quantity[\theta,\sigma^{\pi}]\quad\text{and}\quad A_{\Lambda}\operatorname{Cov}\quantity[\sigma^{\pi},\theta]=B_{\Lambda}Z.$ (17)
This condition is equivalent to either of the following:
$A_{\Lambda}\operatorname{Var}[\sigma^{\pi}]=B_{\Lambda}\operatorname{Cov}[\sigma^{\pi},\theta]\quad\text{and}\quad B_{\Lambda}\operatorname{Var}[\theta\mid\sigma^{\pi}]=O.$ (18)
$A_{\Lambda}\operatorname{Var}[\sigma^{\pi}\mid\theta]=O\quad\text{and}\quad A_{\Lambda}\operatorname{Cov}[\sigma^{\pi},\theta]=B_{\Lambda}Z.$ (19)
Definition 3.
We say that $\Lambda\in\mathcal{S}^{n}_{\rm diag}$ with $\mathbf{M}_{\Lambda}\succeq O$ certifies $\pi\in\Pi^{\rm g}$ if $(\pi,\Lambda)$ satisfies (17).
Proposition 3 characterizes an optimal information structure based on its certifiability by Lagrangian multipliers, defined through two matrix equations in (17). These equations can be understood as generalizing the familiar complementary slackness conditions for linear programming. Moreover, in the presence of either equation in (17), the other can be equivalently stated in the form that involves the “residual variance matrix” of an equilibrium action profile or the payoff state, which result in the two equivalent sets of optimality conditions, (18) and (19). As will be shortly discussed, these alternatives elucidate the statistical configuration of optimal information structures.
The optimality condition (17) is a restatement of strong duality. Thus, any dual feasible $\Lambda$ that certifies some optimal $\pi$ is a solution to the dual problem, and vice versa. This further implies that certifiability satisfies universality; that is, if $\Lambda$ certifies some optimal information structure, then it certifies every optimal information structure. This property will be useful when discussing the uniqueness property of optimal information structures, e.g., Corollary 8.
Corollary 4.
The following are equivalent.
  1. i).
    $\Lambda$ is dual optimal.
  2. ii).
    $\Lambda$ certifies some optimal $\pi\in\Pi^{\rm g}$.
  3. iii).
    $\Lambda$ certifies every optimal $\pi\in\Pi^{\rm g}$.
The strong duality immediately implies that the designer’s optimal value $v^{\rm p}$ is represented as a linear function of the state variance matrix $Z$, which is the common feature of the values of no disclosure and full disclosure discussed in Section 3.3. Moreover, Proposition 3 yields the following equivalent representations of $v^{\rm p}$.
Corollary 5.
Let $\pi\in\Pi^{\rm g}$ be an optimal information structure, certified by $\Lambda\in\mathcal{S}^{n}_{\rm diag}$. Then, it holds that
$v^{\rm p}$ $=C_{\Lambda}\bullet Z$
$=A_{\Lambda}\bullet\operatorname{Var}\quantity[\sigma^{\pi}]=B_{\Lambda}\bullet\operatorname{Cov}\quantity[\sigma^{\pi},\theta]$
$=A_{\Lambda}\bullet\operatorname{Var}\quantity[\operatorname{E}\quantity[\sigma^{\pi}\mid\theta]]=C_{\Lambda}\bullet\operatorname{Var}\quantity[\operatorname{E}\quantity[\theta\mid\sigma^{\pi}]].$
The value $v^{\rm p}$ is expressed as the linear function of the action variance matrix $\operatorname{Var}\quantity[\sigma^{\pi}]$ with a coefficient $A_{\Lambda}$, or that of the action-state covariance matrix $\operatorname{Cov}\quantity[\sigma^{\pi},\theta]$ with a coefficient $B_{\Lambda}$. Further, there are two more equivalent ways of representing $v^{\rm p}$. To interpret these representations, our starting point is decomposing $\operatorname{Var}\quantity[\operatorname{E}\quantity[\sigma^{\pi}\mid\theta]]$ by using the law of total variance,
$\operatorname{Var}\quantity[\operatorname{E}\quantity[\sigma^{\pi}\mid\theta]]=\underbrace{\operatorname{Var}\quantity[\sigma^{\pi}]}_{\begin{subarray}{c}\text{ex-ante}\\ \text{action variance}\end{subarray}}-\ \underbrace{\operatorname{Var}\quantity[\sigma^{\pi}\mid\theta]}_{\begin{subarray}{c}\text{conditional}\\ \text{action variance}\end{subarray}}.$
So, the matrix $\operatorname{Var}\quantity[\operatorname{E}\quantity[\sigma^{\pi}\mid\theta]]$ is statistically interpreted as quantifying the action variance reduction, i.e., the amount of the variance in equilibrium strategies $\sigma^{\pi}$ that can be reduced by conditioning on the true state $\theta$. Likewise, we can interpret $\operatorname{Var}\quantity[\operatorname{E}\quantity[\theta\mid\sigma^{\pi}]]$ as capturing the state variance reduction. Corollary 5 shows that $v^{\rm p}$ can be equally expressed as the linear function of these matrices with coefficients $A_{\Lambda}$ and $C_{\Lambda}$, respectively. Since the action variance reduction reflects the random variation of actions attributed solely to the random variation of the state $\theta$, the representation indicates the absence of extraneous randomness in the optimal value beyond $\theta$. At the same time, the other representation indicates that $v^{\rm p}$ is determined by the state variance reduction conditioning on the entire profile of equilibrium actions, which quantifies the total amount of the information on the state disclosed to the population of agents.
Next, we exploit the optimality conditions in Proposition 3 to delve into the statistical configuration of an optimal information structure $\pi$ along with certifying Lagrangian multipliers $\Lambda$. For mathematical simplicity, we assume that $Z$ has full rank in stating subsequent results. This assumption can be attained in an innocuous fashion if we are allowed to remove any redundant entry of $\theta$ to obtain a minimal representation of the model. [^19] Without the invertibility of $Z$, however, many results can be extended by considering the pseudoinverse $Z^{+}$. General results are presented and proved in the appendix whenever such extensions are feasible. The next lemma shows that when $Z$ has full rank, the dual feasibility of optimal $\Lambda$ simplifies to a lower-dimensional condition that $A_{\Lambda}\succeq O$.
Lemma 2.
Suppose $Z\succ O$. If $\Lambda\in\mathcal{S}^{n}_{\rm diag}$ satisfies (17) with some primal feasible $(X,Y)$, then $\mathbf{M}_{\Lambda}\succeq O$ if and only if $A_{\Lambda}\succeq O$.
Recall that by Fact 1, when an information structure $\pi$ is Gaussian, the associated BNE strategy profile $\sigma^{\pi}$ is jointly normally distributed with $\theta$. The next corollary shows that when $\pi$ is optimal, this joint distribution has to degenerate to some extent. Specifically, it identifies a linear relationship between $\sigma^{\pi}$ and $\theta$. Incidentally, the matrices $A_{\Lambda}$ and $B_{\Lambda}$ in the dual problem serve as the relevant coefficients of the linear relationship, which in turn guarantees the optimality of $\pi$ [^20] Note that (20) implies an alternative expression, which is based on $B_{\Lambda}$ and $C_{\Lambda}$, $B_{\Lambda}^{\top}\quantity(\sigma^{\pi}-\bar{a})=C_{\Lambda}\quantity(\theta-\bar{\theta})\quad{\rm a.s.}$ since $B_{\Lambda}=A_{\Lambda}A_{\Lambda}^{+}B_{\Lambda}$ and $C_{\Lambda}=B_{\Lambda}^{\top}A_{\Lambda}^{+}B_{\Lambda}$ hold for any dual feasible $\Lambda$.
Corollary 6.
Suppose $Z\succ O$. If $\pi\in\Pi^{\rm g}$ is optimal and certified by $\Lambda\in\mathcal{S}^{n}_{\rm diag}$, then
$A_{\Lambda}\quantity(\sigma^{\pi}-\bar{a})=B_{\Lambda}\quantity(\theta-\bar{\theta})\quad{\rm a.s.}$ (20)
Conversely, if $\pi\in\Pi^{\rm g}$ and $\Lambda\in\mathcal{S}^{n}_{\rm diag}$ with $A_{\Lambda}\succeq O$ satisfy (20), then $\pi$ is optimal in $\Pi^{*}$ and certified by $\Lambda$.
This linear relationship implies that any weighted aggregate $\sum_{i=1}^{n}\alpha_{i}\sigma^{\pi}_{i}$ of equilibrium actions is precisely determined as a function of the state $\theta$ when the weights $(\alpha_{1},\ldots,\alpha_{n})$ are taken from the range of $A_{\Lambda}$. To see this, observe that (20) implies
$A_{\Lambda}\operatorname{E}\quantity[\sigma^{\pi}-\bar{a}\mid\theta]=B_{\Lambda}\quantity(\theta-\bar{\theta})\quad{\rm a.s.}$
Thus, if $\alpha=A_{\Lambda}^{\top}\mathbf{x}$ for some $\mathbf{x}\in\mathbb{R}^{n}$, then $\alpha^{\top}\sigma^{\pi}$ can be written as a function of $\theta$. By similar arguments, one can also see that the linear statistics of the state, $\sum_{k=1}^{m}\beta_{k}\sigma^{\pi}_{k}$, is pinned down by the equilibrium strategy profile whenever the vector $\beta=(\beta_{1},\ldots,\beta_{m})$ belongs to the range of $B_{\Lambda}^{\top}$. In particular, Corollary 6 yields important implications for the full identification of $\theta$ from $\sigma^{\pi}$, or vice versa, when $A_{\Lambda}$ or $B_{\Lambda}$ has full rank. Specifically, $\sigma^{\pi}$ uniquely determines $\theta$ through (20) if and only if $B_{\Lambda}$ admits a left inverse (which occurs if and only if $B_{\Lambda}$ has column full rank). Similarly, $\theta$ uniquely determines $\sigma^{\pi}$ through the same equation if and only if $A_{\Lambda}$ is non-singular. These properties of an information structure will be explored in the next section.
4.2 No disclosure, (aggregative) full disclosure, and noise freeness
Our duality results can be utilized to establish necessary and sufficient conditions ensuring the optimality of a given information structure. For example, recall that no disclosure $\underline{\pi}$ results in the action variance matrix $X=O$ and the action-state covariance matrix $Y=O$. Substituting these into (17), the first equation is vacuously satisfied, and the second reduces to $B_{\Lambda}Z=O$. In particular, when $Z$ has full rank, this simplifies to $B_{\Lambda}=O$; namely, for each agent $i$, there is a constant ratio $-\lambda_{i}$ such that $w_{ik}=-\lambda_{i}r_{ik}$ for each $k$, whereas the choice of $\lambda_{1},\ldots,\lambda_{n}$ is subject to the dual feasibility constraint. This requires that each $i$-th row of $W$ forms a proportional vector to the $i$-th row of $R$. Likewise, the optimality of full disclosure $\overline{\pi}$ can be characterized by specifying $X$ and $Y$ as $Q^{-1}RZ(Q^{-1}R)^{\top}$ and $Q^{-1}RZ$, respectively. The next result summarizes these.
Corollary 7.
Suppose $Z\succ O$.
  1. i).
    $\underline{\pi}$ is optimal if and only if there exists $\Lambda\in\mathcal{S}^{n}_{\rm diag}$ such that $A_{\Lambda}\succeq O$ and $B_{\Lambda}=O$.
  2. ii).
    $\overline{\pi}$ is optimal if and only if there exists $\Lambda\in\mathcal{S}^{n}_{\rm diag}$ such that $A_{\Lambda}\succeq O$ and $A_{\Lambda}Q^{-1}R=B_{\Lambda}$.
Both $\underline{\pi}$ and $\overline{\pi}$ have the notable characteristic that the designer does not use randomization beyond the randomness in $\theta$, since $\sigma^{\pi}$ is deterministic conditional on the true state. In general, an information structure $\pi$ has this characteristic if its BNE $\sigma^{\pi}$ is represented as a function of $\theta$, that is, the conditional variance of actions given $\theta$ is zero. If this is the case, we refer to $\pi$ as noise free.
Definition 4.
An information structure $\pi\in\Pi$ is noise free if $\operatorname{Var}[\sigma^{\pi}\mid\theta]=O$ a.s.
Note that a Gaussian information structure $\pi\in\Pi^{\rm g}$ is noise free if and only if $X=YZ^{+}Y^{\top}$ holds for $X=\operatorname{Var}\quantity[\sigma^{\pi}]$ and $Y=\operatorname{Cov}\quantity[\sigma^{\pi},\theta]$ since $\operatorname{Var}[\sigma^{\pi}\mid\theta]=X-YZ^{+}Y^{\top}$. In this case, the first equation in (19) is automatically satisfied regardless of $\Lambda$. Consequently, the optimality of any noise-free information structure can be determined solely by the second equation in (19). On the other hand, the linear relationship (20) in Corollary 6 implies that the equilibrium action profile $\sigma^{\pi}$ can be identified from the state realization if $A_{\Lambda}$ is invertible. Therefore, whenever certifying Lagrangian multipliers can be chosen in such a way of making $A_{\Lambda}\succ O$, we see that the optimal information structure is necessarily noise free. Moreover, in this case, the universality of certification (Corollary 4) implies that the optimal information structure is essentially unique.
Corollary 8.
Suppose $Z\succ O$.
  1. i).
    A noise-free information structure $\pi\in\Pi^{\rm g}$ is optimal if and only if there exists $\Lambda\in\mathcal{S}^{n}_{\rm diag}$ with $A_{\Lambda}\succeq O$ such that $A_{\Lambda}\operatorname{Cov}\quantity[\sigma^{\pi},\theta]=B_{\Lambda}Z$.
  2. ii).
    If (4.1) is solved by $\Lambda\in\mathcal{S}^{n}_{\rm diag}$ such that $A_{\Lambda}\succ O$, then any optimal $\pi\in\Pi^{\rm g}$ is noise free. Moreover, in this case, the solution to (P) is unique.
Next, we consider another yet related property of information structures that extends the idea of full disclosure $\overline{\pi}$. Consider a scenario where a hypothetical observer can access all the agents’ action recommendations forwarded to each of them. Then, we would interpret $\pi$ as disclosing the true state to the entire “population” when the residual variance of $\theta$ is zero conditional on the equilibrium strategy profile $\sigma^{\pi}_{1},\ldots,\sigma^{\pi}_{n}$. In essence, under such $\pi$, even though the true state may remain unknown to each individual, an information structure reveals it to the population in an aggregative form, reflecting the designer’s maximum use of knowledge about the state to guide the agents toward the designer-optimal outcome. This characteristic of information structures can be formalized by means of the zero residual variance of the state.
Definition 5.
An information structure $\pi\in\Pi$ is aggregatively disclosing if $\operatorname{Var}\quantity[\theta\mid\sigma^{\pi}]=O$ a.s.
As is evident from the definition, aggregative disclosing serves as a counterpart to noise-freeness, where the roles of $\sigma^{\pi}$ and $\theta$ are exchanged. Similarly to before, since a Gaussian information structure $\pi$ is aggregatively disclosing if and only if $Z=Y^{\top}X^{+}Y$ holds for $X=\operatorname{Var}\quantity[\sigma^{\pi}]$ and $Y=\operatorname{Cov}\quantity[\sigma^{\pi},\theta]$, the second equation in (18) is automatically satisfied regardless of $\Lambda$. Consequently, the optimality of such an information structure can be determined solely by the first equation in (18). On the other hand, Corollary 6 implies that $\theta$ can be expressed as a function of $\sigma^{\pi}$ when $B_{\Lambda}$ is a left-invertible matrix. Therefore, whenever certifying Lagrangian multipliers can be constructed in this way, the optimal information structure is aggregatively disclosing.
Corollary 9.
Suppose $Z\succ O$.
  1. i).
    An aggregatively-disclosing information structure $\pi\in\Pi^{\rm g}$ is optimal if and only if there exists $\Lambda\in\mathcal{S}^{n}_{\rm diag}$ with $A_{\Lambda}\succeq O$ such that $A_{\Lambda}\operatorname{Var}\quantity[\sigma^{\pi}]=B_{\Lambda}\operatorname{Cov}\quantity[\theta,\sigma^{\pi}]$.
  2. ii).
[…]
It is worth mentioning that the optimality of no disclosure or full disclosure crucially hinges on the dimension of the state. Specifically, both $\underline{\pi}$ and $\overline{\pi}$ are generically suboptimal if the state $\theta$ is multi-dimensional, i.e., $Z\succ O$ and $m\geq 2$. This is because each of the optimality conditions, $B_{\Lambda}=O$ and $A_{\Lambda}Q^{-1}R=B_{\Lambda}$, consists of $nm$ equations, while there are $n$ variables in $\Lambda=\operatorname{Diag}(\lambda_{1},\ldots,\lambda_{n})$. As a result, the number of equations exceeds the number of variables, meaning that the system of equations has no solution generically.
In contrast, when the state is one-dimensional, Corollary 7 presents qualitatively different implications regarding the optimality of no disclosure, as there generally exists a unique set of Lagrangian multipliers $\lambda_{1},\ldots,\lambda_{n}$ that solves $B_{\Lambda}=O$ as long as each entry of $R$ is nonzero. Thus, $\underline{\pi}$ is optimal if and only if $A_{\Lambda}\succeq O$ holds with $\Lambda=\operatorname{Diag}(\lambda_{1},\ldots,\lambda_{n})$, which can be satisfied for a generic set of payoff structures. Interestingly, when no disclosure is suboptimal, an optimal information structure turns out to be aggregatively disclosing.
Corollary 10.
Suppose $m=1$ and $r_{i}\neq 0$ for all $i$, and let $D\equiv\operatorname{Diag}(w_{1}/r_{1},\ldots,w_{n}/r_{n})$, where $r_{i}$ and $w_{i}$ denote the $i$-th entries of $n$-dimensional vectors $R$ and $W$, respectively.
  1. i).
    No disclosure $\underline{\pi}$ is optimal if and only if $O\succeq V+(DQ+Q^{\top}D)/2$.
  2. ii).
    Otherwise, if $O\not\succeq V+(DQ+Q^{\top}D)/2$, then any optimal information structure $\pi\in\Pi^{\rm g}$ is aggregatively disclosing.
The common-value case of Section 2 has the payoff structure with
$Q=\matrixquantity[1&1/2\\ 1/2&1],\quad R=\matrixquantity[-1/2\\ -1/2],\quad V=\matrixquantity[1&c\\ c&1],\quad W=\matrixquantity[0\\ 0].$
This implies that $D=O$, and thus the matrix $V+(DQ+Q^{\top}D)/2$ simplifies to $V$. Since $V$ has positive diagonal entries, $O\not\succeq V$ holds for any $c\in[0,1]$. Therefore, Corollary 10 explains why no disclosure is neither PS-optimal, TS-optimal, nor CS-optimal in the common-value case. On the other hand, as reported in Table 1, every numerical solution leads to an aggregatively disclosing optimal information structure, whereas each firm is kept partially informed when $c=0$.
4.3 State-independent objective and proportional decision criterion
This section considers an objective function $v(a,\theta)$ that does not depend on the state $\theta$, i.e., $W=O$. We call such an objective function state-independent. A state-independent objective function naturally arises in several applications. In addition, some payoff structure can be equivalently transformed to another payoff structures with a state-independent objective function. This notion of equivalence is formalized as follows.
Definition 6.
Two payoff structures, $\mathscr{P}=(Q,R,V,W)$ and $\tilde{\mathscr{P}}=(\tilde{Q},\tilde{R},\tilde{V},\tilde{W})$, are said to be equivalent if any $(X,Y)$ is feasible under $(Q,R)$ if and only if it is feasible under $(\tilde{Q},\tilde{R})$, and if there exists some $c>0$ such that $\tilde{V}\bullet X+\tilde{W}\bullet Y=c(V\bullet X+W\bullet Y)$ holds for all $(X,Y)$ that are feasible under $(Q,R)$.
In words, two payoff structures are equivalent if they share the same set of feasible equilibrium covariance matrices and yield the same value up to a positive scalar multiplication. For example, compare $\mathscr{P}=(Q,R,V,W)$ and $\tilde{\mathscr{P}}=(\Delta Q,\Delta R,cV,cW)$ with a positive constant $c>0$ and a positive diagonal matrix $\Delta\in\mathcal{S}^{n}_{\rm diag}$. Evidently, the value under $\mathscr{P}$ is propotional to that under $\tilde{\mathscr{P}}$. Moreover, for any $(X,Y)\in\mathcal{S}^{n}\times\mathcal{M}^{n,m}$, it holds that
$\operatorname{diag}\quantity(X\quantity(\Delta Q)^{\top})=\Delta\operatorname{diag}\quantity(XQ^{\top})\quad\text{and}\quad\operatorname{diag}\quantity(Y\quantity(\Delta R)^{\top})=\Delta\operatorname{diag}\quantity(YR^{\top}).$
Since every diagonal entry of $\Delta$ is non-zero, it follows that $\operatorname{diag}(XQ^{\top})=\operatorname{diag}(YR^{\top})$ if and only if $\operatorname{diag}(X(\Delta Q)^{\top})=\operatorname{diag}(Y(\Delta R)^{\top})$, thereby $\mathscr{P}$ and $\tilde{\mathscr{P}}$ being equivalent.
Equivalent payoff structures give rise to essentially the same design problem. Thus, the search for optimal information structures can be facilitated by transforming the original payoff structure into simpler equivalent one. This is possible if a payoff structure has the following property.
Definition 7.
Given $\Delta\in\mathcal{S}^{n}_{\rm diag}$, we say that $\mathscr{P}=(Q,R,V,W)$ satisfies the proportional decision criterion (PDC) with respect to $\Delta$ when $W=\Delta R$ holds.
To interpret, recall that $r_{ik}$ expresses the marginal change in agent $i$’s action $a_{i}$ in response to a unit increase of her best estimate about $\theta_{k}$, while $w_{ik}$ captures the marginal contribution of the correlation between $a_{i}$ and $\theta_{k}$ to the designer’s objective. Hence, the proportionality implies that how the designer values the action-state correlations should align with how each agent values each dimension of the state. For instance, if agent $i$ exhibits twice as much sensitivity in her action to $\theta_{2}$ as to $\theta_{1}$, then the designershould aim for agent $i$’s action to be correlated with $\theta_{2}$ twice as much as with $\theta_{1}$. This condition is satisfied trivially when $m=1$ because the proportionality is independently required for different agents. Note that PDC is a necessary condition for the optimality of no disclosure $\underline{\pi}$ since $B_{\Lambda}=O$ implies $W=-\Lambda R$, which requires that each row of $W$ is proportional to the corresponding row of $R$.
The next lemma shows that if $\mathscr{P}$ satisfies PDC, then there exists an equivalent payoff structure with a state-independent objective function, i.e., $W=O$.
Lemma 3.
Given $\Delta\in\mathcal{S}^{n}_{\rm diag}$, if $\mathscr{P}=(Q,R,V,W)$ satisfies PDC with respect to $\Delta$, then it is equivalent to $(Q,R,V_{\Delta},O)$, where $V_{\Delta}\equiv V+(\Delta Q+Q^{\top}\Delta)/2$.
Given a payoff structure $\mathscr{P}$ satisfying PDC, we refer to $\tilde{\mathscr{P}}=(Q,R,V_{\Delta},O)$ as the state-independent form of $\mathscr{P}$. For example, when $m=n$, if $R=\operatorname{Diag}(r_{1},\ldots,r_{n})$ and $W=\operatorname{Diag}(w_{1},\ldots,w_{n})$ are diagonal matrices, the payoff structure satisfies PDC, and the state-independent form is derived through $\Delta=\operatorname{Diag}(w_{1}/r_{1},\ldots,w_{n}/r_{n})$. Alternatively, if $m=1$, then any payoff structure satisfies PDC, and the state-independent form is again derived through $\Delta=\operatorname{Diag}(w_{1}/r_{1},\ldots,w_{n}/r_{n})$, where $r_{i}$ and $w_{i}$ denote the $i$-th entries of $n$-dimensional vectors $R$ and $W$, respectively.
Example 1 (Continued).
Suppose that agents have the identical payoff functions given by (7), and that an objective function is equal to the identical payoff function plus a normalization constant:
$v(a,\theta)=-\quantity(a-\bar{a})^{\top}Q\quantity(a-\bar{a})+2\quantity(a-\bar{a})^{\top}R\quantity(\theta-\bar{\theta}).$ (21)
Then, the payoff structure is $\mathscr{P}=(Q,R,V,W)$ with $V=-Q$ and $W=2R$, which satisfies PDC with respect to $\Delta=2I$. Thus, $\mathscr{P}$ is equivalent to $\tilde{\mathscr{P}}=(Q,R,V_{\Delta},O)$ with $V_{\Delta}=Q$ by Lemma 3. Consequently, the designer’s problem can be identified as the maximization of $Q\bullet X$ over all feasible $(X,Y)$ under $(Q,R)$. Since $Q$ is a positive definite matrix, the objective function is increasing with respect to the matrix ordering $\succeq$, which leads us to predict the optimality of full disclosure. Formally, this can be confirmed from Corollary 7 by taking $\Lambda=2I$ so as to get $A_{\Lambda}=Q$ and $B_{\Lambda}=R$, with which $A_{\Lambda}Q^{-1}R=B_{\Lambda}$ is satisfied.
Motivated by this example, we define the class of payoff structures that are equivalent to the payoff structure that consists of a team and an objective function (21).
Definition 8.
A payoff structure is called a team payoff structure if it is equivalent to some $\mathscr{P}=(Q,R,V,W)$ with $Q\in\mathcal{S}^{n}_{++}$, $V=Q$, and $W=O$.
It is clear that full disclosure is optimal under a team payoff structure. Furthermore, when the dimension of the state is no less than the number of agents, the converse implication holds, up to some qualifications; that is, the optimality of full disclosure necessitates a team payoff structure.
Corollary 11.
Suppose $Z\succ O$ and $m\geq n$. Let $\mathscr{P}=(Q,R,V,W)$ be a payoff structure satisfying PDC with respect to some $\Delta\in\mathcal{S}^{n}_{\rm diag}$. In addition, suppose that $R$ has full row rank and none of the row vectors of $V$ are zero. Then, the following are equivalent:
  1. i).
    $\overline{\pi}$ is optimal.
  2. ii).
    $V_{\Delta}\succ O$ and $V_{\Delta}Q^{-1}\in\mathcal{S}^{n}_{\rm diag}\cap\mathcal{S}^{n}_{++}$.
  3. iii).
    $\mathscr{P}$ is a team payoff structure.
This result implies that full disclosure is generically suboptimal when $m\geq n$ under payoff structures satisfying PDC. However, this is not the case for no disclosure. Specifically, if a payoff structure $\mathscr{P}$ has a state-independent form $\tilde{\mathscr{P}}=(Q,R,V_{\Delta},O)$, the negative semidefiniteness of $V_{\Delta}$ is sufficient for $\underline{\pi}$ to be optimal, which holds for a generic set of payoff structures satisfying PDC.
Corollary 12.
Suppose $Z\succ O$. Let $\mathscr{P}=(Q,R,V,W)$ be a payoff structure satisfying PDC with respect to some $\Delta\in\mathcal{S}^{n}_{\rm diag}$. Then, $\underline{\pi}$ is optimal if $O\succeq V_{\Delta}$.
The private-value case of Section 2 has the payoff structure given by
$Q=\matrixquantity[1&1/2\\ 1/2&1],\quad R=\matrixquantity[-1/2&0\\ 0&-1/2],\quad V=\matrixquantity[1&c\\ c&1],\quad W=\matrixquantity[0&0\\ 0&0],$
which is a state-independent form. Since $V$ cannot be negative semidefinite, the suboptimality of $\underline{\pi}$ is confirmed from Corollary 12. Moreover, we see that the matrix
$VQ^{-1}=\frac{2}{3}\matrixquantity[2-c&2c-1\\ 2c-1&2-c]$
is diagonal and positive definite if and only if $c=1/2$. Thus, Corollary 11 implies the optimality of $\overline{\pi}$ for $c=1/2$, while full disclosure is neither PS-optimal, TS-optimal, nor CS-optimal.
5 Symmetry in network games
5.1 Symmetric information structures
In the network game discussed in Example 3, the payoff coefficient matrix $Q=I+qG$ is associated with a network $G$ describing the physical links between agents. Analogously, we can interpret the covariance matrix of actions $\operatorname{Var}[\sigma^{\pi}]$ as a representation of the informational links between agents. In this section, we show that if the network has some symmetry, the informational links under optimal information structures exhibit similar symmetry, provided that the objective function does not conflict with the connectivity structure in agents’ payoffs. This result serves as a building block for studying the relationship between optimal informational links and physical links in network games. Furthermore, the symmetric structure facilitates the computation of optimal information structures by allowing us to equalize Lagrangian multipliers across symmetric agents.
Let $G\subseteq[n]^{2}$ be a set of edges with $(i,i)\not\in G$ for all $i\in[n]$, which is simply referred to as a network. An automorphism of $G$ is a permutation $\tau:[n]\to[n]$ such that $(i,j)\in G$ if and only if $(\tau(i),\tau(j))\in G$. We denote the collection of all automorphisms by $\operatorname{Aut}\quantity(G)$ and refer it to as the automorphism group of $G$, since it forms a group with respect to a binary operation $(\tau,\rho)\mapsto\tau\rho\in\operatorname{Aut}(G)$ such that $\tau\rho(i)=\rho(\tau(i))$ for each $i\in N$. We define the $G$-orbit of agent $i$ as
$\operatorname{Orb}\quantity(i\mid G)\equiv\quantity{j\in[n]:\tau(i)=j\text{ for some }\tau\in\operatorname{Aut}(G)},$
which is the set of agents who are reachable from agent $i$ by applying some automorphism $\tau\in\operatorname{Aut}(G)$. One can easily see that each orbit constitutes an equivalence class in $[n]$. An automorphism group $\operatorname{Aut}(G)$ is said to be transitive if $\operatorname{Orb}\quantity(i\mid G)=[n]$ holds for all $i$ (equivalently, for some $i$).
Two agents $i$ and $j$ can be regarded as playing symmetric roles in a network $G$ if $j\in\operatorname{Orb}\quantity(i\mid G)$ since swapping their indices leaves the network unchanged. If there is no automorphism other than the identity mapping, the network holds no symmetric pair of agents. On the other hand, if $G$ is the complete graph, then $\operatorname{Aut}(G)$ is the collection of all permutations, implying the symmetric roles of all agents. Let $K_{n}$ denote the complete graph. In general, the size of $\operatorname{Aut}(G)$ can be understood as capturing the degree of symmetry.
For a permutation $\tau:[n]\to[n]$, let $P_{\tau}\equiv[\mathbf{e}_{\tau(1)}\cdots\mathbf{e}_{\tau(n)}]$ be the associated permutation matrix, where the $i$-th column is the $\tau(i)$-th unit vector. For each $n\times n$ matrix $M$, we write $\tau(M)\equiv P_{\tau}MP_{\tau}^{\top}$ with some abuse of notation, which is the matrix obtained by permuting the rows and columns of $M$ according to the permutation $\tau$. A matrix $M$ is said to be $G$-invariant if $M=\tau(M)$ holds for all $\tau\in\operatorname{Aut}(G)$. For example, the adjacency matrix $G$ is $G$-invariant, and thus the coefficient matrix of the network game $Q=I+qG$ is $G$-invariant.
Based on the notion of $G$-invariance, we first formulate the corresponding symmetry of a network game discussed in Example 3, where we assume $m=n$.
Definition 9.
A basic game $(Q,R,Z)$ is a $G$-network game if $m=n$, $Q=Q_{G}\equiv I_{n}+qG$, and $R=I_{n}$, where $I_{n}$ is the $n\times n$ identity matrix. A $G$-network game $(Q_{G},I_{n},Z)$ is a $G$-invariant game if $Z$ is $G$-invariant.
In a $G$-invariant game, the distribution of payoff states is also aligned with the symmetry present in the network $G$.
Similarly, we define the notion of $G$-invariance for an objective function and an information structure.
Definition 10.
When $m=n$, an objective function $v(a,\theta)$ is $G$-invariant if $V$ and $W$ are $G$-invariant; an information structure $\pi$ is $G$-invariant if $\operatorname{Var}[\sigma^{\pi}]$ and $\operatorname{Cov}[\sigma^{\pi},\theta]$ are $G$-invariant.
The next proposition shows that there exists a $G$-invariant optimal information structure under a $G$-invariant game and a $G$-invariant objective function. Moreover, there exists a $G$-invariant diagonal matrix $\Lambda$ certifying the $G$-invariant optimal information structure. This result also gives sufficient conditions for noise-freeness.
Proposition 13.
Suppose that a basic game and an objective function are $G$-invariant. Then, there exists a Gaussian information structure $\pi$ that is $G$-invariant and optimal, and there exists a $G$-invariant diagonal matrix $\Lambda=\operatorname{Diag}(\lambda_{1},\ldots,\lambda_{n})$ that certifies $\pi$, i.e., $\lambda_{i}=\lambda_{\tau(i)}$ for all $i\in[n]$ and $\tau\in\operatorname{Aut}(G)$. If $Z$ is non-singular and $\operatorname{Aut}(G)$ is transitive, the following holds.
  1. i).
    If there is no constant $c\in\mathbb{R}$ such that both $cI_{n}+W$ and $cQ_{G}-V$ are singular, then the optimal $G$-invariant information structure is noise free.
  2. ii).
    If $\mathscr{P}$ satisfies PDC, then the optimal $G$-invariant information structure is noise free.
For example, consider the case of the complete graph $G=K_{n}$. Note that a matrix is $K_{n}$-invariant if and only if all diagonal elements are identical and all off-diagonal elements are identical, since the automorphism group $\operatorname{Aut}(K_{n})$ consists of all permutations. The payoff coefficient matrix of a $K_{n}$-network game is
$Q_{K_{n}}=\matrixquantity[1&q&\cdots&q\\ q&1&\cdots&q\\ \vdots&\vdots&\ddots&\\ q&q&\cdots&1],$ (22)
where $-1/(n-1)<q<1$ since the payoff coefficient matrix is assumed to be positive definite. According to Proposition 13, if $Z$, $V$, and $W$ are $K_{n}$-invariant, then there exists an optimal $K_{n}$-invariant Gaussian information structure $\pi$, where $\operatorname{Var}[\sigma_{i}^{\pi}]$ and $\operatorname{Cov}[\sigma_{i}^{\pi},\theta_{i}]$ are uniform across all agents $i$, and $\operatorname{Cov}[\sigma_{i}^{\pi},\sigma_{j}^{\pi}]$ and $\operatorname{Cov}[\sigma_{i}^{\pi},\theta_{j}]$ are uniform across all pairs of agents $\{i,j\}$. Moreover, there exists a constant $\lambda_{0}\in\mathbb{R}$ such that $\lambda_{0}I_{n}$ certifies the optimal $K_{n}$-invariant information structure.
This result substantially facilitates the computation of an optimal information structure. To illustrate, we assume a payoff structure satisfies PDC, so an equivalent payoff structure is given by
$V=\matrixquantity[v&c&\cdots&c\\ c&v&\cdots&c\\ \vdots&\vdots&\ddots&\\ c&c&\cdots&v],\quad W=O.$ (23)
By Corollaries 11 and 12, full disclosure is optimal if and only if $v>0$ and $q=c/v$; no disclosure is optimal if and only if $V$ is negative semidefinite, i.e., $v\leq c\leq-v/(n-1)$.
We consider the other case, where partial disclosure is optimal. To simplify the computation, we assume $Z=I_{n}$. Using the $K_{n}$-invariance of an optimal information structure, we can write the optimal feasible pair $(X,Y)$ as
$X=\matrixquantity[x&\rho x&\cdots&\rho x\\ \rho x&x&\cdots&\rho x\\ \vdots&\vdots&\ddots&\\ \rho x&\rho x&\cdots&x],\ Y=\matrixquantity[y&\delta y&\cdots&\delta y\\ \delta y&y&\cdots&\delta y\\ \vdots&\vdots&\ddots&\\ \delta y&\delta y&\cdots&y].$ (24)
By Proposition 13, the optimal information structure is noise free, i.e.,
$\operatorname{Var}[\sigma^{\pi}|\theta]=X-YZ^{-1}Y^{\top}=X-YY^{\top}=O,$
which implies $X$ given by
$x=y^{2}((n-1)\delta^{2}+1),\ \rho=\frac{\delta(\delta(n-2)+2)}{\delta^{2}(n-1)+1}.$ (25)
By plugging (25) into the first equation of (17), where
$A_{\Lambda}=\matrixquantity[\lambda_{0}-v&\lambda_{0}q-c&\cdots&\lambda_{0}q-c\\ \lambda_{0}q-c&\lambda_{0}-v&\cdots&\lambda_{0}q-c\\ \vdots&\vdots&\ddots&\\ \lambda_{0}q-c&\lambda_{0}q-c&\cdots&\lambda_{0}-v],\ B_{\Lambda}=\lambda_{0}I_{n},$
we can derive $y$ and $\lambda_{0}$ as functions of $\delta$:
$y=-\frac{\delta((n-2)c+v)+c}{2(\delta-1)(\delta(n-1)+1)(c-qv)},\,\lambda_{0}=\frac{\delta((n-2)c+v)+c}{\delta+\delta(n-2)q+q}.$ (26)
Substituting (25) and (26) into the constraint $\operatorname{diag}(XQ^{\top})=\operatorname{diag}(YR^{\top})$, we obtain the following equation for $\delta$:
$\frac{\delta((n-2)c+v)+c}{4(\delta-1)^{2}(\delta(n-1)+1)^{2}(c-qv)^{2}}\times f(\delta)=0,$ (27)
where $f(\delta)=c_{3}\delta^{3}+c_{2}\delta^{2}+c_{1}\delta+c_{0}=0$ is a cubic function with
$c_{3}$ $=c(n-1)(n-2)^{2}q+c(n-1)(n-2)+(n-1)(n-2)qv+(n-1)v,$
$c_{2}$ $=2(n-1)(c-qv)+3c(n-2)(n-1)q+c(n-1)+2(n-1)qv,$
$c_{1}$ $=-2(n-1)(c-qv)+2c(n-1)q+c(n-2)+2(c-qv)+v,$
$c_{0}$ $=c-2(c-qv).$
One of the solutions to (27) is $\delta=-c/((n-2)c+v)$, which means $x=y=0$ by (25) and (26), i.e., no disclosure is optimal. Since we assume partial disclosure is optimal, the optimal value of $\delta$ is a real solution to the cubic equation $f(\delta)=0$. When $n$, $q$, $v$, and $c$ are given, we can solve the equation to obtain $\delta$ and identify the optimal covariance matrix by plugging $\delta$ back to (25) and (26).
For example, consider the case where $n=2$, $v=1$, and $q={1}/{2}$, corresponding to Section 2. Figure 1(a) illustrates the graphs of $\operatorname{Var}[\sigma_{1}^{\pi}]$ and $\operatorname{Cov}[\sigma_{1}^{\pi},\sigma_{2}^{\pi}]$ as functions of $c$. The covariance $\operatorname{Cov}[\sigma_{1}^{\pi},\sigma_{2}^{\pi}]$ is increasing in $c$ since $c$ is the weight of $\operatorname{Cov}[\sigma_{1}^{\pi},\sigma_{2}^{\pi}]$ in the objective function. In contrast, the variance $\operatorname{Var}[\sigma_{1}^{\pi}]$ is increasing in $c$ if $c<0$ and decreasing in $c$ if $c>0$, since when $c=0$ the objective function equals $\operatorname{Var}[\sigma_{1}^{\pi}]$.
(a) Private-value
(b) Common-value
Figure 1: The graphs of $\operatorname{Var}[\sigma_{1}^{\pi}]$ and $\operatorname{Cov}[\sigma_{1}^{\pi},\sigma_{2}^{\pi}]$ as functions of $c$, where $v=1$ and $q=1/2$.
5.2 The case of a state reduction form
In a $G$-network game in the previous section, each agent has her own payoff state $\theta_{i}$, which is a heterogeneous component of agents’ payoff functions. Although $Z$ is an $n\times n$ matrix, if the rank of $Z$ is less than $n$, we can represent the $G$-network game with a less number of payoff states, and such a representation is more useful in computing an optimal information structure numerically or analytically. For example, if $\theta_{1}=\cdots=\theta_{n}$ almost surely, then every entry of the covariance matrix $Z=\operatorname{Var}[\theta]$ is identical. In this case, it is enough to study an equivalent basic game with $m=1$, i.e., $(Q_{G},\mathbf{1},z)$, where $z$ is the identical entry of $Z$.
To study a general case, let $m<n$ be the rank of $Z$. Then, there exist a full rank $n\times m$ matrix $\tilde{R}$ and i.i.d. random variables $\tilde{\theta}_{1},\ldots,\tilde{\theta}_{m}$ that follow the standard normal distribution such that $\theta=\tilde{R}\tilde{\theta}$ and $Z=\tilde{R}\tilde{R}^{\top}$. Thus, a $G$-network game $(Q_{G},I_{n},\tilde{R}\tilde{R}^{\top})$ is equivalently written as $(Q_{G},\tilde{R},I_{m})$, where agent $i$’s payoff function is
$-a_{i}^{2}+2qa_{i}\sum_{j:(i,j)\in G}a_{j}+2(\tilde{R}\tilde{\theta})_{i}a_{i}.$
We refer to $(Q,\tilde{R},I_{m})$ as a sate reduction form of a $G$-network game. The equivalence of a $G$-network game and its state reduction form is formally stated in the next lemma.
Lemma 4.
Let $(Q_{G},\tilde{R},I_{m})$ be a state reduction form of a $G$-network game $(Q_{G},I_{n},\tilde{R}\tilde{R}^{\top})$. If a pair of matrices $(X,Y)\in\mathcal{S}^{n}\times\mathcal{M}^{n,n}$ is feasible under $(Q_{G},I_{n})$, then there exists $\tilde{Y}\in\mathcal{M}^{n,m}$ such that $Y=\tilde{Y}\tilde{R}^{\top}$ and $(X,\tilde{Y})\in\mathcal{S}^{n}\times\mathcal{M}^{n,m}$ is feasible under $(Q_{G},\tilde{R})$. Conversely, if a pair of matrices $(X,\tilde{Y})\in\mathcal{S}^{n}\times\mathcal{M}^{n,m}$ is feasible under $(Q_{G},\tilde{R})$, then $(X,\tilde{Y}\tilde{R}^{\top})\in\mathcal{S}^{n}\times\mathcal{M}^{n,n}$ is feasible under $(Q_{G},I_{n})$.
This lemma implies
$\max_{\text{$(X,Y)$ is feaslble}}V\bullet X+W\bullet Y=\max_{\text{$(X,\tilde{Y})$ is feaslble}}V\bullet X+W\tilde{R}^{\top}\bullet\tilde{Y}$
because if $Y=\tilde{Y}\tilde{R}^{\top}$ then
[…]
Thus, a feasible pair $(X,Y)$ is optimal under a $G$-network game and an objective function with $(V,W)$ if and only if the corresponding feasible pair $(X,\tilde{Y})$ is optimal under the state reduction form $(Q_{G},\tilde{R},I_{m})$ and the objective function with $(V,W\tilde{R})$. We also refer to the latter objective function as a state reduction form of the original objective function.
The next proposition shows that we can obtain an optimal $G$-invariant information structure by solving the state reduction form of a $G$-invariant game and a $G$-invariant objective function.
Proposition 14.
Consider the state reduction form of a $G$-invariant game and a $G$-invariant objective function. Then, there exists an optimal information structure $\tilde{\pi}$ with an optimal feasible pair $(X,\tilde{Y})$ such that $X$ and $\tilde{Y}\tilde{R}^{\top}$ are $G$-invariant, and there exists a $G$-invariant diagonal matrix $\Lambda=\operatorname{Diag}(\lambda_{1},\ldots,\lambda_{n})$ that certifies $\tilde{\pi}$. In particular, if $\operatorname{Aut}(G)$ is transitive, then $\lambda_{1}=\cdots=\lambda_{n}$.
To illustrate how to use Proposition 14, we assume $m=1$ and calculate an optimal information structure under a $K_{n}$-invariant game $(Q_{K_{n}},I_{n},1)$ and a $K_{n}$-invariant objective function. The state reduction form of the $K_{n}$-invariant game is $(Q_{K_{n}},\mathbf{1},1)$ and that of the $K_{n}$-invariant objective function is given by
$V=\matrixquantity[v&c&\cdots&c\\ c&v&\cdots&c\\ \vdots&\vdots&\ddots&\\ c&c&\cdots&v],\quad W=w\mathbf{1},$
where $w\in\mathbb{R}$. Since a payoff structure satisfies PDC whenever $m=1$, we assume $w=0$ without loss of generality, i.e, $W=\mathbf{0}$. This state reduction form of the problem corresponds to the common-value case of Section 2. By Proposition 14, there exists a constant $\lambda_{0}$ such that $\lambda_{0}I_{n}$ certifies an optimal information structure with an optimal feasible pair $(X,\tilde{Y})$ such that $X$ and $\tilde{Y}\mathbf{1}^{\top}$ are $K_{n}$-invariant.
Full disclosure is optimal if
$\lambda_{0}Q-V\succeq O\text{ and }(\lambda_{0}Q-2V)\mathbf{1}=\mathbf{0}$ (28)
by Corollary 7, which is equivalent to the condition
$c\geq\min\left\{\frac{v(1-(n+1)q)}{(n-1)q+1-2n},-\frac{v}{n-1}\right\}.$
For example, when $v=1$, $q=1/2$, and $n=2$, full disclosure is optimal if and only if $c\geq 1/5$. This explains why full disclosure is TS-optimal ($c=1/3$) and CS-optimal ($c=1$) in the example of Section 2. No disclosure is optimal if and only if $V$ is negative semidefinite, i.e., $v\leq c\leq-v/(n-1)$.
We consider the other case, where partial disclosure is optimal. Using the $K_{n}$-invariance of $X$ and $\tilde{Y}\mathbf{1}$, we can write the optimal feasible pair $(X,\tilde{Y})$ as
$X=\matrixquantity[x&\rho x&\cdots&\rho x\\ \rho x&x&\cdots&\rho x\\ \vdots&\vdots&\ddots&\\ \rho x&\rho x&\cdots&x],\ \tilde{Y}=\matrixquantity[y\\ y\\ \vdots\\ y].$ (29)
We use (19), where we write
$A_{\Lambda}=\matrixquantity[\lambda_{0}-v&\lambda_{0}q-c&\cdots&\lambda_{0}q-c\\ \lambda_{0}q-c&\lambda_{0}-v&\cdots&\lambda_{0}q-c\\ \vdots&\vdots&\ddots&\\ \lambda_{0}q-c&\lambda_{0}q-c&\cdots&\lambda_{0}-v],\ B_{\Lambda}=\lambda_{0}\mathbf{1}.$
Then, we can derive $\lambda_{0}$ and $x$ as functions of $y$ and $\rho$ from the first equation of (19):
$\lambda_{0}=\frac{v-c}{1-q},\ x=\frac{ny^{2}}{(n-1)\rho+1}.$
Plugging this into the second equation of (19), we can solve for
$y=\frac{c-v}{2n(c-qv)}.$
Substituting the above into the constraint $\operatorname{diag}(XQ^{\top})=\operatorname{diag}(\tilde{Y}\tilde{R}^{\top})$, we obtain
$\rho=\frac{c-2qv+v}{(n-1)(c(q-2)+qv)},\ x=\frac{(c-v)(c(q-2)+qv)}{4n(q-1)(c-qv)^{2}}.$
For example, consider the case where $n=2$, $v=1$, and $q={1}/{2}$, corresponding to Section 2. Figure 1(b) illustrates the graphs of $\operatorname{Var}[\sigma_{1}^{\pi}]$ and $\operatorname{Cov}[\sigma_{1}^{\pi},\sigma_{2}^{\pi}]$ as functions of $c$. If $c\leq 1/5$, the covariance $\operatorname{Cov}[\sigma_{1}^{\pi},\sigma_{2}^{\pi}]$ is increasing in $c$, and the variance $\operatorname{Var}[\sigma_{1}^{\pi}]$ is increasing in $c$ if $c<0$ and decreasing in $c$ if $c>0$, which is similar to the private-value case. However, if $c\geq 1/5$, full disclosure is optimal and actions are perfectly correlated, where $\operatorname{Cov}[\sigma_{1}^{\pi},\sigma_{2}^{\pi}]=\operatorname{Var}[\sigma_{1}^{\pi}]$ is constant.
Angeletos and Pavan (2007), Bergemann and Morris (2013), and Ui and Yoshizawa (2015) study similar classes of symmetric LQG games of common value. Angeletos and Pavan (2007) and Ui and Yoshizawa (2015) focus on symmetric Gaussian information structures, where each player’s signal consists of an idiosyncratic private signal and a public signal, and analyze the optimal combination of them from the perspective of social welfare maximization. Bergemann and Morris (2013) characterize the set of all symmetric Bayes correlated equilibria in which actions and a state are jointly normally distributed. Utilizing this result, they derive the optimal symmetric Gaussian information structure in a homogeneous product Cournot game that maximizes the expected total profit. However, the question of whether this structure is also optimal among all information structures, including those that may be asymmetric and non-Gaussian, has not yet been addressed. Proposition 14 together with the above computation confirms that no asymmetric or non-Gaussian information structures can outperform the symmetric Gaussian information structure discussed in Bergemann and Morris (2013).
6 Public information structures
This section assumes that feasible information structures are limited to public ones, an important assumption given real-world contexts where only public information structures are feasible. For example, consider a central bank that distributes information to firms and consumers in an economy. Tailoring this information to each individual firm or consumer is impractical, not only because of technical constraints, but also because of the central bank’s principles of transparency.
An information structure $\pi$ is said to be public if $t_{0}\equiv t_{1}=\cdots=t_{n}$ occurs with probability one, where $t_{0}$ is called a public signal. Let $\Pi^{\rm pb}\subseteq\Pi^{*}$ denote the collection of all public information structures. We study the optimization problem,
$v^{\rm pb}=\max_{\pi\in\Pi^{\rm pb}}\operatorname{E}[v(\sigma^{\pi},\theta)],$
which is referred to as public LQG information design. Henceforth, we assume that $Z\succ O$, and let $Z^{1/2}$ denote the matrix root of $Z$, i.e., the unique matrix satisfying the properties $Z^{1/2}\succ O$ and $Z^{1/2}Z^{1/2}=Z$.
As in the case of private information structures, our first task is to derive the SDP formulation of public LQG information design. A crucial aspect of public information is that agents do not face any strategic uncertainty about the opponents’ actions, which simplify the characterization of a BNE and its induced moments. Specifically, under a public information structure, the first-order condition (8) is
$\sum_{j=1}^{n}q_{ij}\sigma^{\pi}_{j}=\sum_{k=1}^{m}r_{ik}\operatorname{E}\quantity[\theta_{k}\mid t_{0}],\quad\forall i\in N,$
involving no conditional expectations in the left-hand side. From this system of equations, the unique BNE is characterized by $\sigma^{\pi}(t_{0})=Q^{-1}R\operatorname{E}\quantity[\theta\mid t_{0}]$.
Calculating the equilibrium action-state covariance from this explicit expression of BNE, we can express $\mathbf{M}^{\pi}$ as a function of the state variance reduction, $\operatorname{Var}\quantity[E\quantity[\theta\mid t_{0}]]$, which measures the amount of the state variance that can be reduced by observing the public signal $t_{0}$. By statistical requirements, the state variance reduction necessarily satisfies $Z\succeq\operatorname{Var}\quantity[E\quantity[\theta\mid t_{0}]]\succeq O$. Conversely, it is proved in the next lemma that any symmetric matrix that satisfies these matrix inequalities can be induced as an equilibrium action-state covariance matrix under some public Gaussian information structure.
Lemma 5.
For any public information structure $\pi\in\Pi^{\rm pb}$, if we write $S=\operatorname{Var}[\operatorname{E}\quantity[\theta\mid t_{0}]]$, then it holds that $Z\succeq S\succeq O$, and $\mathbf{M}^{\pi}$ coincides with
$\mathbf{M}_{S}\equiv\matrixquantity[Q^{-1}RS(Q^{-1}R)^{\top}&Q^{-1}RS\\ S(Q^{-1}R)^{\top}&Z]\in\mathcal{S}_{+}^{n+m}.$
Conversely, for any $S\in\mathcal{S}^{m}$ such that $Z\succeq S\succeq O$, there exists a public Gaussian information structure $\pi\in\Pi^{\rm pb}\cap\Pi^{\rm g}$ such that $\mathbf{M}^{\pi}=\mathbf{M}_{S}$.
Additionally, the next lemma shows that the state variance reduction serves as a sufficient statistic for determining the designer’s expected value of each possible choice of public information structures.
Lemma 6.
For any public information structure $\pi\in\Pi^{\rm pb}$, if we write $S=\operatorname{Var}[\operatorname{E}\quantity[\theta\mid t_{0}]]$, then the designer’s expected objective is expressed as
$\operatorname{E}\quantity[v(\sigma^{\pi},\theta)]=\overline{C}\bullet S,$
where $\overline{C}\in\mathcal{S}^{m}$ is defined in (15).
As similarly to the previous private information case, these lemmas deliver pivotal implications for simplifying the public LQG information design. First, Lemma 5 implies that the set of inducible covariance matrices under all public information structures coincides with that under all public Gaussian ones, since
$\quantity{\mathbf{M}^{\pi}:\pi\in\Pi^{\rm pb}}=\quantity{\mathbf{M}_{S}:Z\succeq S\succeq O}=\quantity{\mathbf{M}^{\pi}:\pi\in\Pi^{\rm pb}\cap\Pi^{\rm g}}.$
In addition, since Lemma 6 implies that the designer’s objective depends on the choice of public information structures only through the state variance reduction $S$, it follows that the designer can focus on public Gaussian information structures among all public ones without losing optimality.
Second, the public LQG information design problem can be restated as the maximization of the linear objective function $\overline{C}\bullet S$ subject to the two-sided positive semidefiniteness constraints of the form $Z\succeq S\succeq I$. Moreover, by rescaling, this SDP problem turns out to be equivalent to
$v^{\rm pb}=\max_{S\in\mathcal{S}^{m}}\widehat{C}\bullet S\quad\text{subject to}\quad I\succeq S\succeq O,$ (PB)
where
$\widehat{C}\equiv Z^{\frac{1}{2}}\overline{C}Z^{\frac{1}{2}}.$
To see the equivalence, let $\widehat{C}$ be given as above, and let $\widehat{S}=Z^{-1/2}SZ^{-1/2}$. Then, it should be easily confirmed that $Z\succeq S\succeq O$ if and only if $I\succeq\widehat{S}\succeq O$, and that $\overline{C}\bullet S=\widehat{C}\bullet\widehat{S}$, from which we can justify (PB) as the equivalent restatement of the original one.
From (PB), we readily see that if $\widehat{C}$ is negative semidefinite, then $\widehat{C}\bullet S\leq 0$ for all feasible $S$, thereby no disclosure being optimal. Conversely, if $\widehat{C}$ is positive semidefinite, then the objective is non-decreasing with respect to the matrix ordering $\succeq$, whence it is optimal to choose $S=I$, corresponding to full disclosure. For instance, if $m=1$, then $\widehat{C}$ is a scalar variable, whose sign is unambiguously determined. Consequently, our discussion implies that either of the two extremes, $\underline{\pi}$ and $\overline{\pi}$, is optimal in $\Pi^{\rm pb}$ when the state is one-dimensional.
In general, when the state is multi-dimensional, $\widehat{C}$ would be neither positive nor negative semidefinite, thus, partial disclosure can be optimal. Specifically, following the methodology developed by Tamura (2018), we can characterize an optimal public information structure and quantify the informativeness of an optimal public signal by exploiting the eigendecomposition of $\widehat{C}$.
Allowing for geometric multiplicity, we denote by $\gamma_{1},\ldots,\gamma_{m}$ the $m$ real eigenvalues of $\widehat{C}$ and by $\mathbf{u}_{1},\ldots,\mathbf{u}_{m}$ the corresponding eigenvectors that constitute an orthonormal system in $\mathbb{R}^{m}$. We align the eigenvalues in a non-increasing order, and let $k^{*}$ be the last index such that the corresponding eigenvalue exceeds 00, i.e.,
$\gamma_{1}\geq\gamma_{2}\geq\cdots\geq\gamma_{k^{*}}>0\geq\gamma_{k^{*}+1}\geq\cdots\geq\gamma_{m},$
where the convention $k^{*}=0$ is adopted when $\widehat{C}$ is negative semidefinite. Then by the spectral theorem, $\widehat{C}$ can be factorized as follows:
$\widehat{C}=U\operatorname{Diag}(\gamma_{1},\ldots,\gamma_{m})U^{\top},$ (30)
where $U\equiv[\mathbf{u}_{1}\cdots\mathbf{u}_{m}]$ is an $m\times m$ orthogonal matrix, which has each $\mathbf{u}_{k}$ as the $k$-th column vector. In addition, let $U_{k^{*}}\equiv[\mathbf{u}_{1}\cdots\mathbf{u}_{k^{*}}]\in\mathcal{M}^{m,k^{*}}$.
Proposition 15.
Let $\pi$ be a public Gaussian information structure, which sends a public signal $t_{0}=U_{k^{*}}Z^{-1/2}\theta$. Then $\pi$ is optimal in $\Pi^{\rm pb}$, and the associated state variance reduction is given by
$\operatorname{Var}\quantity[\operatorname{E}\quantity[\theta\mid t_{0}]]=Z^{\frac{1}{2}}U_{k^{*}}U_{k^{*}}^{\top}Z^{\frac{1}{2}}.$ (31)
Moreover, $v^{\rm pb}$ is equal to the sum of all positive eigenvalues of $\widehat{C}$, i.e., $\gamma_{1}+\cdots+\gamma_{k^{*}}$. Finally, $k^{*}$ satisfies $p(\mathbf{V})-n\leq k^{*}\leq p(\mathbf{V})$, where $p(\mathbf{V})$ is the number of positive eigenvalues of $\mathbf{V}$.
This proposition characterizes an optimal public information structure by deriving an explicit form of the public signal. Specifically, the optimal public signal is determined by the spectral property of the matrix $\widehat{C}$, which is defined through the primitives. Notice that any optimal public information structure is noise free, as the signal being a linear transformation of the state. In addition, it is shown that the value of public LQG information design equals the sum of the positive eigenvalues of $\widehat{C}$.
If $\widehat{C}$ is positive definite, i.e., if all the eigenvalues are positive, then full disclosure is optimal because (31) reduces to $Z$ when $k^{*}=m$. In contrast, if $\widehat{C}$ is negative semidefinite, i.e., if all the eigenvalues are nonpositive, then no disclosure is optimal. In general, $\widehat{C}$ is neither positive nor negative semidefinite, yet, the optimal public signal is identified as a multi-dimensional statistics, whose dimension coincides with the number of positive eigenvalues of $\widehat{C}$. Specifically, there is a monotonic relationship between the number of positive eigenvalues of $\widehat{C}$ and the informativeness of an optimal public signal, as the associated state variance reduction (31) is increasing in $k^{*}$ with respect to the matrix ordering. In other words, as $\widehat{C}$ gets closer to a positive definite matrix, the optimal signal discloses more information about $\theta$.
Implementing an optimal public information structure requires the designer’s detailed knowledgeable about the basic game in play since the key matrix $\widehat{C}$ hinges on the entire payoff structure. In practice, however, $Q$ and $R$ may be agents’ private information, as being parts of their payoff functions, and the designer may be agnostic on these parameters. Still, in Proposition 15, we derive some relationships between the number of positive eigenvalues of $\widehat{C}$ and that of $\mathbf{V}$. This enables the designer to, at least, partially identify the optimal degree of public information disclosure from her objective function alone. Since the tightness of the bounds on $k^{*}$ depends on $n$, such partial identification tends to be accurate when the number of agents is small relative to $m$.
Tamura (2018) studies an information design problem with a single player under a quadratic specification of the designer’s objective. His model assumes that the individual’s best response is given as the best estimate about a multi-dimensional state, $\operatorname{E}[\theta\mid t]$, upon receiving a signal $t$. He then shows that the designer’s objective maximization is reduced to the SDP problem,
$\max_{S\in\mathcal{S}^{m}}W\bullet S\quad\text{subject to}\quad\operatorname{Var}\quantity[\theta]\succeq S\succeq O,$ (32)
where $W\in\mathcal{S}^{m}$ is a constant matrix, and obtains a closed form solution to the problem. He also demonstrates that this SDP formulation is useful in studying an LQG network game with $n$ players and obtains an optimal public information structure in a special case of an LQG network game when $m=n$. In this regard, Proposition 15 generalizes his result by allowing for an arbitrary $m$.
We provide a full proof of Proposition 15 in the appendix of this paper for the sake of completeness, while the solution method of Tamura (2018) can be utilized to establish the optimality of the public information structure. Our proof is more direct and simpler than that of Tamura (2018) because we use the properties of normal distributions, which he does not assume in solving (32).
To conclude, we address a natural question of what conditions are needed for the optimal public information structure to be “globally optimal” among all information structures, including non-public ones.
Corollary 16.
The optimal public information structure in Proposition 15 attains $v^{\rm p}$ if and only if there exists $\Lambda\in\mathcal{S}^{n}_{\rm diag}$ with $A_{\Lambda}\succeq O$ such that
$\quantity(A_{\Lambda}Q^{-1}R-B_{\Lambda})Z^{\frac{1}{2}}\mathbf{u}_{k}=\mathbf{0},\quad\forall k=1,\ldots,k^{*};\quad{\rm and}$ (33)
$B_{\Lambda}Z^{\frac{1}{2}}\mathbf{u}_{k}=\mathbf{0},\quad\forall k=k^{*}+1,\ldots,m.$ (34)
This corollary results from observing that the optimal public information structure is noise free since it yields a public signal that is a linear transformation of the state. Therefore, we can apply Corollary 8 to derive the condition for the optimal public information structure to be globally optimal in terms of its certifiability. Incidentally, the optimality condition involves both matrix equations in Corollary 7, which characterize the optimality of $\underline{\pi}$ and $\overline{\pi}$, while each equation only needs to be partially satisfied on the linear subspace spanned by $\{Z^{1/2}\mathbf{u}_{1},\ldots,Z^{1/2}\mathbf{u}_{k^{*}}\}$ and $\{Z^{1/2}\mathbf{u}_{k^{*}+1},\ldots,Z^{1/2}\mathbf{u}_{m}\}$, respectively. [^21] Note that when $k^{*}=0$ so that $\underline{\pi}$ is the optimal public information structure, (33) is vacuously satisfied, while (34) is equivalent to $B_{\Lambda}=O$ since $\{Z^{1/2}\mathbf{u}_{1},\ldots,Z^{1/2}\mathbf{u}_{m}\}$ constitutes a basis of $\mathbb{R}^{m}$. Similarly, when $k^{*}=m$, it can be seen that (33) and (34) hold if and only if $A_{\Lambda}Q^{-1}R=B_{\Lambda}$. Therefore, Corollary 16 generalizes Corollary 7. Nevertheless, since (33) and (34) collectively impose $nm$ number of linear equations on $n$ Lagrangian multipliers, the optimal public information structure is not globally optimal for generic payoff structures. This indicates that the designer typically benefits from manipulating each agent’s higher-order belief about others’ private information, aligning with the findings of Arieli and Babichenko (2019) and Hoshino (2022).
Appendix
Appendix A Calculations for Section 2
Recall that in the leading example, the designer’s expected objective is given by
$(1-\lambda)\underbrace{\operatorname{E}\quantity[(1-\sigma^{\pi}_{1}-\sigma^{\pi}_{2})(\sigma^{\pi}_{1}+\sigma^{\pi}_{2})-\theta_{1}\sigma^{\pi}_{1}-\theta_{2}\sigma^{\pi}_{1}]}_{\text{(E-PS)}}+\lambda\underbrace{\operatorname{E}\quantity[\frac{(\sigma^{\pi}_{1}+\sigma^{\pi}_{1})^{2}}{2}]}_{\text{(E-CS)}}.$
We show that this expression is proportional to (2), up to positive scalar multiplications and additions of constant terms.
By the definition of (co)variance, we can calculate (E-PS) as follows:
(E-PS) $=\operatorname{E}\quantity[\sigma^{\pi}_{1}+\sigma^{\pi}_{2}]-\operatorname{E}\quantity|\sigma^{\pi}_{1}+\sigma^{\pi}_{2}|^{2}-\operatorname{E}\quantity[\theta_{1}\sigma^{\pi}_{1}]-\operatorname{E}\quantity[\theta_{2}\sigma^{\pi}_{2}]$
$=-\quantity(\operatorname{E}\quantity|\sigma^{\pi}_{1}|^{2}+2\operatorname{E}\quantity[\sigma_{1}^{\pi}\sigma_{2}^{\pi}]+\operatorname{E}\quantity|\sigma^{\pi}_{2}|^{2}+\operatorname{E}\quantity[\theta_{1}\sigma^{\pi}_{1}]+\operatorname{E}\quantity[\theta_{2}\sigma^{\pi}_{2}])+\operatorname{E}\quantity[\sigma^{\pi}_{1}+\sigma^{\pi}_{2}]$
$=-\quantity(\operatorname{Var}\quantity[\sigma_{1}^{\pi}]+2\operatorname{Cov}\quantity[\sigma_{1}^{\pi},\sigma_{2}^{\pi}]+\operatorname{Var}\quantity[\sigma_{2}^{\pi}]+\operatorname{Cov}\quantity[\sigma_{1}^{\pi},\theta_{1}]+\operatorname{Cov}\quantity[\sigma_{2}^{\pi},\theta_{2}])+L$
$=\operatorname{Var}\quantity[\sigma_{1}^{\pi}]+\operatorname{Var}\quantity[\sigma_{2}^{\pi}]+L,$
where $L$ collectively denotes all terms that depend only on $\operatorname{E}[\sigma_{1}^{\pi}]$ and $\operatorname{E}[\sigma_{2}^{\pi}]$, and the last line is obtained by substituting (1). Likewise, (E-CS) can be calculated as follows:
(E-CS) $=\frac{1}{2}\operatorname{Var}\quantity[\sigma_{1}^{\pi}]+\operatorname{Cov}\quantity[\sigma_{1}^{\pi},\sigma_{2}^{\pi}]+\frac{1}{2}\operatorname{Var}\quantity[\sigma_{2}^{\pi}]+L,$
Therefore, it follows that
[…]
$\quad=\quantity(1-\lambda/2)\quantity(\operatorname{Var}\quantity[\sigma_{1}^{\pi}]+2\cdot\frac{\lambda/2}{1-\lambda/2}\operatorname{Cov}\quantity[\sigma_{1}^{\pi},\sigma_{2}^{\pi}]+\operatorname{Var}\quantity[\sigma_{2}^{\pi}])+L,$
which is confirmed to be proportional to (2) by setting $c\equiv\lambda/(2-\lambda)$.
Appendix B Semidefinite programming
Given $C,A_{1},\ldots,A_{m}\in\mathcal{S}^{n}$ and $b\in\mathbb{R}^{n}$, a canonical SDP problem takes the following form:
$V^{\rm p}=\max_{X\in\mathcal{S}^{n}}C\bullet X\quad\text{subjectto}\quad A_{i}\bullet X=b_{i}\quad\forall i\in[m],\quad X\succeq O.$ (B.1)
The dual problem of this problem takes the following minimization form:
$V^{\rm d}=\min_{y\in\mathbb{R}^{n}}b^{\top}y\quad\text{subject to}\quad\sum_{i\in[m]}y_{i}A_{i}\succeq C.$ (B.2)
Each problem is said to be feasible when if it has a choice variable satisfying all subjective constraints, and it is strictly feasible if there is some feasible choice that strictly satisfies the matrix inequality constraint.
The following lemma summarizes well-known duality results that hold between the primal and dual problems; for reference, see, e.g., Luenberger and Ye (2021).
Lemma 7.
If both primal and dual problems are feasible, then both admit solutions and $-\infty<V^{\rm p}\leq V^{\rm d}<\infty$ holds. In particular, if a primal-dual feasible pair $(X,y)$ satisfies
$X\bullet\quantity(\sum_{i\in[m]}y_{i}A_{i}-C)=0,$ (B.3)
then $X$ attains $V^{\rm p}$ and $y$ attains $V^{\rm d}$. If, in addition, either problem is strictly feasible, then $V^{\rm p}=V^{\rm d}$ holds, and any primal-dual feasible pair $(X,y)$ satisfies (B.3) if and only if $X$ attains $V^{\rm p}$ and $y$ attains $V^{\rm d}$.
It should be clear from construction that the solution set to either problem is convex. The next lemma will be used to prove results in Section 5.
Lemma 8.
If $(X,y)$ and $(X^{\prime},y^{\prime})$ are primal-dual optimal pairs, then so is their convex combination $(\lambda X+(1-\lambda)X^{\prime},\lambda y+(1-\lambda)y^{\prime})$ for any $\lambda\in[0,1]$.
For the sake of later reference, we note the well-known conditional formula for multivariate normal distributions. The following version, which can be found in Rao (1973), does not assume the invertibility of a variance matrix.
Lemma 9.
Let $Z=(Z_{1},Z_{2})$ be a multivariate normal random vector, partitioned into $n$-dimensional $Z_{1}$ and $m$-dimensional $Z_{2}$, with
$\operatorname{E}\matrixquantity[Z_{1}\\ Z_{2}]=\matrixquantity[\mu_{1}\\ \mu_{2}]\in\mathbb{R}^{n+m}\quad\text{and}\quad\operatorname{Var}\matrixquantity[Z_{1}\\ Z_{2}]=\matrixquantity[\Sigma_{11}&\Sigma_{12}\\ \Sigma_{21}&\Sigma_{22}]\in\mathcal{S}^{n+m}_{+}.$
Then, it holds that
$\operatorname{E}\quantity[Z_{1}\mid Z_{2}]=\mu_{1}+\Sigma_{12}\Sigma_{22}^{+}\quantity(Z_{2}-\mu_{2}),$
$\operatorname{Var}\quantity[Z_{1}\mid Z_{2}]=\Sigma_{11}-\Sigma_{12}\Sigma_{22}^{+}\Sigma_{21}.$
Appendix C Omitted proofs
Proof of Lemma 1
Given any information structure $\pi$, Proposition 4 of Miyashita and Ui (2024) implies that $(\operatorname{Var}\quantity[\sigma^{\pi}],\operatorname{Cov}\quantity[\sigma^{\pi},\theta])$ is primal feasible.
Conversely, let $(X,Y)\in\mathcal{S}^{n}\times\mathcal{M}^{n,m}$ be primal feasible. Let $\pi$ be a one-dimensional Gaussian information structure such that $\operatorname{E}\quantity[t]=\bar{a}$, $\operatorname{Var}\quantity[t]=X$, and $\operatorname{Cov}\quantity[t,\theta]=Y$, which exists since $\mathbf{M}_{X,Y}\succeq O$. Then, (11) implies
$\sum_{j=1}^{n}q_{ij}\operatorname{Cov}[t_{i},t_{j}]=\sum_{k=1}^{m}r_{ik}\operatorname{Cov}[t_{i},\theta_{k}].$
Assuming $\operatorname{Var}\quantity[t_{i}]>0$ for a while, by multiplying both sides of the above by $(t_{i}-\bar{a}_{i})/\operatorname{Var}[t_{i}]$, we obtain
$\sum_{j=1}^{n}q_{ij}\quantity(\frac{\operatorname{Cov}[t_{i},t_{j}]}{\operatorname{Var}[t_{i}]}\cdot\quantity(t_{i}-\bar{a}_{i}))=\sum_{k=1}^{m}r_{ik}\quantity(\frac{\operatorname{Cov}[t_{i},\theta_{k}]}{\operatorname{Var}[t_{i}]}\cdot\quantity(t_{i}-\bar{a}_{i})),$
from which Lemma 9 yields
$\sum_{j=1}^{n}q_{ij}\quantity(\operatorname{E}[t_{j}\mid t_{i}]-\bar{a}_{j})=\sum_{k=1}^{m}r_{ik}\quantity(\operatorname{E}[\theta_{k}\mid t_{i}]-\bar{\theta}_{k}).$
By the definition of $\bar{a}$, it follows that
$\sum_{j=1}^{n}q_{ij}\operatorname{E}[t_{j}\mid t_{i}]=\sum_{k=1}^{m}r_{ik}\operatorname{E}[\theta_{k}\mid t_{i}].$ (C.1)
Note that (C.1) holds true even when $\operatorname{Var}\quantity[t_{i}]=0$ since $\operatorname{E}[t_{j}\mid t_{i}]=a_{j}$ and $\operatorname{E}[\theta_{k}\mid t_{i}]=\bar{\theta}_{k}$.
Now, let $\sigma$ be a strategy profile such that $\sigma_{i}(t_{i})=t_{i}$ for all $i\in[n]$ and $t_{i}\in\mathbb{R}$. Then, (C.1) means that $\sigma$ constitutes a BNE under $\pi$, i.e., $\sigma=\sigma^{\pi}$. Moreover, by the distribution of $t$, we have $\mathbf{M}^{\pi}=\mathbf{M}_{X,Y}$, as desired. Q.E.D.
Proof of Proposition 2
In order to apply Lemma 7, our first task is reformulating the problem (P) in the SDP canonical form as in (B.1). For $i,j\in[n+m]$, let $E_{ij}$ denote the $(n+m)\times(n+m)$ square matrix whose $(i,j)$-entry is $1$ and all others are 00. Then, for each $i\in[n]$, let $\Phi_{i}\in\mathcal{S}^{n+m}$ be given as
$\Phi_{i}=\quantity(\sum_{j\in[n]}q_{ij}\frac{E_{ij}+E_{ji}}{2})-\quantity(\sum_{k\in[m]}r_{ik}\frac{E_{i,n+k}+E_{n+k,i}}{2}).$ (C.2)
Also, for each $k,k^{\prime}\in[m]$, let $\Psi_{kk^{\prime}}\in\mathcal{S}^{n+m}$ be given as
$\Psi_{kk^{\prime}}=\frac{E_{n+k,n+k^{\prime}}+E_{n+k^{\prime},n+k}}{2}.$ (C.3)
Now, consider any symmetric block matrix $\mathbf{M}$ of the form
$\mathbf{M}=\matrixquantity[X&Y\\ Y^{\top}&Z]\quad\text{where}\quad X\in\mathcal{S}^{n},\quad Y\in\mathcal{M}^{n,m},\quad Z\in\mathcal{S}^{m}.$
Since the upper-left block $X$ is symmetric, $\Phi_{i}\bullet\mathbf{M}=0$ holds if and only if $\sum_{j\in[n]}q_{ij}x_{ij}=\sum_{k\in[m]}r_{ik}y_{ik}$. Additionally, since $Z$ is symmetric, $\Psi_{kk^{\prime}}\bullet\mathbf{M}=\operatorname{Cov}[\theta_{k},\theta_{k^{\prime}}]$ is satisfied for all $k,k^{\prime}\in[m]$ if and only if $Z=\operatorname{Var}[\theta]$. Therefore, we can rewrite the problem (P) as
$v^{\rm p}=$ $\ \max_{\mathbf{M}\in\mathcal{S}^{n+m}}\quad$ $\mathbf{V}\bullet\mathbf{M}$
subject to $\Phi_{i}\bullet\mathbf{M}=0,\quad\forall i\in[n],$
$\Psi_{kk^{\prime}}\bullet\mathbf{M}=\operatorname{Cov}[\theta_{k},\theta_{k^{\prime}}],\quad\forall k,k^{\prime}\in[m],$
$\mathbf{M}\succeq O,$
which takes the canonical form. Notice that this problem is feasible since all constraints are satisfies by letting $X=O$, $Y=O$, and $Z=\operatorname{Var}[\theta]$.
To derive the dual problem, we denote by $\lambda_{i}$ the Lagrangian multiplier for each equality constraint $\Phi_{i}\bullet\mathbf{M}=0$ and by $\gamma_{kk^{\prime}}$ the one for $\Psi_{kk^{\prime}}\bullet\mathbf{M}=0$. Then, let $\Lambda=\operatorname{Diag}(\lambda_{1},\ldots,\lambda_{n})$ and $\Gamma=[\gamma_{kk^{\prime}}]_{m\times m}$. With these notations, following (B.2), the dual problem can be formulated as follows:
$v^{\rm d}=$ $\min_{\Lambda\in\mathcal{S}^{n}_{\rm diag},\,\Gamma\in\mathcal{M}^{m\times m}}\quad$ $\operatorname{Var}[\theta]\bullet\Gamma$
subject to $\sum_{i\in[n]}\lambda_{i}\Phi_{i}+\sum_{k,k^{\prime}\in[m]}\gamma_{kk^{\prime}}\Psi_{kk^{\prime}}\succeq\mathbf{V}.$
By the definitions of $\Phi_{i}$ and $\Psi_{kk^{\prime}}$, it is not hard to see that
Since $v^{\rm d}$ depends on $\Gamma$ only through their symmetric parts, it is without loss of optimality to let $\Gamma\in\mathcal{S}^{n}$. By doing so, the above-constructed dual problem coincides with (4.1), and note that the problem is feasible since $\Lambda=aI$ and $\Gamma=C_{\Lambda}$ satisfies dual feasibility if $a>0$ is large enough.
Let us show that the dual problem is strictly feasible. To this end, let $a,b>0$ be positive numbers, $\Lambda=aI$, and $\Gamma=abI$. Then, it follows that
which is positive definite if and only if the matrix ()(*)( ∗ ) is positive definite. Moreover, ()(*)( ∗ ) is positive definite if and only if
$\frac{Q+Q^{\top}}{2}-\frac{RR^{\top}}{4b}\succ O.$
By Assumption 1, the above matrix inequality holds true if we let $b$ be sufficiently large. Furthermore, having fixed $b$ in this way, we obtain a strictly feasible dual choice by letting $a$ be sufficiently large. Therefore, by Lemma 7, it follows that $v^{\rm p}=v^{\rm d}$, and both problems are attainable. Q.E.D.
Proof of Proposition 3
Let $\Lambda\in\mathcal{S}^{n}_{\rm diag}$ be any diagonal matrix such that $\mathbf{M}_{\Lambda}\succeq O$ and $Z\bullet C_{\Lambda}=v^{\rm d}$. Then, by Lemma 7, any primal feasible $(X,Y)\in\mathcal{S}^{n}\times\mathcal{M}^{n,m}$ attains $v^{\rm p}$ if and only if $\mathbf{M}_{X,Y}\bullet\mathbf{M}_{\Lambda}=0$. Moreover, since $\mathbf{M}_{X,Y}$ and $\mathbf{M}_{\Lambda}$ are positive semidefinite, $\mathbf{M}_{X,Y}\bullet\mathbf{M}_{\Lambda}=0$ is equivalent to $\mathbf{M}_{X,Y}\mathbf{M}_{\Lambda}=O$, which can be expanded to the following system of matrix equations:
$XA_{\Lambda}=YB_{\Lambda}^{\top},\quad XB_{\Lambda}=YC_{\Lambda},\quad Y^{\top}A_{\Lambda}=ZB_{\Lambda}^{\top},\quad Y^{\top}B_{\Lambda}=ZC_{\Lambda}.$ (C.4)
In the system, the first and third equations are nothing more than (17). The next lemma shows that the rest are redundant in the presence of these two.
Lemma 10.
For any primal feasible $(X,Y)$ and dual feasible $\Lambda$, if they satisfy the first and third equations in (C.4), then they satisfy the second and forth as well.
Proof.
Since $\mathbf{M}_{\Lambda}\succeq O$, Proposition 10.25 of Bernstein (2018) implies $B_{\Lambda}=A_{\Lambda}A_{\Lambda}^{+}B_{\Lambda}$. By the definition of $C_{\Lambda}$, the first equation in (C.4) implies that
$XB_{\Lambda}-YC_{\Lambda}=\underbrace{\quantity(XA_{\Lambda}-YB_{\Lambda}^{\top})}_{=\ O}A_{\Lambda}^{+}B_{\Lambda}=O,$
which implies the second equation in (C.4). Likewise, the third equation in (C.4) yields
$Y^{\top}B_{\Lambda}-ZC_{\Lambda}=\underbrace{\quantity(Y^{\top}A_{\Lambda}-ZB_{\Lambda})}_{=\ O}A_{\Lambda}^{+}B_{\Lambda}=O,$
which implies the fourth equation in (C.4). ∎
By Lemma 10, we conclude that any primal feasible $(X,Y)$ attains $v^{\rm p}$ if and only if there is dual feasible $\Lambda$ for which (17) is satisfied. It remains to prove the equivalence among (17), (18), and (19). Let $(X,Y)$ be primal feasible. Since $\mathbf{M}_{X,Y}\succeq O$, Proposition 10.25 of Bernstein (2018) implies that
$Y=XX^{+}Y=YZ^{+}Z.$ (C.5)
Now, whenever $A_{\Lambda}X=B_{\Lambda}Y^{\top}$ is satisfied, (C.5) implies that
$A_{\Lambda}Y=A_{\Lambda}XX^{+}Y=B_{\Lambda}Y^{\top}X^{+}Y,$
from which we see that $A_{\Lambda}Y=B_{\Lambda}Z$ if and only if $B_{\Lambda}Z=B_{\Lambda}Y^{\top}X^{+}Y$. This establishes the equivalence between (17) and (18). Likewise, whenever $A_{\Lambda}Y=B_{\Lambda}Z$ is satisfied, (C.5) implies that
$B_{\Lambda}Y=B_{\Lambda}ZZ^{+}Y^{\top}=A_{\Lambda}YZ^{+}Y^{\top},$
from which we see that $B_{\Lambda}Y=A_{\Lambda}X$ if and only if $A_{\Lambda}X=A_{\Lambda}YZ^{+}Y^{\top}$. This establishes the equivalence between (17) and (19). Q.E.D.
Proof of Corollary 4
The first assertion is discussed in the main text. Moreover, the equivalence holds because $\mathbf{V}\bullet\mathbf{M}_{X,Y}=Z\bullet C_{\Lambda}$ if and only if $\mathbf{M}_{X,Y}\bullet\mathbf{M}_{\Lambda}=0$ (cf. Lemma 7), and because $\mathbf{M}_{X,Y}\bullet\mathbf{M}_{\Lambda}=0$ if and only if (17) (cf. Lemma 10 and the discussion preceding to it). Q.E.D.
Proof of Corollary 5
Let $\pi\in\Pi^{\rm g}$ be any optimal information structure with $X=\operatorname{Var}\quantity[\sigma^{\pi}]$ and $Y=\operatorname{Cov}\quantity[\sigma^{\pi},\theta]$ that is certified by $\Lambda\in\mathcal{S}^{n}_{\rm diag}$. Note that $(X,Y)$ and $\Lambda$ jointly satisfy (C.4). Then, Proposition 2 implies $v^{\rm p}=C_{\Lambda}\bullet Z$, while (C.4) implies $C_{\Lambda}\bullet Z=B_{\Lambda}\bullet Y=A_{\Lambda}\bullet X$. Moreover, by using Lemma 9, it follows that
[…]
where the last equality holds by $\mathbf{M}_{X,Y}\succeq O$. Similarly, by using the definition of $C_{\Lambda}$, it follows that
[…]
where the last equality holds by $\mathbf{M}_{X,Y}\succeq O$ and $\mathbf{M}_{\Lambda}\succeq O$. Q.E.D.
Proof of Lemma 2
Assuming $Z\succ O$, (17) implies that $B_{\Lambda}=A_{\Lambda}YZ^{-1}$ and $C_{\Lambda}=Z^{-1}Y^{\top}A_{\Lambda}YZ^{-1}$. Hence, it follows that
$\mathbf{M}_{\Lambda}=\matrixquantity[A_{\Lambda}&-A_{\Lambda}YZ^{-1}\\ -Z^{-1}Y^{\top}A_{\Lambda}&Z^{-1}Y^{\top}A_{\Lambda}YZ^{-1}]=\matrixquantity[I&O\\ O&YZ^{-1}]^{\top}\matrixquantity[A_{\Lambda}&-A_{\Lambda}\\ -A_{\Lambda}&A_{\Lambda}]\matrixquantity[I&O\\ O&YZ^{-1}],$
from which we see that $\mathbf{M}_{\Lambda,\Gamma}\succeq O$ if and only if $A_{\Lambda}\succeq O$. Q.E.D.
Proof of Corollary 6
Let $\pi\in\Pi^{\rm g}$ be any optimal information structure with $X=\operatorname{Var}\quantity[\sigma^{\pi}]$ and $Y=\operatorname{Cov}\quantity[\sigma^{\pi},\theta]$ that is certified by $\Lambda\in\mathcal{S}^{n}_{\rm diag}$. Then, Lemma 9 and (18) imply that
$\operatorname{E}\quantity[A_{\Lambda}\quantity(\sigma^{\pi}-\bar{a})\mid\theta]=A_{\Lambda}YZ^{+}\quantity(\theta-\bar{\theta})=B_{\Lambda}ZZ^{+}\quantity(\theta-\bar{\theta}),$
$\operatorname{Var}\quantity[A_{\Lambda}\quantity(\sigma^{\pi}-\bar{a})\mid\theta]=A_{\Lambda}\quantity(X-Y^{\top}Z^{+}Y)A_{\Lambda}=O.$
This implies that $A_{\Lambda}\quantity(\sigma^{\pi}-\bar{a})=B_{\Lambda}ZZ^{+}\quantity(\theta-\bar{\theta})$ a.s. In particular, if $Z\succ O$, then we have $A_{\Lambda}\quantity(\sigma^{\pi}-\bar{a})=B_{\Lambda}\quantity(\theta-\bar{\theta})$ a.s.
Conversely, suppose that $\pi$ and dual feasible $\Lambda$ jointly satisfy (20). Multiplying both sides of (20) by $\quantity(\sigma^{\pi}-\bar{a})^{\top}$ from right and taking unconditional expectations, it follows that
$A_{\Lambda}\quantity(\sigma^{\pi}-\bar{a})\quantity(\sigma^{\pi}-\bar{a})^{\top}=B_{\Lambda}\quantity(\theta-\bar{\theta})\quantity(\sigma^{\pi}-\bar{a})^{\top}\quad\Longrightarrow\quad A_{\Lambda}\operatorname{Var}\quantity[\sigma^{\pi}]=B_{\Lambda}\operatorname{Cov}\quantity[\theta,\sigma^{\pi}].$
Analogously, multiplying both sides of (20) by $\quantity(\theta-\bar{\theta})^{\top}$ from right and taking unconditional expectations, we get
$A_{\Lambda}\quantity(\sigma^{\pi}-\bar{a})\quantity(\theta-\bar{\theta})^{\top}=B_{\Lambda}\quantity(\theta-\bar{\theta})\quantity(\theta-\bar{\theta})^{\top}\quad\Longrightarrow\quad A_{\Lambda}\operatorname{Cov}\quantity[\sigma^{\pi},\theta]=B_{\Lambda}Z.$
Hence, $\pi$ and $\Lambda$ satisfy (17), from which Proposition 3 concludes that $\pi$ is optimal. Q.E.D.
Proof of Corollary 7
It has already been discussed in the main text that $\underline{\pi}$ is optimal if and only if it is certified by $\Lambda$ such that $\mathbf{M}_{\Lambda}\succeq O$ and $B_{\Lambda}=O$, while Lemma 2 implies $\mathbf{M}_{\Lambda}\succeq O$ is equivalent to $A_{\Lambda}\succeq O$ when $Z\succ O$. For the optimality of $\overline{\pi}$, taking $X=Q^{-1}RZ(Q^{-1}R)$ and $Y=Q^{-1}RZ$, the first equality in (19) is automatically satisfied irrespective of $\Lambda$, and the second equality becomes $A_{\Lambda}Q^{-1}R=B_{\Lambda}Z$. Q.E.D.
Proof of Corollary 8
The first assertion is confirmed by applying the optimality condition (19) in Proposition 3. If there exists a dual optimal $\Lambda$ such that $A_{\Lambda}\succ O$, then Corollary 6 implies that any $\pi$ is noise free. Moreover, in that case, any primal solutions $(X,Y)$ and $(\tilde{X},\tilde{Y})$ have to satisfy
$A_{\Lambda}X=B_{\Lambda}Y^{\top},\quad A_{\Lambda}Y=B_{\Lambda}Z,\quad A_{\Lambda}\tilde{X}=B_{\Lambda}\tilde{Y}^{\top},\quad A_{\Lambda}\tilde{Y}=B_{\Lambda}Z.$
By the second and forth equations, we have $A_{\Lambda}(Y-\tilde{Y})=O$, from which $Y=\tilde{Y}$ since $A_{\Lambda}$ is non-singular, Then, by the first and third equations, we have $A_{\Lambda}(X-\tilde{X})=O$, from which $X=\tilde{X}$ for the same reason. Q.E.D.
Proof of Corollary 9
[…]
Proof of Corollary 10
Suppose $m=1$. As for the optimality of no disclosure $\underline{\pi}$, notice that $B_{\Lambda}=O$ simplifies to $\lambda_{i}r_{i}+w_{i}=0$ for each $i\in[n]$, thereby dictating $\Lambda=-D$, where $D=\operatorname{Diag}\quantity(w_{1}/r_{1},\ldots,w_{n}/r_{n})$. Hence, by Corollary 7, it follows that $\underline{\pi}$ is optimal if and only if $A_{-D}\succeq O$, or equivalently, $O\succeq V+(DQ+Q^{\top}D)/2$.
Now, suppose that $O\not\succeq V+(DQ+Q^{\top}D)/2$. By the previous conclusion, it follows that $\underline{\pi}$ is suboptimal so that $v^{\rm p}>0$. By Corollary 5, we need to have $B_{\Lambda}\neq O$ whenever $\Lambda$ certifies an optimal information structure. Since this implies $B_{\Lambda}$ has full column rank, Corollary 9 implies that any optimal information structure is aggregatively disclosing. Q.E.D.
Proof of Lemma 3
Suppose that $W=\Delta R$ for some $\Delta\in\mathcal{S}^{n}_{\rm diag}$. Fix any $(X,Y)$ that is feasible under $(Q,R)$. Then, it follows that
[…]
Moreover, since $X$ is symmetric, we have $\Delta Q\bullet X=Q^{\top}\Delta\bullet X$, from which $W\bullet Y=(\Delta Q+Q^{\top}\Delta)/2\bullet X$. Hence $V\bullet X+W\bullet Y=V_{\Delta}\bullet X$. Q.E.D.
Proof of Corollary 11
Let $\mathscr{P}$ be a payoff structure satisfying all assumptions of the corollary, and let $\tilde{\mathscr{P}}=(Q,R,V_{\Delta},O)$ be the state-independent form of $\mathscr{P}$.
(iii) $\Longrightarrow$ (i). As discussed in the main text, $\overline{\pi}$ is optimal under a team payoff structure.
(i) $\Longrightarrow$ (ii). Since $\overline{\pi}$ is optimal under $\tilde{\mathscr{P}}$ as well, Corollary 7 implies there exists $\Lambda\in\mathcal{S}^{n}_{\rm diag}$ such that
$\underbrace{\quantity(\frac{\Lambda Q+Q^{\top}\Lambda}{2}-V_{\Delta})}_{\succeq\ O}Q^{-1}R=\frac{\Lambda R}{2},$
which can be rewritten as
$\quantity(2Q^{\top}\Lambda-V_{\Delta})Q^{-1}R=O.$
Since $R$ has full rank by assumption, $R$ has a right inverse, from which we have $Q^{\top}\Lambda=2V_{\Delta}$. Moreover, since $V_{\Delta}$ is symmetric, it follows that $\Lambda Q=2V_{\Delta}$. Thus we have $V_{\Delta}Q^{-1}=\Lambda/2$, which implies that $V_{\Delta}Q^{-1}$ is a diagonal matrix. In addition, since $V_{\Delta}=(\Lambda Q+Q^{\top}\Lambda)/2-V_{\Delta}$, we have $V_{\Delta}\succeq O$.
It remains to show that $V_{\Delta}Q^{-1}\succ O$ and $V_{\Delta}\succ O$. By $\Lambda Q=2V_{\Delta}$, we can see that the $i$-th row of $V_{\Delta}$ is given by $(\lambda_{i}q_{i1}/2,\ldots,\lambda_{i}q_{in}/2)$, where $\lambda_{i}$ is the $i$-th diagonal entry of $\Lambda$. Since all row vectors of $V_{\Delta}$ are non-zero by assumption, it follows that $\lambda_{i}\neq 0$ for all $i$. Moreover, since $V_{\Delta}\succeq O$ and $Q+Q^{\top}\succ O$, we have $\lambda_{i}q_{ii}/2\geq 0$ and $q_{ii}>0$, whence $\lambda_{i}>0$ for all $i$. To show $V\succ O$, seeking a contradiction, assume that $V$ has 00 as its eigenvalue so that $V{\mathbf{x}}=0$ holds for some non-zero ${\mathbf{x}}\in\mathbb{R}^{n}$. By $V=Q^{\top}\Lambda/2$, this implies that $Q^{\top}{\mathbf{y}}=0$ with ${\mathbf{y}}=\Lambda{\mathbf{x}}/2$. Since ${\mathbf{x}}\neq 0$, and since $\lambda_{i}\neq 0$ for all $i$, we have ${\mathbf{y}}\neq 0$. Thus, $Q^{\top}{\mathbf{y}}$ implies that 00 is an eigenvalue of $Q^{\top}$, contradicting the non-singularity of $Q^{\top}$.
(ii) $\Longrightarrow$ (iii). Suppose that $V_{\Delta}\succ O$ and $V_{\Delta}Q^{-1}\in\mathcal{S}^{n}_{\rm diag}\cap\mathcal{S}^{n}_{++}$. As discussed in the main text, by multiplying $Q$ and $R$ by $V_{\Delta}Q^{-1}$ from right, the original payoff structure $\mathscr{P}$ is equivalently transformed to $(V_{\Delta},V_{\Delta}Q^{-1}R,V_{\Delta},O)$. Since $V_{\Delta}\succ O$, this shows that $\mathscr{P}$ is a team payoff structure. Q.E.D.
Proof of Corollary 12
Let $\mathscr{P}$ satisfy PDC with respect to $\Delta$. By Corollary 7, no information disclosure $\underline{\pi}$ is optimal under $\mathscr{P}$ if and only if it is optimal under the equivalent payoff structure $(Q,R,V_{\Delta},O)$, which occurs if and only if there exists $\Lambda\in\mathcal{S}^{n}_{\rm diag}$ such that
$\frac{\Lambda Q+Q^{\top}\Lambda}{2}\succeq V_{\Delta}\quad\text{and}\quad\Lambda R=O.$
If $O\succeq V_{\Delta}$, these conditions can be satisfied with $\Lambda=O$. Q.E.D.
Proof of Proposition 13
Let $(X,Y)$ be a solution to (P), and let $(\Lambda,\Gamma)$ be a solution to (D), which exist by Proposition 2.
For any $\tau\in\operatorname{Aut}(G)$, $(\tau(X),\tau(Y))$ satisfies the constraints of (P) and attains the same value; that is, $(\tau(X),\tau(Y))$ is also a solution to (P). To see this, observe first that $(\tau(X),\tau(Y))$ satisfies the constraints
$\matrixquantity[\tau(X)&\tau(Y)\\ \tau(Y)^{\top}&Z]=\matrixquantity[\tau(X)&\tau(Y)\\ \tau(Y)^{\top}&\tau(Z)]\succeq O$
and $\operatorname{diag}(\tau(X)Q^{\top})=\operatorname{diag}(\tau(X)\tau(Q)^{\top})=\operatorname{diag}(\tau(XQ^{\top}))=\operatorname{diag}(\tau(Y))$. In addition, $(\tau(X),\tau(Y))$ attains the same value since
[…]
and similarly, $W\bullet Y=W\bullet\tau(Y)$. Consider the average of $\tau(X)$ and $\tau(Y)$ over $\operatorname{Aut}(G)$:
$X^{*}\equiv\sum_{\tau\in\operatorname{Aut}(G)}\tau(X)/|\operatorname{Aut}(G)|$ and $Y^{*}\equiv\sum_{\tau\in\operatorname{Aut}(G)}\tau(\tilde{Y})/|\operatorname{Aut}(G)|$.
Then, $(X^{*},Y^{*})$ is a solution to (P) because the set of solutions is a convex set. By construction, $X^{*}$ and $Y^{*}$ are $G$-invariant, which implies the existence of a $G$-invariant optimal Gaussian information structure $\pi$.
Similarly, one can show that $(\tau(\Lambda),\tau(\Gamma))$ is a solution to (D), and thus $(\Lambda^{*},\Gamma^{*})$ given by $\Lambda^{*}=\sum_{\tau\in\operatorname{Aut}(G)}\tau(\Lambda)/|\operatorname{Aut}(G)|$ and $\Gamma^{*}=\sum_{\tau\in\operatorname{Aut}(G)}\tau(\Gamma)/|\operatorname{Aut}(G)|$ is also a solution. Thus, the $G$-invariant diagonal matrix $\Lambda^{*}$ certifies $\pi$.
We prove (i) and (ii). Since $\operatorname{Aut}(G)$ is transitive, we can write $\Lambda^{*}=\lambda_{0}^{*}I$, and thus $A_{\Lambda^{*}}=\lambda_{0}^{*}Q_{G}-V$ and $B_{\Lambda^{*}}=(\lambda_{0}^{*}I_{n}+W)/2$. Under the assumption of (i), we show that $A_{\Lambda^{*}}$ is non-singular. Seeking a contradiction, suppose $A_{\Lambda^{*}}=\lambda_{0}^{*}Q_{G}-V$ is singular. Then, $Y^{*\top}A_{\Lambda^{*}}$ is singular, so $B_{\Lambda^{*}}=(\lambda_{0}^{*}I_{n}+W)/2$ is singular by (17), which contradicts the assumption.
Under the assumption of (ii), if $V_{\Delta}$ is negative semidefinite, then no information disclosure is optimal by Corollary 12, and it is noise free. Suppose $V_{\Delta}$ is not negative semidefinite. If $\lambda_{0}^{*}=0$, then $A_{\Lambda^{*}}=-V_{\Delta}\succeq 0$ since $\Lambda^{*}$ certifies $\pi$, which contradicts to the assumption. Thus, $B_{\Lambda^{*}}=\lambda_{0}^{*}I_{n}/2$ is non-singular, which implies aggregative disclosing and thus noise-freeness in the case of $m=n$. Q.E.D.
Proof of Proposition 14
By Proposition 13, there exists a $G$-invariant information structure $\pi$ in the original problem with an optimal feasible pair $(X,Y)$, and there exists a $G$-invariant diagonal matrix $\Lambda^{*}$ that certifies $\pi$. By Lemma 4, there exists an optimal information structure $\tilde{\pi}$ in the state reduction form of the problem with an optimal feasible pair $(X,\tilde{Y})$ such that $X$ and $\tilde{Y}\tilde{R}$ are $G$-invariant, where $Y=\tilde{Y}\tilde{R}^{\top}$.
We show that $\Lambda^{*}$ certifies $\tilde{\pi}$. Let $\Gamma^{*}\in\mathcal{S}^{n}$ be such that $(\Lambda^{*},\Gamma^{*})$ is a solution to the original dual problem. Define $\tilde{\Gamma}^{*}\equiv\tilde{R}^{\top}\Gamma^{*}\tilde{R}\in\mathcal{S}^{m}$. In the sequel, we show that $(\Lambda^{*},\tilde{\Gamma}^{*})$ is a solution to the dual problem of the reduced form, which implies that $\Lambda^{*}$ certifies $\tilde{\pi}$.
The original dual problem is
and the reduced dual problem is
Thus, $(\Lambda^{*},\Gamma^{*})$ is feasible in the reduced dual problem since
[…]
and
$\matrixquantity[A_{\Lambda^{*}}&-B_{\Lambda^{*}}\tilde{R}\\ -(B_{\Lambda^{*}}\tilde{R})^{\top}&\tilde{\Gamma}^{*}]=\matrixquantity[I_{n}&O\\ O&\tilde{R}]^{\top}\matrixquantity[A_{\Lambda^{*}}&-B_{\Lambda^{*}}\\ -B_{\Lambda^{*}}^{\top}&\Gamma^{*}]\matrixquantity[I_{n}&O\\ O&\tilde{R}]\succeq O,$
which implies $\tilde{v}^{\rm d}\leq v^{\rm d}$. On the other hand, $v^{\rm d}$ equals the value of the original primal problem $v^{\rm p}$, which equals the value of the reduced primal problem $\tilde{v}^{\rm p}$ by Lemma 4; that is,
$\tilde{v}^{\rm p}=\tilde{v}^{\rm d}\leq v^{\rm d}=v^{\rm p}=\tilde{v}^{\rm p}.$
Therefore, $(\Lambda^{*},\Gamma^{*})$ is a solution to the reduced dual problem. Q.E.D.
Proof of Lemma 4
To prove the first statement, suppose that $(X,Y)$ is feasible under $(Q_{G},I_{n},Z)$, where $Z=\tilde{R}\tilde{R}^{\top}$:
$\operatorname{diag}(XQ_{G}^{\top})=\operatorname{diag}(Y)\quad\text{and}\quad\matrixquantity[X&Y\\ Y^{\top}&Z]\succeq O.$ (C.6)
The latter condition implies the existence of $\tilde{M}\in\mathcal{M}^{n,n}$ such that $Y^{\top}=Z\tilde{M}=\tilde{R}\tilde{R}^{\top}\tilde{M}$. Let $\tilde{Y}=\tilde{M}^{\top}\tilde{R}$. Then, $Y=\tilde{Y}\tilde{R}^{\top}$, and (C.6) implies
$\operatorname{diag}(XQ_{G}^{\top})=\operatorname{diag}(\tilde{Y}\tilde{R}^{\top})\quad\text{and}\quad\matrixquantity[X&\tilde{Y}\tilde{R}^{\top}\\ (\tilde{Y}\tilde{R}^{\top})^{\top}&\tilde{R}\tilde{R}^{\top}]\succeq O,$ (C.7)
and thus
$\operatorname{diag}(XQ_{G}^{\top})=\operatorname{diag}(\tilde{Y}\tilde{R}^{\top})\quad\text{and}\quad\matrixquantity[X&\tilde{Y}\\ \tilde{Y}^{\top}&I_{m}]\succeq O.$ (C.8)
Therefore, $(X,\tilde{Y})$ is feasible in $(Q_{G},\tilde{R},I_{m})$.
Conversely, suppose that $(X,\tilde{Y})$ is feasible under $(Q_{G},\tilde{R},I_{m})$, i.e., (C.8) holds. Then, (C.7) holds as well, which implies (C.6). Q.E.D.
Proof of Proposition 14
By Proposition 13, there exists a $G$-invariant information structure $\pi$ in the original problem with an optimal feasible pair $(X,Y)$, and there exists a $G$-invariant $\Lambda$ that certifies $\pi$. By Lemma 4, there exists an optimal information structure $\tilde{\pi}$ in the state reduction form of the problem with an optimal feasible pair $(X,\tilde{Y})$ such that $X$ and $\tilde{Y}\tilde{R}$ are $G$-invariant, where $Y=\tilde{Y}\tilde{R}^{\top}$.
We show that $\Lambda$ certify $\tilde{\pi}$. Let $\Gamma\in\mathcal{S}^{n}$ be such that $(\Lambda,\Gamma)$ is a solution to the original dual problem. Define $\tilde{\Gamma}\equiv\tilde{R}^{\top}\Gamma\tilde{R}\in\mathcal{S}^{m}$. Then, it is straightforward to show that $(\Lambda,\tilde{\Gamma})$ is a solution to the dual problem of the reduced form, which implies that $\Lambda$ certify $\tilde{\pi}$.
Proof of Lemma 5
As discussed in the main text, a unique BNE is characterized as $\sigma^{\pi}=Q^{-1}R\operatorname{E}\quantity[\theta\mid t_{0}]$ for an arbitrary public information structure $\pi\in\Pi^{\rm pb}$. From this, we can readily confirm that $\mathbf{M}^{\pi}=\mathbf{M}_{\operatorname{Var}\quantity[\operatorname{E}\quantity[\theta\mid t_{0}]]}$. Moreover, $Z\succeq\operatorname{Var}\quantity[\operatorname{E}\quantity[\theta\mid t_{0}]]\succeq O$ holds by the law of total variance.
Conversely, take any $S\in\mathcal{S}^{m}$ such that $Z\succeq S\succeq O$. Then, we let $\pi\in\Pi^{\rm pb}\cap\Pi^{\rm g}$ be a public Gaussian information structure, sending an $m$-dimensional public signal $t_{0}$, such that $t_{0}$ and $\theta$ are jointly normally distributed with $\operatorname{Var}\quantity[t_{0}]=\operatorname{Cov}\quantity[t_{0},\theta]=S$. Such a random vector exists because we have
as confirmed by the positive semidefiniteness of the Schur complement $Z-SS^{+}S=Z-S$. Using the conditional Gaussian formula, the unique BNE under $\pi$ is computed as
$\sigma^{\pi}\quantity(t_{0})=Q^{-1}R\operatorname{E}\quantity[\theta\mid t_{0}]=Q^{-1}RSS^{+}t_{0}-\underbrace{Q^{-1}R\quantity(\bar{\theta}+SS^{+}\operatorname{E}\quantity[t_{0}])}_{\rm constant}.$
Hence, it follows that
$\operatorname{Var}\quantity[\sigma^{\pi}(t_{0})]=Q^{-1}RSS^{+}\operatorname{Var}\quantity[t_{0}]S^{+}S\quantity(Q^{-1}R)^{\top}=Q^{-1}RS\quantity(Q^{-1}R)^{\top},$
$\operatorname{Cov}\quantity[\sigma^{\pi}(t_{0}),\theta]=Q^{-1}RSS^{+}\operatorname{Cov}\quantity[t_{0},\theta]=Q^{-1}RSS^{+}.$
Thus we have $\mathbf{M}^{\pi}=\mathbf{M}_{S}$, as desired. Q.E.D.
Proof of Lemma 6
Given any $\pi\in\Pi^{\rm pb}$, let $S=\operatorname{Var}\quantity[\operatorname{E}\quantity[\theta\mid t_{0}]]$. Substituting $\mathbf{M}^{\pi}=\mathbf{M}_{S}$ into (14), we perform some elementary calculations to obtain
$\operatorname{E}\quantity[v(\sigma^{\pi},\theta)]$ $=\mathbf{V}\bullet\mathbf{M}_{S}$
$=V\bullet Q^{-1}RS(Q^{-1}R)^{\top}+\frac{1}{2}W\bullet Q^{-1}RS+\frac{1}{2}W^{\top}\bullet S(Q^{-1}R)^{\top}$
[…]
$=\overline{C}\bullet S,$
as desired. Q.E.D.
Proof of Proposition 15
Consider the eigendecomposition $\widehat{C}=UDU^{\top}$, as introduced in the main text, where we denote $D=\operatorname{Diag}(\gamma_{1},\ldots,\gamma_{m})$. We first show that the sum of positive eigenvalues of $\widehat{C}$ serves as an upper bound on the value of any feasible choice.
Lemma 11.
For any $S\in\mathcal{S}^{m}$, if $I\succeq S\succeq O$, then all diagonal entries of $U^{\top}SU$ are confined in $[0,1]$. Consequently,
$\widehat{C}\bullet S\leq\sum_{k=1}^{k^{*}}\gamma_{k}.$
Proof.
Denote by $s_{1},\ldots,s_{m}$ the diagonal entries of $U^{\top}SU$. Since $S\succeq O$, we have $U^{\top}SU\succeq O$, thereby all of $s_{1},\ldots,s_{m}$ being nonnegative. Moreover, since $I\succeq S$, and since $U$ is orthogonal, it follows that $I-U^{\top}SU=U^{\top}(I-S)U\succeq O$. This implies that all diagonal entries of $I-U^{\top}SU$ are nonnegative, which in turn implies that all of $s_{1},\ldots,s_{m}$ are weakly less that $1$. Hence, it follows that
[…]
where the inequality holds since $\gamma_{k}>0$ if and only if $k\leq k^{*}$ and $s_{k}\in[0,1]$ for all $k$. ∎
Let us show that the upper bound of Lemma 11 is achieved by public signal $t_{0}=U_{k^{*}}^{\top}Z^{-1/2}\theta$. By construction, we have $\operatorname{Var}\quantity[t_{0}]=U_{k^{*}}^{\top}U_{k^{*}}$ and $\operatorname{Cov}\quantity[t_{0},\theta]=U_{k^{*}}^{\top}Z^{1/2}$. Then, the conditional Gaussian formula yields
$\operatorname{Var}\quantity[\operatorname{E}\quantity[\theta\mid t_{0}]]$ $=Z^{\frac{1}{2}}U_{k^{*}}\quantity(U_{k^{*}}^{\top}U_{k^{*}})^{-1}U_{k^{*}}^{\top}Z^{\frac{1}{2}}=Z^{\frac{1}{2}}U_{k^{*}}U_{k^{*}}^{\top}Z^{\frac{1}{2}},$
where the second equality holds since the column vectors of $U_{k^{*}}$ constitutes an orthonormal system, thereby $U_{k^{*}}^{\top}U_{k^{*}}$ being an identify matrix. Hence, by Lemma 6, the designer’s expected value can be calculated as follows:
[…]
We note that the last equality holds because $U^{\top}U_{k^{*}}=[\delta_{ij}]_{m\times k^{*}}$, where $\delta_{ij}$ denotes the Kronecker delta.
Lastly, we shall show that $p(\mathbf{V})-n\leq p(\widehat{C})\leq p(\mathbf{V})$. Since $\widehat{C}$ and $\overline{C}$ are congruent, it follows that $p(\widehat{C})=p(\overline{C})$ by Sylvester’s law of inertia. Moreover, by arranging (15), we can express $\overline{C}$ as
$\overline{C}=\mathbf{Q}^{\top}\mathbf{V}\mathbf{Q}\quad{\rm where}\quad\mathbf{Q}\equiv\matrixquantity[Q^{-1}R\\ I]\in\mathcal{M}^{n+m,m}.$
[…]
Now, suppose that the $k$-th largest eigenvalue of $\overline{C}$ is strictly positive. By the Courant–Fischer theorem, it follows that there exist $\epsilon>0$ and a $k$-dimensional linear subspace $\mathcal{U}\subseteq\mathbb{R}^{m}$ such that
$\min_{\mathbf{u}\in\mathcal{U}\setminus\{\mathbf{0}\}}\frac{\mathbf{u}^{\top}\overline{C}\mathbf{u}}{\mathbf{u}^{\top}\mathbf{u}}\geq\epsilon.$
Since $\ker\quantity(\mathbf{Q})=\{\mathbf{0}\}$, it follows that
$\frac{\quantity(\mathbf{Q}\mathbf{u})^{\top}\mathbf{V}\quantity(\mathbf{Q}\mathbf{u})}{\quantity(\mathbf{Q}\mathbf{u})^{\top}\quantity(\mathbf{Q}\mathbf{u})}\cdot\frac{\quantity(\mathbf{Q}\mathbf{u})^{\top}\quantity(\mathbf{Q}\mathbf{u})}{\mathbf{u}^{\top}\mathbf{u}}\geq\epsilon,\quad\forall\mathbf{u}\in\mathcal{U}\setminus\{\mathbf{0}\},$
from which we have
$\min_{\mathbf{u}\in\mathcal{U}\setminus\{\mathbf{0}\}}\frac{\quantity(\mathbf{Q}\mathbf{u})^{\top}\mathbf{V}\quantity(\mathbf{Q}\mathbf{u})}{\quantity(\mathbf{Q}\mathbf{u})^{\top}\quantity(\mathbf{Q}\mathbf{u})}\geq\frac{\epsilon}{\delta}\quad{\rm where}\quad\delta\equiv\max\quantity{\frac{\quantity(\mathbf{Qu})^{\top}\mathbf{Qu}}{\mathbf{u}^{\top}\mathbf{u}}:\mathbf{u}\in\mathbb{R}^{m}\setminus\{\mathbf{0}\}}.$
Notice that $\delta>0$ since $\mathbf{Q}^{\top}\mathbf{Q}$ is a positive definite matrix. Then, since $\{\mathbf{Qu}\in\mathbb{R}^{n+m}:\mathbf{u}\in\mathcal{U}\}$ is a $k$-dimensional linear subspace of $\mathbb{R}^{n+m}$, again by the Courant–Fischer theorem, the above equation suggests that the $k$-th largest eigenvalue of $\mathbf{V}$ is strictly positive. Thus we have $p(\overline{C})\leq p(\mathbf{V})$.
If we denote by $p^{\prime}(\cdot)$ the number of non-negative eigenvalues of a given square matrix, we can also show that $p^{\prime}(\overline{C})\leq p^{\prime}(\mathbf{V})$ by slightly changing the preceding proof. Moreover, applying the same result to $-\overline{C}$ and $-\mathbf{V}$, it follows that $p^{\prime}(-\overline{C})\leq p^{\prime}(-\mathbf{V})$. Then, noticing that $p^{\prime}(-\overline{C})=m-p(\overline{C})$ and $p^{\prime}(-\mathbf{V})=(n+m)-p(\mathbf{V})$, we obtain
$m-p(\overline{C})\leq(n+m)-p(\mathbf{V}),$
from which $p(\mathbf{V})-n\leq p(\overline{C})$. Q.E.D.
Proof of Corollary 16
Let $\pi$ be the optimal public information structure as in Proposition 15, and observe that $\pi$ is noise free with
$\operatorname{Cov}\quantity[\sigma^{\pi},\theta]=Q^{-1}RZ^{\frac{1}{2}}U_{k^{*}}U_{k^{*}}^{\top}Z^{\frac{1}{2}}.$
Hence, by Corollary 8, we see that $\pi$ is optimal in $\Pi^{*}$ if and only if there exists $\Lambda\in\mathcal{S}^{n}_{\rm diag}$ with $A_{\Lambda}\succeq O$ such that
$A_{\Lambda}Q^{-1}RZ^{\frac{1}{2}}U_{k^{*}}U_{k^{*}}^{\top}Z^{\frac{1}{2}}=B_{\Lambda}Z.$
By construction, we have $U_{k^{*}}U_{k^{*}}^{\top}=\sum_{k=1}^{k^{*}}\mathbf{u}_{k}\mathbf{u}_{k}^{\top}$ and $I=UU^{\top}=\sum_{k=1}^{m}\mathbf{u}_{k}\mathbf{u}_{k}^{\top}$. Hence, the above condition can be equivalently rewritten as follows:
$A_{\Lambda}Q^{-1}RZ^{\frac{1}{2}}U_{k^{*}}U_{k^{*}}^{\top}=B_{\Lambda}Z^{\frac{1}{2}}$
$\quad\iff\quad A_{\Lambda}Q^{-1}RZ^{\frac{1}{2}}\quantity(\sum_{k=1}^{k^{*}}\mathbf{u}_{k}\mathbf{u}_{k}^{\top})=B_{\Lambda}Z^{\frac{1}{2}}\quantity(\sum_{k=1}^{m}\mathbf{u}_{k}\mathbf{u}_{k}^{\top})$
$\quad\iff\quad\quantity(A_{\Lambda}Q^{-1}RZ^{\frac{1}{2}}-B_{\Lambda})Z^{\frac{1}{2}}\sum_{k=1}^{k^{*}}\mathbf{u}_{k}\mathbf{u}_{k}^{\top}-B_{\Lambda}Z^{\frac{1}{2}}\sum_{k=k^{*}+1}^{m}\mathbf{u}_{k}\mathbf{u}_{k}^{\top}=O.$ (C.9)
It is evident that (C.9) holds if both (33) and (34) are satisfied. Conversely, suppose that (C.9) is satisfied. Then, since $\{\mathbf{u}_{1},\ldots,\mathbf{u}_{m}\}$ constitutes an orthonormal basis in $\mathbb{R}^{m}$, multiplying (C.9) by $\mathbf{u}_{k}$ from right, we obtain (33) when $k\leq k^{*}$, and (34) when $k>k^{*}$. This implies that (C.9) is equivalent to (33) and (34), which concludes the proof. Q.E.D.