### v3

Abstract
This paper addresses information design in a workhorse model of network games, where agents have linear best responses, the information designer optimizes a quadratic objective, and the payoff state follows a multivariate Gaussian distribution. We formulate the problem as semidefinite programming (SDP) and utilize the duality principle to characterize an optimal information structure. A Gaussian information structure is shown to be optimal among all information structures. A necessary and sufficient condition for optimality is that the induced equilibrium strategy profile and the state jointly satisfy a linear constraint derived from complementary slackness conditions. Consequently, the true state is typically revealed to the entire population of agents, even though individual agents remain only partially informed. In symmetric network games, an optimal information structure inherits the same degree of symmetry, which facilitates its computation.
JEL classification: C72, D82
Keywords: linear-quadratic-Gaussian, information design, network interventions, semidefinite programming, strong duality
1 Introduction
We study games among agents who strategically interact with one another under incomplete information. The optimal action of each agent—e.g., a level of investment—is dependent on other agents’ action choices through payoff externalities, as well as on exogenous payoff-relevant states, such as random returns on investment. An extra player of the game, called an information designer, chooses an information structure—i.e., a joint distribution of payoff states and agents’ private signals—to manipulate agents’ incentives. Our aim is to understand how the information designer can optimally design such information structures for her objective maximization.
Information design with multiple agents is fundamentally different from the single-agent case because agents face not only uncertainty about economic fundamentals but also strategic uncertainty about the behavior of others. In such environments, providing information to some agents can generate “spillover effects” for others: When agent $i$ receives information, her neighbor $j$ may adjust their strategy in response to the anticipated change in $i$’s behavior. Moreover, these effects can cascade through the network, as agent $i$ might further update her strategy in response to the anticipated change in $j$’s behavior, and so forth. These interdependent incentives make characterizing the optimal information structure a challenging task.
To bring insights into the relationship between optimal information structures and economic environments, this paper addresses a tractable class of information design problems that are neither too specific nor too general. To this end, we focus on linear-quadratic Gaussian (LQG) games, where agents’ payoff functions and designer’s objective function are quadratic in continuous actions and payoff states, which follow a multivariate Gaussian distribution. Since payoff functions are arbitrary as long as they are quadratic, the class can encapsulate a wide array of strategic situations. [^3] The origin of LQG games dates back to the seminal work of Radner (1962). Several instances of LQG games include Cournot and Bertrand oligopoly (Vives, 1984, 1999), financial markets and beauty contests (Morris and Shin, 2002, Angeletos and Pavan, 2007), public goods provision (Bramoullé and Kranton, 2007), and network games (Ballester et al., 2006, Bramoullé et al., 2014) among other applications. In addition, when payoffs are quadratic, an objective can typically be quadratic, as when a utilitarian planner aims to maximize the sum of agents’ payoffs. The information designer has the ability to select and commit to an information structure from a feasible set, fixing the payoff functions and the prior distribution of states. LQG information design is a problem of finding an optimal information structure that maximizes the expected value of the objective function.
No a priori restrictions are imposed on the distribution of private signals. However, our foundational observation is that among all information structures, there always exists an optimal Gaussian information structure, under which states and signals are jointly normally distributed. This arises from the nature of quadratic objective and payoff functions: In expectation, the designer’s objective is determined solely by the covariance matrix of equilibrium actions and states. Moreover, we show that the class of Gaussian information structures is versatile enough to implement any covariance matrix that could emerge under any information structure. Consequently, it is without loss of the designer’s optimality to restrict attention to Gaussian information structures, despite many other alternatives.
A Gaussian information structure is determined solely by its first and second moments, whereas the first is shown to be constant across all information structures. Hence, we can formulate LQG information design as the problem of choosing an optimal covariance matrix. This formulation entails maximizing a linear objective function of a matrix subject to two fundamental constraints: a system of linear constraints regulating each agent’s incentives, and a positive semidefiniteness constraint stemming from the statistical requirement. This problem structure mirrors that of semidefinite programming (SDP), allowing us to employ several methodological tools developed in convex optimization. [^4] SDP is known as a generalization of linear programming; see, e.g., Vandenberghe and Boyd (1996) for further details.
Leveraging the duality principle of SDP, we derive the dual problem of the designer’s maximization problem, formulated as the minimization of a function of Lagrangian multipliers corresponding to each agent’s incentive constraint. Proposition 2 then asserts that the dual problem provides a tight upper bound on the information designer’s optimal value; in other words, strong duality holds. This strong duality leads to the characterization of an optimal information structure based on its “certifiability” by Lagrangian multipliers. Our main technical result, Proposition 3, states that an action-state covariance matrix is optimal if and only if it satisfies the system of matrix equations together with an appropriate choice of Lagrangian multipliers. Moreover, by exploiting the statistical interpretation of these equations, in Corollary 6, we show that the optimal information structure exhibits a linear relationship between the equilibrium strategy profile and the state.
Our characterization is useful particularly when one’s interest lies in checking the optimality of a given information structure. For example, Corollary 7 obtains necessary and sufficient conditions for two extreme forms of information disclosure—no disclosure (i.e., nobody receives an informative signal) and full disclosure (i.e., everyone is fully informed of the true state)—in terms of payoff parameters. It turns out that the optimality of these cases critically hinges on the dimension of the payoff state. Specifically, when the state is multidimensional, each of them is suboptimal for generic payoff structures because the corresponding optimality condition imposes a greater number of linear equations than the number of Lagrangian multipliers that can be chosen as certifiers. In contrast, in the one-dimensional case, we obtain a “bang-bang” result, Corollary 10, which asserts that an optimal information structure is either no disclosure or “aggregative” full disclosure, meaning that each individual may remain partially informed, yet the true state is fully revealed to the entire population.
In general, it is not an easy task to analytically derive optimal information structures since our characterization is based on a high-dimensional system of equations. However, there are several known optimization tools that calculate numerical solutions to SDP problems. Moreover, by assuming a certain degree of symmetry across agents’ payoffs, we demonstrate that the dimension of our problem is effectively reduced so that the analytical characterization of an optimal information structure can be sharpened. For example, in Corollary 13, we show that an optimal information structure continues to hold the aforementioned bang-bang property even when agents are not exposed to a one-dimensional common state.
Lastly, we explore the case where the information designer is limited to public information structures—under which all agents receive identical signals—and study an optimal information structure within this class. This problem holds practical importance when individualized information disclosure is unfeasible due to technological constraints or transparency principles. Incidentally, public LQG information design reduces to a Bayesian persuasion problem with a single representative agent (Kamenica and Gentzkow, 2011), which is analogous to the problem studied by Tamura (2018). Leveraging his solution method, Proposition 16 provides a closed-form expression for the optimal public information structure. It is noteworthy that the optimal public information structure is generically suboptimal within the class of all information structures, including non-public ones. This indicates that the information designer typically benefits from manipulating agents’ higher-order belief regarding others’ private information. [^5] Arieli and Babichenko (2019) and Hoshino (2022) also highlight the importance of leveraging strategic uncertainty in games with finite actions and binary states, while their settings do not exhibit the generic gap between private and public information structures as in our result.
This paper contributes to the growing literature on information design, which has yielded a wide array of economic applications in various contexts. [^6] See survey papers by Bergemann and Morris (2019) and Kamenica (2019). Applications of information design with multiple agents include matching markets (Ostrovsky and Schwarz, 2010), voting games (Alonso and Câmara, 2016), congestion games (Das et al., 2017), auctions (Bergemann et al., 2017a), contests (Zhang and Zhou, 2016), and stress testing (Inostroza and Pavan, 2023). While the literature has investigated several applications and general approaches, research on information design with multiple interacting agents remains relatively sparse due to technical challenges arising from agents’ intricately interlinked incentives in the face of strategic uncertainty. The main contribution of this paper lies in formulating the information designer’s problem in the SDP form and in developing the duality-based approach to characterize the optimum.
Our approach is built on an outcome-based approach, which involves the following decomposition of the problem into two steps: [^7] See Bergemann and Morris (2013, 2016a, 2016b, 2019), Bergemann et al. (2015, 2017b), and Taneva (2019). Mathevet et al. (2020) develop a different belief-based approach, which is a natural analogue of the concavification method of Kamenica and Gentzkow (2011) for Bayesian persuasion problems. First, identify the feasible set of outcomes—i.e., action-state joint distributions—that could be induced by the designer. Second, determine which of these outcomes would maximize the designer’s objective. In symmetric LQG games with agents sharing identical payoff functions, Bergemann and Morris (2013) characterize the feasible outcomes in terms of linear constraints on the first and second moments of outcomes by restricting attention to a stylized class of conditionally i.i.d. symmetric information structures. [^8] Bergemann et al. (2017b) and Miyashita (2024) extend these constraints to asymmetric LQG games with finite/continuum populations, respectively. We further show that the constraints are necessarily satisfied even when signals are not Gaussian; this result serves as a building block to establish the optimality of Gaussian information structures. However, it has remained unclear (i) whether there exists an asymmetric or non-Gaussian information structure that surpasses the optimal conditional i.i.d. information structure, and (ii) how to solve the second step optimization when agents exhibit asymmetry in terms of payoffs and/or information. We advance the analysis of Bergemann and Morris (2013) and subsequent studies by broadening the class of environments considered and by recasting the problem in the SDP form. This comprehensive extension enables us to establish the “global” optimality of Gaussian information structures and characterize it via duality.
The duality methodology is widely used in several optimization problems, including information design. [^9] For example, Kolotilin (2018), Dworczak and Martini (2019), and Kolotilin et al. (forthcoming) employ the duality methodology to study persuasion problems with a single receiver. Among the closest to our work is Smolin and Yamashita (2023), which study information design in games with strictly concave utility functions. In their framework, the information design problem is formulated as infinite-dimensional linear programming, with the objective being linear in a probability measure over actions and states, subject to linear constraints. They derive the associated dual problem and establish weak duality. In contrast, under quadratic specifications, we formulate the problem as SDP with respect to an action-state covariance matrix, involving only finitely many variables. This reduction in dimensionality enhances tractability, allowing us to establish strong duality and complementary slackness conditions. Furthermore, all economic applications provided by Smolin and Yamashita (2023) are nested within our framework, and we demonstrate that their insights regarding the optimality of Gaussian information structures extend generally to all LQG environments.
This paper also relates to the literature on network interventions, an active topic studied across many disciplines. [^10] See, for example, Valente (2012) for a general introduction to the subject. In economics, the seminal work by Ballester et al. (2006) formulate the intervention problem as a network game and quantify the structural importance of each individual by comparing equilibrium group behavior with and without the individual in the network. Galeotti et al. (2020) consider the planner’s intervention of changing targeted individuals’ base action levels. Within our model, these interventions can be precisely captured as exogenous changes in payoff structures. In contrast, our focus lies in altering information structures by considering the effect of information disclosure on equilibrium behavior. [^11] This type of informational intervention has garnered significant attention in recent empirical literature. For example, Banerjee et al. (2013) examine the effect of information transmission on microfinance participation using demographic and social network data. Thus, our study provides complementary insights for planners aiming to influence group behavior by manipulating individuals’ incentives through direct interventions and indirect strategic spillovers to their neighbors.
The rest of the paper is organized as follows. Section 2 conducts numerical simulations in a leading example to motivate our general analysis. Section 3 introduces the model of LQG information design and characterizes an equilibrium through moment restrictions. Section 4 formulates LQG information design as an SDP problem, sets up the dual problem, and offers our main results. Sections 5 and 6 focus on special cases: symmetric environments with general information structures and general environments with public information structures, respectively. All omitted proofs can be found in the appendix.
2 Leading example
We consider a variant of network games studied by Galeotti et al. (2020). There are finitely many agents $[n]\equiv\{1,\ldots,n\}$. Let $G\equiv[g_{ij}]_{n\times n}$ be an $n\times n$ binary matrix with $g_{ij}\in\{0,1\}$ for each $i,j\in[n]$. This matrix $G$, called a network, captures the strategic externality among agents. Specifically, $g_{ij}=1$ is interpreted as indicating the existence of a directed link to agent $i$ from $j$. Assume that $g_{11}=\cdots=g_{nn}=0$, i.e., there are no self loops.
For now, we focus on “symmetric” networks, where any pair of agents play symmetric roles, in order to facilitate numerical calculations. Specifically, the two symmetric networks considered in this section are an island-shaped one and a cyclical one, illustrated in Figure 1. In each of these networks, although agents are not entirely homogeneous, as they have different sets of neighbors, there is an intuitive sense in which they appear symmetric. Indeed, for any distinct pair of agents $(i,j)$, we can find a permutation, which sends $i$ to $j$, and which keeps the network structure unchanged by relabelling the names of agents according to the permutation.121212 We formalize this notion of symmetry in Section 5 basced on the automorphism group of a network.
123456 126354
Figure 1: In the island-shaped network (left), six agents are divided into two subgroups of size $3$, with each agent receiving strategic influences from the other two agents within their own subgroup. In the (directed) cyclical network (right), six agents are arranged on a circle, with each agent receiving strategic influences from the two adjacent agents in a clockwise direction.
Each agent $i$ simultaneously chooses an investment level $a_{i}\in\mathbb{R}$ at a cost $a_{i}^{2}/2$. The private marginal return on investment depends on an unknown state of the world $\theta_{i}$ and the localized aggregate of other agents’ investment levels as follows:
$\phi_{i}(a,\theta_{i})\equiv\theta_{i}-q\sum_{j\neq i}g_{ij}a_{j},\quad\text{where}\quad q\in\mathbb{R}.$
Here, $q$ represents the degree of strategic externality. The case of $q>0$ captures strategic substitutes, where agents have incentives to be differentiated from their neighbors (Vives, 1999, Goyal and Moraga-Gonzalez, 2001). In contrast, the case of $q<0$ reflects strategic complements (Ballester et al., 2006), where each agent’s marginal return is enhanced when her neighbors choose higher investment levels.
For an action profile $a\equiv(a_{1},\ldots,a_{n})$, the agent $i$’s payoff is given by
$u_{i}(a,\theta_{i})\equiv\phi_{i}(a,\theta_{i})\cdot a_{i}-\frac{a_{i}^{2}}{2}.$
This payoff structure bears similarity to the one in Galeotti et al. (2020), but a crucial difference is that each agent’s base-level marginal return $\theta_{i}$ is a random variable. We assume that $\theta\equiv(\theta_{1},\ldots,\theta_{n})$ follows a multivariate normal distribution with mean zero and the covariance matrix $Z$ such that $\operatorname{Var}[\theta_{i}]=1$ and $\operatorname{Cov}[\theta_{i},\theta_{j}]=\rho$ for any $i\neq j$. Here, $\rho\in[0,1]$ represents the correlation coefficient between different agents’ payoff-relevant states. The case with $\rho=1$ is referred to as the common-value case; in this case, $\theta_{i}$ and $\theta_{j}$ are perfectly correlated with each other. The other cases with $\rho<1$ are referred to as private-value cases.
Each agent does not know the state realization when deciding on her investment level. Instead, there is an (information) designer who can send a private signal $t_{i}$ to each agent $i$ according to a pre-announced joint probability distribution $\pi$ over $(t,\theta)\equiv(t_{1},\ldots,t_{n},\theta_{1},\ldots,\theta_{n})$, referred to as an information structure. The information designer can manipulate agents’ incentives by choosing an information structure, as their private signals can be correlated with one another and with the state. For simplicity, assume that the information designer aims at maximizing the social welfare—i.e., the sum of agents’ payoffs—by anticipating that they behave in a way of maximizing their “own” payoffs.
Given an information structure $\pi$, agent $i$’s strategy $\sigma_{i}$ is a function that assigns $\sigma_{i}(t_{i})\in\mathbb{R}$ to each realization of a private signal $t_{i}$. A strategy profile $(\sigma_{1}^{\pi},\ldots,\sigma^{\pi}_{n})$ is said to be a Bayesian Nash equilibrium (under $\pi$) if each agent’s strategy $\sigma^{\pi}_{i}$ maximizes her conditional expected payoff given $t_{i}$, provided that the opponents’ adopt the strategies $\sigma^{\pi}_{-i}=(\sigma^{\pi}_{j})_{j\neq i}$. Evaluated in the Bayesian Nash equilibrium under the chosen $\pi$, it can be shown that the information designer’s objective function is proportional to
$\sum_{i\in[n]}\operatorname{Var}\quantity[\sigma^{\pi}_{i}].$ (1)
Also, the best-response condition for each agent can be reduced to the following obedience condition (Bergemann and Morris, 2013): [^13] Together with positive semidefiniteness, our Lemma 2 shows that this obedience condition is not only necessary but also sufficient for any covariance matrix to be inducible in BNE under some information structure.
$\operatorname{Var}\quantity[\sigma^{\pi}_{i}]=\operatorname{Cov}\quantity[\sigma^{\pi}_{i},\theta_{i}]-q\sum_{j\neq i}g_{ij}\operatorname{Cov}\quantity[\sigma^{\pi}_{i},\sigma^{\pi}_{j}],\quad\forall i\in[n].$ (2)
Notably, the objective function (1), as well as each agent’s best response condition (2), depends on the choice of $\pi$ only through the covariance matrix of the action-state joint distribution that is induced by $\pi$. Specifically, if we denote by $x_{ij}=\operatorname{Cov}[\sigma^{\pi}_{i},\sigma^{\pi}_{j}]$ and $y_{ij}=\operatorname{Cov}[\sigma^{\pi}_{i},\theta_{j}]$ for $i,j\in[n]$, then (1) can be rewritten as
$\sum_{i\in[n]}x_{ii}.$ (3)
Also, (2) can be rewritten as
$x_{ii}=y_{ii}-q\sum_{j\neq i}g_{ij}x_{ij},\quad\forall i\in[n].$ (4)
In addition, since $X=[x_{ij}]_{n\times n}$ and $Y=[y_{ij}]_{n\times n}$ are components of the covariance matrix, they necessarily satisfy the following statistical requirement:
$\matrixquantity[X&Y\\ Y^{\top}&Z]\text{ is positive semidefinite}.$ (5)
Therefore, the information designer’s problem is mathematically equivalent to the constrained maximization of (3) subject to (4) and (5).
This problem structure entails the maximization of a linear objective function (3) subject to a bunch of linear constraints (4) and a positive semidefiniteness constraint (5). In fact, these are the defining mathematical features of optimization problems known as semidefinite programming (SDP), admitting several optimization tools that are readily available. Specifically, the current problem can be numerically solved by utilizing CVXPY, a Python-based modeling language for convex optimization. Below, we present some findings from these numerical simulations.
A fundamental finding is that when a network exhibits symmetry across agents, the optimal information structure also reflects this symmetry. Specifically, in both island-shaped and cyclical networks, each numerically obtained optimal information structure results in an action-state covariance matrix that remains invariant under the relabeling of agents’ indexes. This allows us to focus on a representative agent, say agent 1, to describe the configuration of the numerical solutions.
Figure 2: $P_{\rm agg}$, $P_{\rm indv}$, and $N_{\rm indv}$ are plotted as functions of $q$ for different values of $\rho$. These graphs are obtained for a directed cyclical network with $n=6$, as shown in the right panel of Figure 1. Similar qualitative results are observed for other symmetric networks, such as island-shaped and undirected cyclical networks.
Second, any optimal information structure fully discloses the state to the entire population of agents in “aggregate” form, while each individual agent may remain partially informed. To illustrate this finding, we use the following measure:
$P_{\rm indv}\equiv\frac{\operatorname{Var}\quantity[\theta_{1}]-\operatorname{Var}[\theta_{1}\mid\sigma^{\pi}_{1}]}{\operatorname{Var}\quantity[\theta_{1}]},$
which quantifies the informativeness of the first agent’s private information as the percentage reduction in the variance of $\theta_{1}$ conditioning on the agent’s information set. Similarly, the informativeness of the entire population’s private information about $\theta_{1}$ is measured by
$P_{\rm agg}\equiv\frac{\operatorname{Var}\quantity[\theta_{1}]-\operatorname{Var}[\theta_{1}\mid\sigma^{\pi}_{1},\ldots,\sigma^{\pi}_{n}]}{\operatorname{Var}\quantity[\theta_{1}]}.$
In Figure 2, the values of $P_{\rm indv}$ and $P_{\rm agg}$ are plotted as functions of $q$. We can see that $P_{\rm agg}$ consistently equals $1$, indicating that the realization of $\theta_{1}$ is always fully revealed to the entire population. In contrast, in the common-value case with $\rho=1$, the individual state variance reduction $P_{\rm indv}$ is less than $1$ when $q$ is positive and large, meaning each individual agent remains partially informed when exposed to large strategic substitutes. Moreover, in all private-value cases with $\rho<1$, the agent is only partially informed, except when $q=0$ so that no strategic externalities are present.
Thus, the information designer should fully disclose the state to the population, while typically keeping each agent partially informed. This highlights the importance of private information in optimal interventions, as such aggregate full disclosure is only feasible when personalized information can be provided to different agents.
Third, whether an optimal information structure features “non-fundamental volatility” depends on the statistical nature of the fundamental uncertainty. To explore this, we define the following measure:
$N_{\rm indv}\equiv\frac{\operatorname{Var}\quantity[\sigma^{\pi}_{1}\mid\theta_{1},\ldots,\theta_{n}]}{\operatorname{Var}\quantity[\sigma^{\pi}_{1}]},$
which quantifies the non-fundamental volatility in the agent’s equilibrium strategy as the percentage of variance in $\sigma^{\pi}_{1}$ that can be explained by conditioning on fundamental uncertainty. Note that $N_{\rm indv}$ ranges from 0 to 1, with lower values indicating less non-fundamental volatility in the agent’s strategy. As shown in Figure 2, in the common-value case, $N_{\rm indv}$ takes positive values unless the state is fully disclosed to each individual agent. In contrast, $N_{\rm indv}$ equals $1$ in all private-value cases, meaning the equilibrium never exhibits non-fundamental volatility.
This finding suggests that the way the information designer optimally achieves aggregate full disclosure depends on whether agents’ payoff-relevant states are common or private. In the common-value case, non-fundamental volatility in $\sigma^{\pi}_{1}$ implies that each agent’s signal contains idiosyncratic noise, which cancels out when aggregated. In contrast, each agent’s signal is deterministic in $\theta_{1},\ldots,\theta_{n}$ in private-value cases, meaning the designer need not introduce additional randomness beyond the fundamental uncertainty. Instead, the optimal information structure sends signals that incorporate not only the agent’s own payoff-relevant state but also those of others, thereby guiding strategies toward the optimal outcome by leveraging their strategic interactions.
3 Model
The following matrix notations are used. For a positive integer $n\in\mathbb{N}$, let $[n]\equiv\{1,\ldots,n\}$. Given $n,m\in\mathbb{N}$, we denote by $\mathcal{M}^{n,m}$ the set of $n\times m$ matrices, $\mathcal{S}^{n}$ the set of $n\times n$ symmetric matrices, and $\mathcal{S}^{n}_{\rm diag}$ the set of $n\times n$ diagonal matrices. Denote by $A^{\top}\in\mathcal{M}^{m,n}$ the transpose of $A\in\mathcal{M}^{n,m}$. For $A,B\in\mathcal{S}^{n}$, we write $A\succeq B$ (resp. $A\succ B$) when $A-B$ is positive semidefinite (resp. positive definite). Let $\mathcal{S}^{n}_{+}$ (resp. $\mathcal{S}^{n}_{++}$) be the subset of $\mathcal{S}^{n}$ collecting all $A\in\mathcal{S}^{n}$ such that $A\succeq O$ (resp. $A\succ O$). For $A\in\mathcal{S}_{++}^{n}$, let $A^{1/2}$ be the matrix root of $A$, i.e., the unique matrix such that $A^{1/2}\in\mathcal{S}^{n}_{++}$ and $A^{1/2}A^{1/2}=I$. Denote by $A^{-1/2}=(A^{1/2})^{-1}$.
[…]
3.1 Basic games and information structures
There are $n$ agents indexed by $i\in[n]$, each of whom chooses a real number $a_{i}\in\mathbb{R}$ as an action. Agent $i$’s payoff is quadratically dependent on an action profile $a\equiv(a_{1},\ldots,a_{n})\in\mathbb{R}^{n}$ and a payoff state $\theta\equiv(\theta_{1},\cdots,\theta_{m})\in\mathbb{R}^{m}$, where $m\in\mathbb{N}$. Specifically, let the payoff function be given as
$u_{i}(a,\theta)\equiv-\underbrace{\frac{q_{ii}}{2}\cdot a_{i}^{2}}_{\begin{subarray}{c}\text{private cost}\\ \text{of own action}\end{subarray}}+\ \underbrace{a_{i}\cdot\quantity(\sum_{k\in[m]}r_{ik}\theta_{k}-\sum_{j\in[n]\setminus\{i\}}q_{ij}a_{j})}_{\text{returns from own action}}+\underbrace{h_{i}(a_{-i},\theta)}_{\text{pure externalities}},$ (6)
where $q_{ij}$ and $r_{ik}$ are constant, and $h_{i}(a_{-i},\theta)$ is an arbitrary function of the opponents’ actions $a_{-i}\equiv(a_{j})_{j\neq i}$ and a payoff state $\theta$. Let $Q\equiv[q_{ij}]_{n\times n}$ and $R\equiv[r_{ik}]_{n\times m}$ be the matrices consisting of the coefficients in (6).
Throughout this paper, we always assume the following regularity conditions.
Assumption 1.
The matrix $Q+Q^{\top}$ is positive definite. A payoff state $\theta$ follows a multivariate normal distribution with mean $\bar{\theta}\equiv(\operatorname{E}[\theta_{1}],\ldots,\operatorname{E}[\theta_{m}])$ and variance $Z\equiv\operatorname{Var}[\theta]=[\operatorname{Cov}[\theta_{k},\theta_{k^{\prime}}]]_{m\times m}\neq O$.
The positive definiteness of $Q+Q^{\top}$ is satisfied in many applications and guarantees the unique existence of an equilibrium under an arbitrary information structure. That $Z\neq O$ is a minimal assumption to make our problem non-trivial. In the payoff function (6), $h_{i}$ has no influence on agent’s decisions so that their strategic concerns can be summarized by the matrices $Q$ and $R$ alone. Also, since $\theta$ is normally distributed, its distribution can be summarized by the mean vector $\bar{\theta}$ and the covariance matrix $Z$, while $\bar{\theta}$ will be of no importance in our problem. In this regard, we call the profile $(Q,R,Z)$ a basic game.
An (information) designer sends a private signal $t_{i}$ to each agent $i$ according to a pre-announced joint probability distribution over $(t,\theta)\equiv(t_{1},\ldots,t_{n},\theta_{1},\ldots,\theta_{n})$. This distribution $\pi$ is called an information structure. When $t$ and $\theta$ are jointly normally distributed, $\pi$ is called a Gaussian information structure. Let $\Pi^{*}$ denote the set of all information structures, and let $\Pi^{\rm g}\subseteq\Pi^{*}$ be its subset consisting of all Gaussian ones.
Together with a basic game $(Q,R,Z)$, an information structure $\pi$ defines a Bayesian game. In particular, we say that the game is linear-quadratic-Gaussian (LQG) when $\pi$ is Gaussian.
3.2 Bayesian Nash equilibrium
Given any information structure $\pi$, agent $i$’s strategy is given as a measurable function $\sigma_{i}$ of $t_{i}$ with $\operatorname{E}|\sigma_{i}(t_{i})|^{2}<\infty$. A strategy profile $\sigma\equiv(\sigma_{1},\ldots,\sigma_{n})$ is a Bayesian Nash equilibrium (BNE) under $\pi$ if each agent’s strategy maximizes her conditional expected payoff given the opponents’ strategies; that is,
$\sigma_{i}(t_{i})\in\mathop{\rm arg~{}max}\limits_{a_{i}\in\mathbb{R}}\operatorname{E}^{\pi}\quantity[u_{i}((a_{i},\sigma_{-i}(t_{-i})),\theta)\mid t_{i}],\quad\forall i\in[n]$
hods almost surely (a.s.), where $\sigma_{-i}(t_{-i})\equiv(\sigma_{j}(t_{j}))_{j\neq i}$ and $\operatorname{E}^{\pi}[\cdot\mid t_{i}]$ is the conditional expectation operator given $t_{i}$ with respect to $\pi$. We simply write $\operatorname{E}\quantity[\cdot]$ instead of $\operatorname{E}^{\pi}\quantity[\cdot]$ as long as there is no risk of confusion. It is known that a unique BNE exists for any information structure under Assumption 1; see Proposition 4 of Ui (2016). In light of this fact, we denote by $\sigma^{\pi}$ the unique BNE under each information structure $\pi$.
Lemma 1.
If $Q+Q^{\top}$ is positive definite, then a unique BNE $\sigma^{\pi}$ exists for each $\pi\in\Pi^{*}$. In particular, if $\pi\in\Pi^{\rm g}$, then $(\sigma^{\pi},\theta)$ is jointly normally distributed.
Since $u_{i}$ is quadratic and concave, as specified in (6), we can see that a strategy profile $\sigma$ is a BNE if and only if
$\sum_{j=1}^{n}q_{ij}\operatorname{E}\quantity[\sigma_{j}\mid t_{i}]=\sum_{k=1}^{m}r_{ik}\operatorname{E}\quantity[\theta_{k}\mid t_{i}],\quad\forall i\in[n],$ (7)
where $\sigma_{j}=\sigma_{j}(t_{j})$. This condition displays linear constraints that the conditional expectations of the state and equilibrium actions have to obey. Moreover, applying the law of iterated expectations, we can show that the unconditional counterpart of (7) holds for expectations and covariances. Specifically, the unique equilibrium $\sigma^{\pi}$ under any (possibly, non-Gaussian) information structure necessary satisfies the following moment restrictions:
$\sum_{j=1}^{n}q_{ij}\operatorname{E}[\sigma_{j}^{\pi}]=\sum_{k=1}^{m}r_{ik}\operatorname{E}[\theta_{k}],\quad\forall i\in[n],$ (8)
$\sum_{j=1}^{n}q_{ij}\operatorname{Cov}[\sigma_{i}^{\pi},\sigma_{j}^{\pi}]=\sum_{k=1}^{m}r_{ik}\operatorname{Cov}[\sigma_{i}^{\pi},\theta_{k}],\quad\forall i\in[n].$ (9)
In matrix form, these restrictions are rewritten as
$Q\operatorname{E}\quantity[\sigma^{\pi}]=R\bar{\theta}\quad\text{and}\quad\operatorname{diag}(Q\operatorname{Var}\quantity[\sigma])=\operatorname{diag}\quantity(R\operatorname{Cov}\quantity[\theta,\sigma]).$
Since $Q$ is invertible under Assumption 1, the first equation uniquely pins down the expectation of an action profile as $\bar{a}\equiv\operatorname{E}\quantity[\sigma^{\pi}]=Q^{-1}R\bar{\theta}$ across all information structures. From the information designer’s perspective, this indicates that there is no room for changing the expected action levels by manipulating information structures.
The second equation places some restrictions on the variance matrix of an equilibrium action-state joint distribution, but in a less stringent manner than the first equation. Specifically, let $\mathbf{M}^{\pi}$ denote the variance matrix that arises in the unique equilibrium under an information structure $\pi$; that is,
Since $\mathbf{M}^{\pi}$ is a symmetric matrix, and since its lower-right block is exogenously fixed by the distribution of $\theta$, it contains ${n(n+2m+1)}/{2}$ free variables, while there are only $n$ linear equations in (9) to be satisfied by these variables.
In the next lemma, we establish that these restrictions are not only necessary but also sufficient for a given $(n+m)\times(n+m)$ positive semidefinite matrix to be inducible as the covariance matrix of an equilibrium action-state joint distribution under some Gaussian information structure. This extends the result of Bergemann and Morris (2013) to the present setting with heterogeneous payoff functions and a multi-dimensional state with arbitrary information structures. [^14] Bergemann and Morris (2013) focus on symmetric Gaussian information structures and symmetric basic games with $m=1$ and show that the constraint (10) is necessary and sufficient for the equilibrium covariance matrices under Gaussian information structures. Proposition 5 of Miyashita and Ui (2024) provides a similar sufficiency result in an asymmetric setting with $m=1$ and a continuum of agents.
Lemma 2.
For any information structure $\pi\in\Pi^{*}$, if we write $X=\operatorname{Var}[\sigma^{\pi}]$ and $Y=\operatorname{Cov}[\sigma^{\pi},\theta]$, then it holds that
$\operatorname{diag}(XQ^{\top})=\operatorname{diag}(YR^{\top})\quad\text{and}\quad\mathbf{M}_{X,Y}\equiv\matrixquantity[X&Y\\ Y^{\top}&Z]\in\mathcal{S}_{+}^{n+m}.$ (10)
Conversely, if $X\in\mathcal{S}^{n}$ and $Y\in\mathcal{M}^{n\times m}$ satisfy (10), then there exists a Gaussian information structure $\pi\in\Pi^{\rm g}$ such that $\mathbf{M}^{\pi}=\mathbf{M}_{X,Y}$.
Definition 1.
A pair of matrices $(X,Y)\in\mathcal{S}^{n}\times\mathcal{M}^{n,m}$ is said to be (primal) feasible under $(Q,R)$ if it satisfies (10).
We simply say that $(X,Y)$ is feasible when $(Q,R)$ is clear from the context.
Lemma 2 implies that the set of the covariance matrices under BNE over all information structures coincides with that over all Gaussian ones, since
$\quantity{\mathbf{M}^{\pi}:\pi\in\Pi^{*}}=\quantity{\mathbf{M}_{X,Y}:\text{$(X,Y)$ is primal feasible}}=\quantity{\mathbf{M}^{\pi}:\pi\in\Pi^{\rm g}}.$ (11)
This observation will guarantee the optimality of Gaussian information structures in our information design problem, where the designer’s objective is given as a quadratic function in the state and actions.
3.3 LQG information design
Given a basic game $(Q,R,Z)$, we now introduce an (information) designer, who chooses an information structure $\pi$ from a set of feasible information structures $\Pi\subseteq\Pi^{*}$. The designer can make a commitment to provide information according to the following timeline, which is a standard assumption in the literature: First, the designer chooses $\pi\in\Pi$ and informs all agents of $\pi$. Second, the realization of $(t,\theta)$ is drawn according to $\pi$, and each agent $i$ observes $t_{i}$. Third, agents follow the unique BNE $\sigma^{\pi}$ under the chosen information structure $\pi$.
The designer aims to maximize the expected value of a given objective function $v(a,\theta)$ by anticipating that the unique BNE is played under the chosen information structure. We assume that $v$ is quadratically dependent on the state $\theta$ and an action profile $a$ as follows:
Here, $\bar{v}$ is any linear function of $(a,\theta)$, and $\mathbf{V}$ is a symmetric matrix consisting of the coefficients of quadratic terms in $v(a,\theta)$.
LQG information design is the problem of finding an information structure that maximizes the designer’s expected objective over the set of feasible information structures $\Pi$:
$\max_{\pi\in\Pi}\quad\operatorname{E}\quantity[v(\sigma^{\pi},\theta)].$ (12)
An information structure $\pi\in\Pi$ is said to be optimal in $\Pi$ if it is a solution to this problem. When a feasible set is unrestricted, i.e., $\Pi=\Pi^{*}$, we simply say that such $\pi$ is optimal, and the value of (12) is denoted by $v^{\rm p}$.
This problem gains tractability by reformulating the objective function as a linear function of the action-state covariance matrix. Specifically, for any $\pi\in\Pi^{*}$, since the trace of a matrix is commutative,
[…]
By the linearity of $\bar{v}$, the objective can then be rewritten as $\mathbf{V}\bullet\mathbf{M}^{\pi}+\bar{v}(\bar{a},\bar{\theta})$, wherein the second term is uncontrollable for the designer since $\bar{\theta}$ is exogenously fixed and $\bar{a}$ remains constant across all information structures. Thus, it is without loss of generality to adopt the following assumption.
Assumption 2.
An objective function $v$ is given by (3.3) with $\bar{v}(\bar{a},\bar{\theta})=0$.
Under this assumption, the objective function in (12) can be written as
$\operatorname{E}\quantity[v(\sigma^{\pi},\theta)]$ $=\mathbf{V}\bullet\mathbf{M}^{\pi}$
$=V\bullet\operatorname{Var}\quantity[\sigma^{\pi}]+W\bullet\operatorname{Cov}\quantity[\sigma^{\pi},\theta]$
$=\sum_{i=1}^{n}\sum_{j=1}^{n}v_{ij}\operatorname{Cov}\quantity[\sigma^{\pi}_{i},\sigma^{\pi}_{j}]+\sum_{i=1}^{n}\sum_{k=1}^{m}w_{ik}\operatorname{Cov}\quantity[\sigma^{\pi}_{i},\theta_{k}],$ (13)
where $v_{ij}$ is the $(i,j)$-entry of $V$, representing the marginal contribution of the covariance between agent $i$’s action and agent $j$’s action to the objective. Similarly, $w_{ik}$ is the $(i,k)$-entry of $W$, representing the marginal contribution of the covariance between agent $i$’s action and the $k$-th dimension of the state.
We call $\mathbf{V}\bullet\mathbf{M}^{\pi}$ the value of $\pi$. For example, consider no disclosure, denoted by $\underline{\pi}$, where all agent are informed of nothing. Since agents’s equilibrium strategies are constant at $\bar{a}$ under $\underline{\pi}$, the value of $\underline{\pi}$ equals to 00. Alternatively, consider full disclosure, denoted by $\overline{\pi}$, where each agent is fully informed of the state realization (and nothing else). Since $\theta$ becomes common knowledge under $\overline{\pi}$, agents follow the action profile $Q^{-1}R\theta$ in each state $\theta$. The value of $\overline{\pi}$ is then given as $\overline{C}\bullet Z$, where
$\overline{C}\equiv(Q^{-1}R)^{\top}VQ^{-1}R+\frac{(Q^{-1}R)^{\top}W+W^{\top}Q^{-1}R}{2}$ (14)
is a symmetric matrix defined by the payoff parameters. Notably, the values of $\underline{\pi}$ and $\overline{\pi}$ can both be expressed as a Frobenius inner product $C\bullet Z$, where $C=O$ for $\underline{\pi}$ and $C=\overline{C}$ for $\overline{\pi}$. As we will see in Corollary 5, this is a general property of optimal information structures.
These two information structures represent extremes in terms of the informativeness of private signals about the state. We refer to all other cases as partial disclosure. The optimal choice of $\pi$ depends on the agents’ payoff parameters, $Q$ and $R$, and the designer’s payoff parameters, $V$ and $W$. We collect all these parameters in a profile $\mathscr{P}\equiv(Q,R,V,W)$, referred to as the payoff structure of the LQG information design.
4 Main results
In this section, we seek for an optimal information structure by assuming that all information structures are feasible, i.e., $\Pi=\Pi^{*}$. In doing so, Lemma 2 brings us crucial implications. As indicated by (11), the set of equilibrium covariance matrices over all information structures coincides with that over all Gaussian ones, which is a closed convex set. Thus, the information designer can focus exclusively on Gaussian information structures without losing optimality since the objective function of LQG information design is given as a function of an equilibrium covariance matrix.
Proposition 1.
There exists a Gaussian information structure that attains $v^{\rm p}$.
This insight is novel in the literature, although some have predicted similar results to hold in more specific contexts, e.g. Bergemann and Morris (2013), Bergemann et al. (2015), Tamura (2018), and Smolin and Yamashita (2023). In light of Proposition 1, henceforth, we can restrict our attention to the class $\Pi^{\rm g}$ of Gaussian information structures rather than considering all information structures in $\Pi^{*}$. We note, however, that an optimal Gaussian information structure is generally not uniquely optimal in $\Pi^{*}$. As such, a non-Gaussian information structure is also optimal as long as it induces the same action-state covariance matrix as the optimal Gaussian information structure.
4.1 Characterization via duality
Lemma 2 further establishes an equivalent representation of LQG information design that allows applications of mathematical tools from convex optimization. Recall that a matrix $\mathbf{M}$ is an equilibrium covariance matrix under some Gaussian information structure $\pi\in\Pi^{\rm g}$ if and only if $\mathbf{M}=\mathbf{M}_{X,Y}$ for some feasible pair of matrices $(X,Y)$. By means of this observation, we can equivalently state the problem (12) as follows:
$v^{\rm p}=\max_{X\in\mathcal{S}^{n},\,Y\in\mathcal{M}^{n,m}}\mathbf{V}\bullet\mathbf{M}_{X,Y}\quad\text{subject to}\quad\operatorname{diag}\quantity(XQ^{\top})=\operatorname{diag}\quantity(YR^{\top}),\quad\mathbf{M}_{X,Y}\succeq O,$ (P)
where the objective function is linear in a feasible pair of matrices $(X,Y)$.
This formulation exhibits commonality with convex optimization known as semidefinite programming (SDP). SDP is an extension of classical linear programming, characterized as the maximization of a linear function of a positive semidefinite matrix subject to linear constraints. The associated dual problem of (P) is given by the following, where the optimization variables are a diagonal matrix $\Lambda=\operatorname{Diag}\quantity(\lambda_{1},\ldots,\lambda_{n})$ and a symmetric matrix $\Gamma=\quantity[\gamma_{kk^{\prime}}]_{m\times m}$:
Definition 2.
A pair of matrices $(\Lambda,\Gamma)\in\mathcal{S}^{n}_{\rm diag}\times\mathcal{S}^{m}$ is said to be (dual) feasible if it satisfies the constraint of (4.1).
The interpretation is that each $\gamma_{kk^{\prime}}$ is the shadow price of $\operatorname{Cov}\quantity[\theta_{k},\theta_{k^{\prime}}]$, and $\lambda_{i}$ is the Lagrangian multiplier capturing the value of relaxing agent $i$’s incentive constraint. The problem then takes the form of minimizing the total expenditure, $Z\bullet\Gamma$ subject to the dual feasibility constraint. As such, the dual feasibility constraint essentially says that the total expenditure is no less than the designer’s value, $\mathbf{V}\bullet\mathbf{M}_{X,Y}$, for all primal feasible $(X,Y)$. The next proposition formally proves this. Furthermore, it establishes strong duality, i.e., there is no gap between $v^{\rm p}$ and $v^{\rm d}$.
Proposition 2.
Both primal problem (P) and dual problem (4.1) admit solutions. For any primal feasible $(X,Y)$ and any dual feasible $(\Lambda,\Gamma)$, it holds that $Z\bullet\Gamma\geq\mathbf{V}\bullet\mathbf{M}_{X,Y}$. Moreover, strong duality holds, i.e., $v^{\rm p}=v^{\rm d}$.
We can simplify the dual problem (4.1) using the following augmented incentive matrices:
$A_{\Lambda}\equiv\frac{\Lambda Q+Q^{\top}\Lambda}{2}-V\quad{\rm and}\quad B_{\Lambda}\equiv\frac{\Lambda R+W}{2},$ (15)
by which (4.1) is rewritten as
In addition, let $C_{\Lambda}$ be defined by
$C_{\Lambda}\equiv B_{\Lambda}^{\top}A_{\Lambda}^{+}B_{\Lambda},$
where $A_{\Lambda}^{+}$ denotes the Moore–Penrose pseudoinverse of $A_{\Lambda}$. [^15] The Moore–Penrose inverse of $M\in\mathcal{M}^{m,n}$ is the unique matrix $M^{+}\in\mathcal{M}^{n,m}$ that satisfies the following conditions: $MM^{+}M=M$, $(MM^{+})^{\top}=MM^{+}$, $M^{+}MM^{+}=M^{+}$, and $(M^{+}M)^{\top}=M^{+}M$. Then, the block structure in (4.1) implies $(\Lambda,\Gamma)$ is dual feasible if and only if (i) $A_{\Lambda}$is positive semidefinite, (ii) the generalized Schur complement $\Gamma-C_{\Lambda}$ is positive semidefinite, and (iii) $B_{\Lambda}=A_{\Lambda}A_{\Lambda}^{+}B_{\Lambda}$holds true. [^16] See Proposition 10.2.5 of Bernstein (2018), which we also record as Lemma 7 in Appendix A. The last condition, $B_{\Lambda}=A_{\Lambda}A_{\Lambda}^{+}B_{\Lambda}$, essentially says that each column vector of $B_{\Lambda}$ belongs to the range of $A_{\Lambda}$, which is trivially satisfied when $A_{\Lambda}$ has full rank. Therefore, since $Z\succeq O$, the optimal value of (4.1) shall take the the form of $Z\bullet C_{\Lambda}$ that can be achieved when $\Gamma$ equals the “minimum” matrix $C_{\Lambda}$. Hence, we can reduce (4.1) as
We simply say that $\Lambda\in\mathcal{S}^{n}_{\rm diag}$ is (dual) feasible when $(\Lambda,C_{\Lambda})$ is dual feasible, and it is (dual) optimal if it is feasible and $v^{\rm d}=Z\bullet C_{\Lambda}$.
We are now ready to present the main result of this paper, which identifies an optimality condition in terms of Lagrange multipliers. This condition enables us to determine the optimality of an information structure through the notion of “certifiability” by $\Lambda$.
Proposition 3.
A Gaussian information structure $\pi\in\Pi^{\rm g}$ is optimal if and only if there exists a dual feasible matrix $\Lambda\in\mathcal{S}^{n}_{\rm diag}$ that satisfies
$A_{\Lambda}\operatorname{Var}\quantity[\sigma^{\pi}]=B_{\Lambda}\operatorname{Cov}\quantity[\theta,\sigma^{\pi}]\quad\text{and}\quad A_{\Lambda}\operatorname{Cov}\quantity[\sigma^{\pi},\theta]=B_{\Lambda}Z.$ (16)
This condition is equivalent to either of the following:
$A_{\Lambda}\operatorname{Var}[\sigma^{\pi}]=B_{\Lambda}\operatorname{Cov}[\sigma^{\pi},\theta]\quad\text{and}\quad B_{\Lambda}\operatorname{Var}[\theta\mid\sigma^{\pi}]=O.$ (17)
$A_{\Lambda}\operatorname{Var}[\sigma^{\pi}\mid\theta]=O\quad\text{and}\quad A_{\Lambda}\operatorname{Cov}[\sigma^{\pi},\theta]=B_{\Lambda}Z.$ (18)
Definition 3.
A dual feasible matrix $\Lambda\in\mathcal{S}^{n}_{\rm diag}$ certifies $\pi\in\Pi^{\rm g}$ if $(\pi,\Lambda)$ satisfies (16).
Proposition 3 characterizes an optimal information structure based on its certifiability by Lagrangian multipliers, defined through two matrix equations in (16). These equations can be understood as generalizing the familiar complementary slackness conditions for linear programming. Moreover, in the presence of either equation in (16), the other can be equivalently stated in the form that involves the “residual variance matrix” of an equilibrium action profile or the payoff state, which result in the two equivalent sets of optimality conditions, (17) and (18). As will be shortly discussed, these alternatives elucidate the statistical configuration of optimal information structures.
The optimality condition (16) is a restatement of strong duality. Thus, any dual feasible $\Lambda$ that certifies some optimal $\pi$ is a solution to the dual problem, and vice versa. This further implies that certifiability satisfies universality; that is, if $\Lambda$ certifies some optimal information structure, then it certifies every optimal information structure. This property will be useful when discussing the uniqueness of an optimal information structure.
Corollary 4.
The following are equivalent.
  1. i).
    $\Lambda$ is dual optimal.
  2. ii).
    $\Lambda$ certifies some optimal $\pi\in\Pi^{\rm g}$.
  3. iii).
    $\Lambda$ certifies every optimal $\pi\in\Pi^{\rm g}$.
The strong duality immediately implies that the designer’s optimal value is represented as a linear function of the state variance matrix $Z$, which is the common feature of the values of no disclosure and full disclosure, as discussed in Section 3.3. Moreover, Proposition 3 yields the following equivalent representations of $v^{\rm p}$.
Corollary 5.
Let $\pi\in\Pi^{\rm g}$ be an optimal information structure, certified by $\Lambda\in\mathcal{S}^{n}_{\rm diag}$. Then,
$v^{\rm p}$ $=C_{\Lambda}\bullet Z$
$=A_{\Lambda}\bullet\operatorname{Var}\quantity[\sigma^{\pi}]=B_{\Lambda}\bullet\operatorname{Cov}\quantity[\sigma^{\pi},\theta]$
$=A_{\Lambda}\bullet\operatorname{Var}\quantity[\operatorname{E}\quantity[\sigma^{\pi}\mid\theta]]=C_{\Lambda}\bullet\operatorname{Var}\quantity[\operatorname{E}\quantity[\theta\mid\sigma^{\pi}]].$
The value $v^{\rm p}$ is expressed as the linear function of the action variance matrix $\operatorname{Var}\quantity[\sigma^{\pi}]$ with a coefficient $A_{\Lambda}$, or that of the action-state covariance matrix $\operatorname{Cov}\quantity[\sigma^{\pi},\theta]$ with a coefficient $B_{\Lambda}$. Further, there are two more equivalent ways of representing $v^{\rm p}$. To interpret these representations, we decompose $\operatorname{Var}\quantity[\operatorname{E}\quantity[\sigma^{\pi}\mid\theta]]$ by using the law of total variance,
$\operatorname{Var}\quantity[\operatorname{E}\quantity[\sigma^{\pi}\mid\theta]]=\underbrace{\operatorname{Var}\quantity[\sigma^{\pi}]}_{\begin{subarray}{c}\text{ex-ante}\\ \text{action variance}\end{subarray}}-\ \underbrace{\operatorname{Var}\quantity[\sigma^{\pi}\mid\theta]}_{\begin{subarray}{c}\text{conditional}\\ \text{action variance}\end{subarray}}.$
So, the matrix $\operatorname{Var}\quantity[\operatorname{E}\quantity[\sigma^{\pi}\mid\theta]]$ is statistically interpreted as quantifying the action variance reduction, i.e., the amount of the variance in equilibrium strategies $\sigma^{\pi}$ that can be reduced by conditioning on the true state $\theta$. Likewise, we can interpret $\operatorname{Var}\quantity[\operatorname{E}\quantity[\theta\mid\sigma^{\pi}]]$ as capturing the state variance reduction. Corollary 5 shows that $v^{\rm p}$ can be equally expressed as the linear function of these matrices with coefficients $A_{\Lambda}$ and $C_{\Lambda}$, respectively. Since the action variance reduction reflects the random variation of actions attributed solely to the random variation of the state $\theta$, the representation indicates the absence of extraneous randomness in the optimal value beyond $\theta$. At the same time, the other representation indicates that $v^{\rm p}$ is determined by the state variance reduction conditioning on the entire profile of equilibrium actions, which quantifies the total amount of the information on the state disclosed to the population of agents.
Next, we exploit the optimality conditions in Proposition 3 to delve into the statistical configuration of an optimal information structure $\pi$ along with certifying Lagrangian multipliers $\Lambda$. For mathematical simplicity, we assume that $Z$ has full rank in stating subsequent results. This assumption can be attained in an innocuous fashion if we are allowed to remove any redundant entry of $\theta$ to obtain a minimal representation of the model. [^17] Even without the invertibility of $Z$, many results can be extended by considering the pseudoinverse $Z^{+}$. General results are presented and proved in the appendix whenever such extensions are available. The next lemma shows that when $Z$ has full rank, the dual feasibility of optimal $\Lambda$ simplifies to a lower-dimensional condition that $A_{\Lambda}\succeq O$ than $\mathbf{M}_{\Lambda}\succeq O$.
Lemma 3.
Suppose $Z\succ O$. If $\Lambda\in\mathcal{S}^{n}_{\rm diag}$ satisfies (16) with some primal feasible $(X,Y)$, then $\mathbf{M}_{\Lambda}\succeq O$ if and only if $A_{\Lambda}\succeq O$.
Recall that when an information structure $\pi$ is Gaussian, the associated BNE strategy profile $\sigma^{\pi}$ is jointly normally distributed with $\theta$. The next corollary shows that when $\pi$ is optimal, the action-state joint distribution degenerates so that $\sigma^{\pi}$ and $\theta$ exhibit a linear relationship. Incidentally, the matrices $A_{\Lambda}$ and $B_{\Lambda}$ in the dual problem serve as the relevant coefficients of this linear relationship.
Corollary 6.
Suppose $Z\succ O$. If $\pi\in\Pi^{\rm g}$ is optimal and certified by $\Lambda\in\mathcal{S}^{n}_{\rm diag}$, then
$A_{\Lambda}\quantity(\sigma^{\pi}-\bar{a})=B_{\Lambda}\quantity(\theta-\bar{\theta})\quad{\rm a.s.}$ (19)
Conversely, if any $\pi\in\Pi^{\rm g}$ and dual feasible $\Lambda\in\mathcal{S}^{n}_{\rm diag}$ satisfy (19), then $\pi$ is optimal and certified by $\Lambda$.
This linear relationship implies that any weighted aggregate $\sum_{i=1}^{n}\alpha_{i}\sigma^{\pi}_{i}$ of equilibrium actions is precisely determined as a function of the state $\theta$ when the weights $(\alpha_{1},\ldots,\alpha_{n})$ are taken from the range of $A_{\Lambda}$. To see this, observe that (19) implies
$A_{\Lambda}\operatorname{E}\quantity[\sigma^{\pi}-\bar{a}\mid\theta]=B_{\Lambda}\quantity(\theta-\bar{\theta})\quad{\rm a.s.}$
Thus, if $\alpha=A_{\Lambda}^{\top}\mathbf{x}$ for some $\mathbf{x}\in\mathbb{R}^{n}$, then $\alpha^{\top}\sigma^{\pi}$ can be written as a function of $\theta$. By similar arguments, one can also see that the linear statistics of the state, $\sum_{k=1}^{m}\beta_{k}\theta_{k}$, is pinned down by the equilibrium strategy profile whenever the vector $\beta=(\beta_{1},\ldots,\beta_{m})$ belongs to the range of $B_{\Lambda}^{\top}$. In particular, Corollary 6 yields important implications for the full identification of $\theta$ from $\sigma^{\pi}$, or vice versa, when $A_{\Lambda}$ or $B_{\Lambda}$ has full rank. Specifically, $\sigma^{\pi}$ uniquely determines $\theta$ through (19) if and only if $B_{\Lambda}$ admits a left inverse (which occurs if and only if $B_{\Lambda}$ has column full rank). Similarly, $\theta$ uniquely determines $\sigma^{\pi}$ through the same equation if and only if $A_{\Lambda}$ is non-singular. These properties of an information structure will be explored in the next subsection.
4.2 No disclosure, (aggregative) full disclosure, and noise-freeness
Our duality results can be utilized to establish necessary and sufficient conditions ensuring the optimality of a given information structure. For example, recall that no disclosure $\underline{\pi}$ results in the action variance matrix $X=O$ and the action-state covariance matrix $Y=O$. Substituting these into (16), the first equation is vacuously satisfied, and the second reduces to $B_{\Lambda}Z=O$. In particular, when $Z$ has full rank, this simplifies to $B_{\Lambda}=O$; namely, for each agent $i$, there is a constant ratio $-\lambda_{i}$ such that $w_{ik}=-\lambda_{i}r_{ik}$ for each $k$, whereas the choice of $\lambda_{1},\ldots,\lambda_{n}$ is subject to the dual feasibility constraint. This requires that each $i$-th row of $W$ is proportional to the corresponding $i$-th row of $R$. Likewise, the optimality of full disclosure $\overline{\pi}$ can be characterized by specifying $X$ and $Y$ as $Q^{-1}RZ(Q^{-1}R)^{\top}$ and $Q^{-1}RZ$, respectively. The next result summarizes these.
Corollary 7.
Suppose $Z\succ O$.
  1. i).
    $\underline{\pi}$ is optimal if and only if there exists $\Lambda\in\mathcal{S}^{n}_{\rm diag}$ such that $A_{\Lambda}\succeq O$ and $B_{\Lambda}=O$.
  2. ii).
    $\overline{\pi}$ is optimal if and only if there exists $\Lambda\in\mathcal{S}^{n}_{\rm diag}$ such that $A_{\Lambda}\succeq O$ and $A_{\Lambda}Q^{-1}R=B_{\Lambda}$.
Both $\underline{\pi}$ and $\overline{\pi}$ have the notable characteristic that the designer does not use randomization beyond the randomness in $\theta$, since $\sigma^{\pi}$ is deterministic conditional on the true state. In general, an information structure $\pi$ has this characteristic if its BNE $\sigma^{\pi}$ is represented as a function of $\theta$, that is, the conditional variance of actions given $\theta$ is zero. If this is the case, we refer to $\pi$ as noise-free.
Definition 4.
We say that an information structure $\pi$ is noise-free if $\operatorname{Var}[\sigma^{\pi}\mid\theta]=O$ a.s.
Note that a Gaussian information structure $\pi$ is noise-free if and only if $X=YZ^{+}Y^{\top}$ holds for $X=\operatorname{Var}\quantity[\sigma^{\pi}]$ and $Y=\operatorname{Cov}\quantity[\sigma^{\pi},\theta]$ by the conditional covariance formula. As a result, the first equation in (18) is automatically satisfied regardless of $\Lambda$, and thus the optimality of such an information structure is determined solely by
$A_{\Lambda}Y=B_{\Lambda}Z.$
In other words, the noise-free information structure $\pi$ is optimal if and only if the above equation holds with some dual feasible $\Lambda$.
Leveraging Corollary 6, we can also derive a condition under which “any” optimal information structure is noise-free. Specifically, the linear relationship (19) implies if $A_{\Lambda}$ is invertible, then $\sigma^{\pi}$ is written as a function of $\theta$, meaning that the equilibrium action profile can be identified from the true state. Thus, the optimal information structure is necessarily noise-free whenever certifying Lagrangian multipliers can be chosen in such a way of making $A_{\Lambda}\succ O$. Moreover, in this case, the universality of certification (Corollary 4) implies that the optimal information structure is essentially unique. [^18] We do not require $Z\succ O$ for this result and subsequent ones in the rest of this section. Consequently, these results are applicable for our network-game analysis in Section 5, where the non-singularity of $Z$ may conflict with modeling assumptions.
Corollary 8.
If (4.1) is solved by $\Lambda\in\mathcal{S}^{n}_{\rm diag}$ such that $A_{\Lambda}\succ O$, then any optimal $\pi\in\Pi^{\rm g}$ is noise-free. Moreover, in this case, the solution to (P) is unique.
Next, we consider another yet related property of information structures that extends the idea of full disclosure $\overline{\pi}$. Consider a scenario where a hypothetical observer can access all the agents’ action recommendations forwarded to each of them. Then, we would interpret $\pi$ as disclosing the true state to the entire “population” when the residual variance of $\theta$ is zero conditional on the equilibrium strategy profile $\sigma^{\pi}$. In essence, under such $\pi$, even though the true state may remain unknown to each individual, the information structure reveals it to the entire population in an aggregative form, reflecting the designer’s maximum use of knowledge about the state to guide the agents toward the designer-optimal outcome. This characteristic of information structures can be formalized by means of the zero residual variance of the state.
Definition 5.
We say that an information structure $\pi\in\Pi$ is aggregatively disclosing if $\operatorname{Var}\quantity[\theta\mid\sigma^{\pi}]=O$ a.s.
As is evident from the definition, the notion of aggregative disclosure serves as a counterpart to noise-freeness, where the roles of $\sigma^{\pi}$ and $\theta$ are exchanged. Again, a Gaussian information structure $\pi$ is aggregatively disclosing if and only if $Z=Y^{\top}X^{+}Y$ holds with $X=\operatorname{Var}\quantity[\sigma^{\pi}]$ and $Y=\operatorname{Cov}\quantity[\sigma^{\pi},\theta]$. In this case, the second equation in (17) is automatically satisfied regardless of $\Lambda$, and thus the optimality of such an information structure is determined solely by
$A_{\Lambda}X=B_{\Lambda}Y^{\top}.$
In other words, the aggregatively disclosing information structure $\pi$ is optimal if and only if the above equation holds with some dual feasible $\Lambda$.
Similarly to before, Corollary 6 can provide a condition under which any optimal information structure is aggregatively disclosing. Specifically, (19) implies that $\theta$ can be expressed as a function of $\sigma^{\pi}$ when $B_{\Lambda}$ is a left-invertible matrix, which is the case if and only if $B_{\Lambda}$ has column full rank (note that this necessitates $m\leq n$). Thus, whenever certifying Lagrangian multipliers can be constructed in this way, any optimal information structure is aggregatively disclosing.
Corollary 9.
[…]
Now, let us revisit the two extreme disclosure policies in Corollary 7. It is worth mentioning that the optimality of no disclosure $\underline{\pi}$ or full disclosure $\overline{\pi}$ crucially hinges on the dimension of the payoff state. Specifically, if the state $\theta$ is multi-dimensional, i.e., $Z\succ O$ and $m\geq 2$, then both $\underline{\pi}$ and $\overline{\pi}$ are generically suboptimal. This is because each of the optimality conditions, $B_{\Lambda}=O$ and $A_{\Lambda}Q^{-1}R=B_{\Lambda}$, in Corollary 7 consists of $nm$ equations, while $\Lambda$ contains only $n$ variables. Thus, the system of equations generically has no solution, as the number of equations exceeds the number of variables.
In contrast, when the state is one-dimensional, Corollary 7 presents qualitatively different implications regarding the optimality of no disclosure. Specifically, there generally exists a unique set of Lagrangian multipliers $\lambda_{1},\ldots,\lambda_{n}$ that solves $B_{\Lambda}=O$ as long as each entry of $R$ is nonzero. Constructing $\Lambda$ in this way, it follows that $\underline{\pi}$ is optimal if and only if $A_{\Lambda}\succeq O$. Importantly, this condition is met for generic payoff parameters, unlike the generic suboptimality of $\underline{\pi}$ in the multi-dimensional case. More interestingly, whenever no disclosure is suboptimal, it is shown that an optimal information structure must be aggregatively disclosing.
Corollary 10.
Suppose $m=1$ and $r_{i}\neq 0$ for all $i$, and let $D=\operatorname{Diag}(w_{1}/r_{1},\ldots,w_{n}/r_{n})$ and $\tilde{V}=V+(DQ+Q^{\top}D)/2$, where $r_{i}$ and $w_{i}$ denote the $i$-th entries of $n$-dimensional vectors $R$ and $W$, respectively.
  1. i).
    No disclosure is optimal if and only if $O\succeq\tilde{V}$.
  2. ii).
    Otherwise, if $O\not\succeq\tilde{V}$, any optimal information structure is aggregatively disclosing.
5 Symmetry in network games
Now, we focus on the model of network games, which extends the leading example from Section 2. In this model, each agent $i$’s best-response strategy is given by linear the equation
$\sigma^{\pi}_{i}+q\sum_{j\neq i}g_{ij}\operatorname{E}_{i}\quantity[\sigma^{\pi}_{j}]=\operatorname{E}_{i}\quantity[\hat{\theta}_{i}],\quad\text{where}\quad\hat{\theta}_{i}\equiv\sum_{k=1}^{m}r_{ik}\theta_{k}.$ (20)
In words, the agent $i$’s optimal strategy $\sigma_{i}^{\pi}$ is chosen in such a way of keeping the localized aggregate of her own and neighbors’ strategies equal to her personal state $\hat{\theta}_{i}$ that is defined as the weighted aggregate of multi-dimensional random variables $(\theta_{1},\ldots,\theta_{n})$. In this section, we treat these personal states $\hat{\theta}=(\hat{\theta}_{1},\ldots,\hat{\theta}_{n})$, rather than $(\theta_{1},\ldots,\theta_{m})$, as the primitive of the model, and interpret $Z$ as $\operatorname{Var}[\hat{\theta}]$. In reference to (20), this treatment allows us to reduce the modeling parameters to let $n=m$ and $R=I$. Thus, this section studies the following class of LQG games.
Definition 6.
Let $G=[g_{ij}]_{n\times n}$ be a binary matrix with $g_{ii}=0$ for every $i\in[n]$. A $G$-network game is a basic game $(Q,R,Z)$ such that $n=m$, $Z=\operatorname{Var}[\hat{\theta}]\neq O$ is an $n\times n$ positive semidefinite matrix, and $Q$ and $R$ are given by
$Q=Q_{G}\equiv I+qG\quad\text{and}\quad R=I,$
where $q\in\mathbb{R}$ is a parameter such that $Q_{G}+Q_{G}^{\top}\succ O$ [^19] The condition $Q_{G}+Q_{G}^{\top}\succ O$ holds if $|q|$ is less than the reciprocal of the spectral radius of $(G+G^{\top})/2$.
A few remarks are in order. First, note that $Z=\operatorname{Var}[\hat{\theta}]$ is any $n\times n$ is generally not invertible, even if the original payoff state $\theta=(\theta_{1},\ldots,\theta_{m})$ entails an invertible variance matrix. For example, if all agents exhibit the same response to the original $\theta$, the rank of $Z$ is necessarily one. Second, in consistency with the formulation (3.3) of the designer’s objective function, the reduction to personal states necessitates that $v$ is dependent on $\theta$ only through $\hat{\theta}$. This requirement is satisfied, for example, if the designer’s objective is given as a function of agents’ payoffs.
5.1 Optimality of symmetric information structures
A network $G$ describes the physical links between agents. Analogously, we can interpret the action covariance matrix $\operatorname{Var}[\sigma^{\pi}]$ as a representation of the informational links between agents. In this section, we show that when the network has some symmetry, the informational links under optimal information structures exhibit similar symmetry, provided that the designer’s objective function does not conflict with the connectivity structure in agents’ payoffs. This result serves as a building block for studying the relationship between optimal informational links and physical links in network games. Furthermore, the symmetric structure facilitates the computation of optimal information structures by allowing us to equalize Lagrangian multipliers across symmetric agents.
[…]
$\operatorname{Orb}\quantity(i\mid G)\equiv\quantity{j\in[n]:\tau(i)=j\text{ for some }\tau\in\operatorname{Aut}(G)},$
which is the set of agents who are reachable from agent $i$ by applying some automorphism $\tau\in\operatorname{Aut}(G)$. One can easily see that each orbit constitutes an equivalence class in $[n]$. An automorphism group $\operatorname{Aut}(G)$ is said to be transitive if $\operatorname{Orb}\quantity(i\mid G)=[n]$ holds for all $i$ (equivalently, for some $i$).
When $j\in\operatorname{Orb}\quantity(i\mid G)$, two agents $i$ and $j$ can be considered to play symmetric roles in $G$, as swapping their indices leaves the network unchanged. In general, the coarseness of the orbits as a partition of $[n]$ is determined by the richness of $\operatorname{Aut}\quantity(G)$, which captures the degree of symmetry that $G$ entails. Specifically, if there is no automorphism other than the identity mapping, then $\operatorname{Orb}\quantity(i\mid G)=\{i\}$ so that the network holds no symmetric pair of agents. On the other extreme, for the complete network $K=[k_{ij}]_{n\times n}$ with $k_{ij}=1$ for every $i\neq j$, $\operatorname{Aut}\quantity(K)$ equals to the collection of all permutations, implying the symmetric roles of all agents. Aside from the complete network, each symmetric network considered in Section 2 also entails a transitive automorphism group.
Given any permutation $\tau:[n]\to[n]$, we define the associated permutation matrix $P_{\tau}\equiv[\mathbf{e}_{\tau(1)}\cdots\mathbf{e}_{\tau(n)}]$, whose $i$-th column is given as the $\tau(i)$-th unit vector. For each $n\times n$ matrix $M$, we abuse the notation to write $\tau(M)\equiv P_{\tau}MP_{\tau}^{\top}$, which is the matrix obtained by permuting the rows and columns of $M$ according to $\tau$. We say that $M$ is invariant with respect to $G$, or $G$-invariant in short, when $M=\tau(M)$ holds for all $\tau\in\operatorname{Aut}(G)$. For example, $M$ is $G$-invariant for any network $G$ if (and only if) its diagonal entries are identical, and its off-diagonal entries are identical as well.
Note that any network $G$ itself is a $G$-invariant matrix. Consequently, in the $G$-network game, agents’ payoff parameters entail $G$-invariance; that is $Q=Q_{G}$ and $R=I$ are $G$-invariant matrices. The next assumption further postulates that the designer’s objective function and the distribution of personal states are aligned with the symmetry present in the network $G$.
Assumption 3.
Given a network $G$, the matrices $V$, $W$, and $Z$ are $G$-invariant.
In addition, we define the notion of $G$-invariance for an information structure based on the $G$-invariance of the action-state variance matrix that the information structure induces.
Definition 7.
An information structure $\pi$ is $G$-invariant if $\operatorname{Var}[\sigma^{\pi}]$ and $\operatorname{Cov}[\sigma^{\pi},\theta]$ are $G$-invariant matrices.
[…]
Proposition 11.
Under Assumption 3, there exists an information structure $\pi\in\Pi^{\rm g}$ that is $G$-invariant and optimal. Moreover, $\pi$ is certified by $\Lambda=\operatorname{Diag}(\lambda_{1},\ldots,\lambda_{n})$ such that $\lambda_{i}=\lambda_{j}$ whenever $j\in\operatorname{Orb}\quantity(i\mid G)$. In particular, $\lambda_{1}=\cdots=\lambda_{n}$ if $\operatorname{Aut}\quantity(G)$ is transitive.
This proposition establishes that if LQG information design exhibits $G$-invariance, there exists a $G$-invariant optimal information structure, and it can be certified by some $G$-invariant diagonal matrix. In particular, when the underlying network $G$ entails a transitive automorphism group, the certifying Lagrangian multipliers can be uniform across all agents. This finding is useful to reduce the dimension of an SDP problem, thereby facilitating both general analysis and explicit calculation of optimal information structures in symmetric environments.
Now, we examine some qualitative properties of optimal symmetric information structures. The next result offers a sufficient condition under which an optimal information structure is noise-free when $\operatorname{Aut}\quantity(G)$ is transitive and $Z$ is non-singular.
Corollary 12.
In addition to Assumption 3, suppose that $\operatorname{Aut}\quantity(G)$ is transitive and $Z\succ O$. If there exists no constant $\delta\in\mathbb{R}$ such that (i) both $\delta I+W$ and $\delta/2\cdot(Q_{G}+Q_{G}^{\top})-V$ are singular, and (ii) $\delta/2\cdot(Q_{G}+Q_{G}^{\top})\succeq V$, then any optimal information structure is noise-free, the solution to (P) is unique, and (4.1) is solved by some scalar matrix $\Lambda=\lambda I$ with $A_{\Lambda}\succ O$.
To evaluate this result, suppose that $\delta$ as in Corollary 12 exists; namely, $\delta$ must be a common eigenvalue of two matrices,
$-W\quad{\rm and}\quad\quantity(\frac{Q_{G}+Q_{G}^{\top}}{2})^{-1/2}V\quantity(\frac{Q_{G}+Q_{G}^{\top}}{2})^{-1/2},$
and, particularly, $\delta$ serves as the maximal eigenvalue of the second matrix. Now, if we perturb $W$ by adding a scalar matrix $\epsilon I$, the eigenvalues of the first matrix uniformly decrease by $\epsilon$, while keeping the eigenvalues of the second matrix unchanged. This indicates that for generic choices of parameters, an optimal information structure is noise-free and essentially unique when $\operatorname{Aut}\quantity(G)$ is transitive and $Z$ is non-singular.
In $G$-network games, each personal state $\hat{\theta}_{i}$ is exclusively relevant to agent $i$’s payoff and has no impact on the payoffs of others. In this setting, it is natural to postulate that the designer’s objective does not depend on the cross term of $\hat{\theta}_{i}$ and $a_{j}$ for $i\neq j$. This means that $W$ is a diagonal matrix, and particularly, it is a scalar matrix of the form $W=wI$ if $\operatorname{Aut}\quantity(G)$ is transitive. The next corollary identifies key qualitative features of optimal information structures in this case by establishing a bang-bang property akin to Corollary 10. Moreover, it refines the noise-freeness condition in Corollary 12 and highlights that noise-freeness is generically attained at the optimum as long as agents’ personal states are not perfectly correlated.
Corollary 13.
In addition to Assumption 3, suppose that $\operatorname{Aut}\quantity(G)$ is transitive and $W=wI$ for some $w\in\mathbb{R}$. Let $\tilde{V}=V+w/2\cdot(Q_{G}+Q_{G}^{\top})$.
  1. i).
    No disclosure is optimal if and only if $O\succeq\tilde{V}$.
  2. ii).
    Otherwise, if $O\not\succeq\tilde{V}$, any optimal information structure is aggregatively disclosing.
If, in addition, $Z\succ O$ and $\tilde{V}$ is non-singular, then any optimal information structure is noise-free, and the solution to (P) is unique.
This corollary accounts for our numerical findings in Section 2. Specifically, since the designer’s objective is given as the sum of individual action variances, the leading example maps to the LQG information design with $V=I$ and $W=O$. As a result, $\tilde{V}$ equals $I$—which is positive definite—implying, by Corollary 13, that an optimal information structure is aggregatively disclosing, irrespective of agents’ payoffs. Notably, this result holds regardless of the distribution of agents’ personal states. On top of that, in any private-value case, $Z$ is non-singular, and thus the optimal information structure is noise-free and essentially unique.
5.2 Optimal information structures on the complete network
To illustrate our findings, we now focus on $K$-network games, which correspond to the special case with $G$ equal to the complete network $K$. Specifically, this section studies the following parametrized model.
Assumption 4.
Suppose $n\geq 2$.
  • The agents’ payoff parameters are given by
    $Q=\matrixquantity[1&q&\cdots&q\\ q&1&\cdots&q\\ \vdots&\vdots&\ddots&\vdots\\ q&q&\cdots&1]\quad{\rm and}\quad R=I,\quad{\rm where}\quad-\frac{1}{n-1}<q<1.$
  • The designer’s payoff parameters are given by
    $V=\matrixquantity[v&c&\cdots&c\\ c&v&\cdots&c\\ \vdots&\vdots&\ddots&\vdots\\ c&c&\cdots&v]\quad{\rm and}\quad W=wI.$
  • The variance matrix of personal states is given by
    $Z=\matrixquantity[z&\rho_{z}z&\cdots&\rho_{z}z\\ \rho_{z}z&z&\cdots&\rho_{z}z\\ \vdots&\vdots&\ddots&\vdots\\ \rho_{z}z&\rho_{z}z&\cdots&z],\quad{\rm where}\quad z>0\quad{\rm and}\quad-\frac{1}{n-1}<\rho_{z}\leq 1.$
The parametric restrictions on $(z,\rho_{z})$ ensures that $Z\succeq O$, while excluding the trivial case of $\operatorname{Var}[\hat{\theta}_{1}]=\cdots=\operatorname{Var}[\hat{\theta}_{n}]=0$. Note that $\rho_{z}$ is allowed to take $1$, in which case agents’ personal states are perfectly correlated with one another. This case is called the common-value case. The other case of $\rho_{z}<1$ is called the private-value case.
According to Corollary 13, an optimal information structure reveals either nothing or everything to the entire set of agents, depending on the sign of $\tilde{V}\equiv V+wQ$. This matrix $\tilde{V}$ holds constant diagonal entires, as well as constant off-diagonal entries, which are given by
$\alpha\equiv v+w\quad{\rm and}\quad\beta\equiv c+qw,$
respectively. These parameters $\alpha$ and $\beta$ serve as sufficient determinants for which of no disclosure or aggregative disclosure is optimal since $\tilde{V}$ is negative semidefinite if and only if
$\alpha\leq\beta\leq-\frac{\alpha}{n-1}.$ (21)
Specifically, no disclosure $\underline{\pi}$ is optimal if and only if (21) is satisfied. This condition necessitates $\alpha\leq 0$, which reflects the designer’s preference for minimizing the volatility of agents’ individual actions. In addition, $\beta$ is required to have a small absolute value relative to $\alpha$, meaning that the designer receives a small marginal value from the correlation between different agents’ actions.
In contrast, if (21) is violated, any optimal information structure is aggregatively disclosing. While it remains ambiguous how much information is revealed to each individual agent, Proposition 11 implies that $\operatorname{Var}[\sigma^{\pi}]$ and $\operatorname{Cov}[\sigma^{\pi},\theta]$ can be $K$-invariant at the optimum, leading us to postulate that a primal solution $(X,Y)$ to (P) takes the form,
$X=\matrixquantity[x&\rho_{x}x&\cdots&\rho_{x}x\\ \rho_{x}x&x&\cdots&\rho_{x}x\\ \vdots&\vdots&\ddots&\vdots\\ \rho_{x}x&\rho_{x}x&\cdots&x]\quad{\rm and}\quad Y=\matrixquantity[y&\rho_{y}y&\cdots&\rho_{y}y\\ \rho_{y}y&y&\cdots&\rho_{y}y\\ \vdots&\vdots&\ddots&\vdots\\ \rho_{y}y&\rho_{y}y&\cdots&y],$ (22)
which are parametrized by four variables $(x,y,\rho_{x},\rho_{y})$. To ensure that $(X,Y)$ is primal feasible, these variables need to satisfy
$y=x+(n-1)q\cdot\rho_{x}x.$ (23)
Proposition 11 also allows us to let certifying Lagrangian multipliers be uniform across all agents, $\lambda\equiv\lambda_{1}=\cdots=\lambda_{n}$. Then, we can analytically characterize the optimal information structure by solving for five variables $(x,y,\rho_{x},\rho_{y},\lambda)$ by using one of the optimality conditions in Proposition 3, coupled with the primal feasibility constraint (23). In what follows, we perform detailed calculations by distinguishing between two scenarios regarding the distribution of personal states.
5.2.1 Common-value case
We first consider the common-value case, i.e., $\rho_{z}=1$. In this case, $\hat{\theta}_{0}\equiv\hat{\theta}_{1}=\cdots=\hat{\theta}_{n}$ occurs with probability $1$, where $\hat{\theta}_{0}$ is called the common state, and we necessarily have that $\rho_{y}=1$. Then, by solving for the remaining variables, the next result offers a closed-form expression for an optimal information structure in the common value case.
Corollary 14.
In addition to Assumption 4, suppose that $\rho_{z}=1$.
  1. i).
    No disclosure $\underline{\pi}$ is optimal if and only if (21) holds.
  2. ii).
    Full disclosure $\overline{\pi}$ is optimal if and only if
    $\beta\geq-\frac{\alpha}{n-1}\quad{\rm and}\quad\beta\geq\underbrace{\quantity(\frac{(n+1)q-1}{2n-1-(n-1)q})}_{\equiv\ b(q)}\alpha,$ (24)
    where $b(q)$ satisfies $-1/n<b(q)<q$.
  3. iii).
    Otherwise, an optimal information structure $\pi$ is partial disclosure. The action-state joint distribution under $\pi$ is characterized by
    $x=\frac{(\alpha-\beta)(q\alpha-(2-q)\beta)z}{4n(1-q)(q\alpha-\beta)^{2}},\quad y=\frac{(\alpha-\beta)z}{2n(q\alpha-\beta)},\quad\rho_{x}=\frac{(1-2q)\alpha+\beta}{(n-1)(q\alpha-(2-q)\beta)}.$
    In particular, each agent’s action recommendation reduces the common state variance as follows:
    $\frac{z-\operatorname{Var}\quantity[\hat{\theta}_{0}\mid\sigma^{\pi}_{i}]}{z}=\frac{(1-q)(\alpha-\beta)}{n(q\alpha-(2-q)\beta)}\equiv P_{1}(\alpha,\beta).$
Figure 3: In the common-value case ($\rho_{z}=1$), an optimal information structure is either no disclosure (white), full disclosure (dark red), or partial disclosure (yellow-red gradient), depending on the values of $\alpha$ and $\beta$. In the gradient region, the color range expresses the value of $P_{1}(\alpha,\beta)$. The diagram is obtained by setting $n=3$ and $q=1/3$.
The implications of Corollary 14 can be visualized using a color map, as shown in Figure 3, with $\alpha$ and $\beta$ on the horizontal and vertical axes, respectively. The diagram is partitioned into three regions based on the linear relationships between $\alpha$ and $\beta$ that are identified in the corollary. The white region, which is determined by the condition (21), corresponds to the set of $(\alpha,\beta)$ for which $\underline{\pi}$ is optimal. On the other hand, $\overline{\pi}$ is optimal in the dark red region that is determined by (24). Partial disclosure is optimal in the remaining region, shaded in a gradient from yellow to red.
When partial disclosure is optimal, an optimal information structure is aggregatively disclosing, while the informativeness of each agent’s private signal depends on $\alpha$ and $\beta$. Specifically, the color range represents the value of $P_{1}(\alpha,\beta)$, which is the percentage of the amount in the ex-ante variance of the common state that is reduced by conditioning on each agent’s action recommendation. The color range gets deeper with $P_{1}(\alpha,\beta)$ increasing from lower values to higher values. Isoquant curves are given as rays going through the origin, reflecting the fact that $P_{1}(\alpha,\beta)$ is dependent on $\alpha$ and $\beta$ only through the ratio between them. As the ratio $\beta/\alpha$ gets closer to the threshold of $b(q)$ from $-1$, the informativeness of an individual signal increases. It is worth emphasizing that an optimal information structure in this case is not noise-free. Indeed,
$\operatorname{Var}\quantity[\sigma^{\pi}\mid\hat{\theta}_{0}]=x-\frac{y^{2}}{z}=\frac{(\alpha-\beta)(b(q)\cdot\alpha-\beta)}{4n(1-q)(q\alpha-\beta)^{2}}\cdot z$
is strictly positive if both (21) and (24) are violated. Therefore, in the common-value case, the designer can achieve the optimal form of aggregative disclosure by adding some noise to each agent’s signal, while the optimal information structure is designed in such a way that those noise terms offset with one another in aggregate.
Several papers in the literature have investigated information design problems in symmetric common-value LQG games that are similar to the model of this section. Angeletos and Pavan (2007) and Ui and Yoshizawa (2015) focus on symmetric Gaussian information structures, where the player’s signal consists of a conditionally i.i.d. private signal and a public signal, and characterize the optimal combination of them. Bergemann and Morris (2013) derive an optimal symmetric Gaussian information structure in a homogeneous product Cournot game to maximize the expected total profit by focusing on symmetric Bayes correlated equilibria in which actions and a one-dimensional state are jointly normally distributed. However, it has not been addressed whether these information structures are also optimal among “all” information structures, including asymmetric or non-Gaussian information ones. At this point, Proposition 11 ensures the robust optimality of the symmetric Gaussian information structure in our Corollary 14; that is, it cannot be outperformed even when other alternative disclosure policies are available.
5.2.2 Private-value case
Next, we consider the private-value case, i.e., $\rho_{z}<1$, under which each agent’s personal state contains some idiosyncratic component. Since $Z$ is non-singular in this case, we can apply Corollary 12 to discuss the uniqueness and noise-freeness properties of an optimal information structure. The following result demonstrates that if $(\alpha,\beta)$ does not lie on the boundary of the optimality region for $\underline{\pi}$, these properties are attained at the optimum, and the dual problem is solved by a strictly feasible choice.
Corollary 15.
In addition to Assumption 4, suppose that $\rho_{z}<1$.
  1. i).
    No disclosure $\underline{\pi}$ is optimal if and only if (21) holds.
  2. ii).
    Full disclosure $\overline{\pi}$ is optimal if and only if $\alpha\geq 0$ and $\beta=q\alpha$.
  3. iii).
    Otherwise, partial disclosure is optimal. An optimal information structure is aggregatively disclosing and noise-free.
Moreover, if $(\alpha,\beta)$ does not lie on the boundary of the region defined by (21)—i.e., if either (a) $\alpha>0$or (b) $\beta\neq\alpha$and $\beta\neq-\alpha/(n-1)$ —then the solution to (P) is unique, and (4.1) is solved by a scalar matrix $\Lambda=\lambda I$ with $A_{\Lambda}\succ O$.
Figure 4: In the private-value case ($\rho_{z}<1$), optimal information structures are visualized as a function of $(\alpha,\beta)$. The diagrams are obtained by setting $n=3$, $q=1/3$, and $\rho_{z}=0$.
In Figure 4, we graphically describe an optimal information structure $\pi$ in the private-value case by computing the reduction of agent’s own state variance, as well as that of the opponent’s state variance, defined as follows:
$P_{0}^{\rm own}\equiv\frac{z-\operatorname{Var}\quantity[\hat{\theta}_{i}\mid\sigma^{\pi}_{i}]}{z}\quad{\rm and}\quad P_{0}^{\rm opp}\equiv\frac{z-\operatorname{Var}\quantity[\hat{\theta}_{j}\mid\sigma^{\pi}_{i}]}{z},\quad i\neq j.$
The left (resp. right) panel depicts the value of $P_{0}^{\rm own}$ (resp. $P_{0}^{\rm opp}$), where the color range is interpreted in the same manner as in Figure 3.
[…]
Combined with our general strong duality results, Corollary 15 uncovers the essential features of an optimal information structure in the private-value case, beyond the situation where $\underline{\pi}$ or $\overline{\pi}$ is optimal. Below, in steps, we demonstrate how to calculate $(x,y,\rho_{x},\rho_{y})$ under the optimal information structure. The calculation signifies the usefulness of our optimality conditions by reducing the designer’s optimization problem to a systematic procedure for solving polynomial equations. For simplicity, we focus on the pure private-value case, i.e., $\rho_{z}=0$, but the same approach can be applied to other cases.
First, by Proposition 11, it is without loss of generality to let $\Lambda=\lambda I$ be a scalar matrix in the dual problem. Then, we set up auxiliary matrices,
$A_{\Lambda}=\lambda Q-(V+wQ)\quad{\rm and}\quad B_{\Lambda}=\frac{\lambda I}{2}.$ (25)
Corollary 15 tells us that $A_{\Lambda}$ must be positive definite at the optimum. In particular, since $A_{\Lambda}$ holds $\lambda-\alpha$ and $q\lambda-\beta$ as its on-diagonal and off-diagonal entries, $A_{\Lambda}\succ O$ if and only if
$\lambda>\max\quantity{\frac{\alpha-\beta}{1-q},\frac{\alpha+(n-1)\beta}{1+(n-1)q}}.$ (26)
Second, we express $(y,\rho_{y})$ in terms of $\lambda$ by using one of the optimality conditions in Proposition 3. By substituting $A_{\Lambda}$ and $B_{\Lambda}$ into the second equation of (16), we obtain
$y=\frac{z}{2}\cdot\frac{\lambda f(\lambda)}{\quantity(\lambda-\alpha)f(\lambda)-(n-1)\quantity(\beta-q\lambda)^{2}}\quad{\rm and}\quad\rho_{y}=\frac{\beta-q\lambda}{f(\lambda)},$ (27)
where
$f(\lambda)=\quantity(1+(n-2)q)\lambda-\quantity(\alpha+(n-2)\beta).$
Third, we characterize $(x,\rho_{x})$ by using the fact that the optimal information structure is noise-free; that is, $X=YZ^{-1}Y$ holds. This implies that $x$ and $\rho_{x}$ can be expressed in terms of $y$ and $\rho_{y}$ as follows:
$x=\frac{y^{2}+(n-1)\rho_{y}^{2}y^{2}}{z}\quad\text{and}\quad\rho_{x}=\frac{2\rho_{y}+(n-2)\rho_{y}^{2}}{1+(n-1)\rho_{y}^{2}}.$ (28)
Fourth, we solve for $\lambda$ by using the primal feasibility constraint (23). Specifically, by substituting (28) into (23), it follows that
$\quantity((n-1)\rho_{y}+1)^{2}y=z.$
We then plug (27) into this equation to obtain a polynomial equation about $\lambda$, whose solution is used as a certifier of the optimal information structure. We remark that the equation can have multiple solutions, but the dual feasibility constraint (26) serves as a criterion to choose a unique solution, say $\lambda^{*}$.
Lastly, by substituting $\lambda^{*}$ back into (27), we can recover $y$ and $\rho_{y}$, which are then substituted into (28) to yield $x$ and $\rho_{x}$. The resulting profile $(x,y,\rho_{x},\rho_{y})$ characterizes the action-state joint distribution under the optimal information structure. Furthermore, by taking the inverse of $A_{\Lambda^{*}}$ with $\Lambda^{*}=\lambda^{*}I$, Corollary 6 recovers the profile of action recommendations as follows:
$\sigma^{\pi}=\bar{a}+\frac{A_{\Lambda^{*}}^{-1}\quantity(\hat{\theta}-\operatorname{E}[\hat{\theta}])}{2}.$ (29)
Now, we discuss some pivotal properties of the optimal information structure in the private-value case. As we can see from (29), each agent’s action recommendation is given as a linear transformation of payoff-relevant random variables, including both the agent’s own personal state and those of her opponents. Also, using the implication (28) of noise-freeness, we can calculate $P_{0}^{\rm own}$ and $P_{0}^{\rm opp}$ as follows:
$P_{0}^{\rm own}=\frac{1}{1+(n-1)\rho_{y}^{2}}\quad{\rm and}\quad P_{0}^{\rm opp}=\frac{\rho_{y}^{2}}{1+(n-1)\rho_{y}^{2}}.$
These quantities depend on the optimal information structure only through $\rho_{y}^{2}$, with $P_{0}^{\rm own}$ decreasing and $P_{0}^{\rm opp}$ increasing in $\rho_{y}^{2}$. Moreover, it holds that
$P_{0}^{\rm own}+(n-1)P_{0}^{\rm opp}=1,$
which indicates that the average of $\operatorname{Var}[\hat{\theta}_{1}\mid\sigma^{\pi}_{i}],\ldots,\operatorname{Var}[\hat{\theta}_{n}\mid\sigma^{\pi}_{i}]$ is kept constant.
The above discussion indicates that as long as the optimal information structure is not $\underline{\pi}$, each agent’s action recommendation typically mixes up her own payoff state and the opponents’ payoff states, yet the entire profile of action recommendations is designed to fully eliminate fundamental uncertainty in aggregate, as being signified by Corollary 13. This method of implementing aggregative disclosure is qualitatively different from the common-value case, where an action recommendation is generated as a perturbation of the common state by some noise, with noise terms offsetting in aggregate. In contrast, in the private-value case, all fundamental uncertainty is fully resolved in aggregate without introducing any noise into individual signals. Instead, the designer leverages each agent’s own state as well as those of their opponents to guide their behavior towards the designer’s optimal outcome.
6 Public information structures
In this section, we assume that feasible information structures are limited to public ones, which may be plausible in several real-world contexts. For example, consider a central bank that distributes information to firms and consumers in an economy. In this case, tailoring this information to each individual agent is impractical, not only because of technical constraints, but also because of the central bank’s principles of transparency.
An information structure $\pi$ is said to be public if $t_{0}\equiv t_{1}=\cdots=t_{n}$ occurs with probability one, where $t_{0}$ is called a public signal. Let $\Pi^{\rm pb}\subseteq\Pi^{*}$ denote the collection of all public information structures. We study the optimization problem,
$\max_{\pi\in\Pi^{\rm pb}}\quad\operatorname{E}[v(\sigma^{\pi},\theta)],$
Let $v^{\rm pb}$ denote the value of this problem, referred to as public LQG information design. Henceforth, we assume that $Z\succ O$.
As in the case of private information structures, our first task is to derive the SDP formulation of public LQG information design. A crucial aspect of public information is that agents do not face any strategic uncertainty about the opponents’ actions, which simplify the characterization of a BNE and its induced moments. Specifically, under a public information structure, the first-order condition (7) is
$\sum_{j=1}^{n}q_{ij}\sigma^{\pi}_{j}(t_{0})=\sum_{k=1}^{m}r_{ik}\operatorname{E}\quantity[\theta_{k}\mid t_{0}],\quad\forall i\in N,$
involving no conditional expectations in the left-hand side. From this system of equations, the unique BNE is obtained as $\sigma^{\pi}(t_{0})=Q^{-1}R\operatorname{E}\quantity[\theta\mid t_{0}]$.
Calculating the equilibrium action-state covariance from this closed-form expression for BNE, we can express $\mathbf{M}^{\pi}$ as a function of the state variance reduction, $\operatorname{Var}\quantity[E\quantity[\theta\mid t_{0}]]$, measuring the amount of the state variance that can be reduced by observing the public signal $t_{0}$. This matrix necessitates the statistical requirement of $Z\succeq\operatorname{Var}\quantity[E\quantity[\theta\mid t_{0}]]\succeq O$. Conversely, in the next lemma, we show that any symmetric matrix that satisfies these matrix inequalities can be induced as an equilibrium action-state covariance matrix under some public Gaussian information structure. Moreover, it shows that the state variance reduction serves as a sufficient statistic for determining the designer’s expected value of each public information structure.
Lemma 4.
For any public information structure $\pi\in\Pi^{\rm pb}$, if we write $S=\operatorname{Var}[\operatorname{E}\quantity[\theta\mid t_{0}]]$, then it holds that $Z\succeq S\succeq O$, and $\mathbf{M}^{\pi}$ coincides with
$\mathbf{M}_{S}\equiv\matrixquantity[Q^{-1}RS(Q^{-1}R)^{\top}&Q^{-1}RS\\ S(Q^{-1}R)^{\top}&Z]\in\mathcal{S}_{+}^{n+m}.$
Conversely, for any $S\in\mathcal{S}^{m}$ such that $Z\succeq S\succeq O$, there exists a public Gaussian information structure $\pi\in\Pi^{\rm pb}\cap\Pi^{\rm g}$ such that $\mathbf{M}^{\pi}=\mathbf{M}_{S}$. Moreover, the designer’s expected objective is expressed as
$\operatorname{E}\quantity[v(\sigma^{\pi},\theta)]=\overline{C}\bullet S,$
where $\overline{C}\in\mathcal{S}^{m}$ is the matrix defined in (14).
As similarly to the private information case, these lemmas deliver pivotal implications for simplifying the public LQG information design. First, Lemma 4 implies that the set of inducible covariance matrices under all public information structures coincides with that under all public Gaussian ones, since
$\quantity{\mathbf{M}^{\pi}:\pi\in\Pi^{\rm pb}}=\quantity{\mathbf{M}_{S}:Z\succeq S\succeq O}=\quantity{\mathbf{M}^{\pi}:\pi\in\Pi^{\rm pb}\cap\Pi^{\rm g}}.$
In addition, since the designer’s objective depends on the choice of public information structures only through the state variance reduction, it is without loss of optimality to focus on public Gaussian information structures among all public ones.
Second, the public LQG information design problem can be restated as the maximization of the linear objective function $\overline{C}\bullet S$ subject to the two-sided positive semidefiniteness constraints of the form $Z\succeq S\succeq I$. Moreover, by rescaling, this SDP problem turns out to be equivalent to
$\max_{S\in\mathcal{S}^{m}}\quad\widehat{C}\bullet S\quad\text{subject to}\quad I\succeq S\succeq O,\quad\text{where}\quad\widehat{C}\equiv Z^{\frac{1}{2}}\overline{C}Z^{\frac{1}{2}}.$ (PB)
To see the equivalence, let $\widehat{C}$ be given as above, and let $\widehat{S}=Z^{-1/2}SZ^{-1/2}$. Then, it should be easily confirmed that $Z\succeq S\succeq O$ if and only if $I\succeq\widehat{S}\succeq O$, and that $\overline{C}\bullet S=\widehat{C}\bullet\widehat{S}$, from which we can justify (PB) as the equivalent restatement of the original one.
From (PB), we readily see that if $\widehat{C}$ is negative semidefinite, then $\widehat{C}\bullet S\leq 0$ holds for all feasible $S$, thereby no disclosure being optimal. Conversely, if $\widehat{C}$ is positive semidefinite, then the objective is non-decreasing with respect to the matrix ordering $\succeq$, whence it is optimal to choose $S=I$, which corresponds to full disclosure. In particular, when $m=1$, the matrix $\widehat{C}$ reduces to a scalar variable, and its sign is unambiguously determined. As a result, either of the two extreme choices, $\underline{\pi}$ and $\overline{\pi}$, must be optimal in $\Pi^{\rm pb}$ when the state is one-dimensional.
In general, when the state is multi-dimensional, $\widehat{C}$ may be neither positive nor negative semidefinite, thus, partial disclosure can be optimal. Specifically, following the methodology developed by Tamura (2018), we can characterize an optimal public information structure and quantify the informativeness of an optimal public signal by utilizing the eigendecomposition of $\widehat{C}$.
Allowing for geometric multiplicity, let us denote by $\gamma_{1},\ldots,\gamma_{m}$ the $m$ real eigenvalues of $\widehat{C}$ and by $\mathbf{u}_{1},\ldots,\mathbf{u}_{m}$ the corresponding eigenvectors, which constitute an orthonormal system in $\mathbb{R}^{m}$. We align the eigenvalues in a non-increasing order, and let $k^{*}$ be the last index such that the corresponding eigenvalue exceeds 00, i.e.,
$\gamma_{1}\geq\gamma_{2}\geq\cdots\geq\gamma_{k^{*}}>0\geq\gamma_{k^{*}+1}\geq\cdots\geq\gamma_{m},$
where the convention $k^{*}=0$ is adopted when $\widehat{C}$ is negative semidefinite. By the spectral theorem, $\widehat{C}$ can be factorized as follows:
$\widehat{C}=U\operatorname{Diag}(\gamma_{1},\ldots,\gamma_{m})U^{\top},$ (30)
where $U\equiv[\mathbf{u}_{1}\cdots\mathbf{u}_{m}]$ is an $m\times m$ orthogonal matrix, which has each $\mathbf{u}_{k}$ as the $k$-th column vector. In addition, let $U_{k^{*}}\equiv[\mathbf{u}_{1}\cdots\mathbf{u}_{k^{*}}]\in\mathcal{M}^{m,k^{*}}$.
Proposition 16.
Let $\pi$ be a public Gaussian information structure, which sends a public signal $t_{0}=U_{k^{*}}Z^{-1/2}\theta$. Then $\pi$ is optimal in $\Pi^{\rm pb}$, and the associated state variance reduction is given by
$\operatorname{Var}\quantity[\operatorname{E}\quantity[\theta\mid t_{0}]]=Z^{\frac{1}{2}}U_{k^{*}}U_{k^{*}}^{\top}Z^{\frac{1}{2}}.$ (31)
Moreover, $v^{\rm pb}$ is equal to the sum of all positive eigenvalues of $\widehat{C}$, i.e., $\gamma_{1}+\cdots+\gamma_{k^{*}}$. Finally, $k^{*}$ satisfies $p(\mathbf{V})-n\leq k^{*}\leq p(\mathbf{V})$, where $p(\mathbf{V})$ is the number of positive eigenvalues of $\mathbf{V}$.
This proposition characterizes an optimal public information structure by deriving an explicit form of the public signal. Specifically, the optimal public signal is determined by the spectral property of the matrix $\widehat{C}$, which is defined through the primitives. Notice that any optimal public information structure is noise-free, as the signal being a linear transformation of the state. In addition, it is shown that the value of public LQG information design equals the sum of the positive eigenvalues of $\widehat{C}$.
If $\widehat{C}$ is positive definite, i.e., if all the eigenvalues are positive, then full disclosure is optimal because (31) reduces to $Z$ when $k^{*}=m$. In contrast, if $\widehat{C}$ is negative semidefinite, i.e., if all the eigenvalues are nonpositive, then no disclosure is optimal. In general, $\widehat{C}$ is neither positive nor negative semidefinite, yet, the optimal public signal is identified as a multi-dimensional statistics, whose dimension coincides with the number of positive eigenvalues of $\widehat{C}$. Specifically, there is a monotonic relationship between the number of positive eigenvalues of $\widehat{C}$ and the informativeness of an optimal public signal, as the associated state variance reduction (31) is increasing in $k^{*}$ with respect to the matrix ordering. In other words, as $\widehat{C}$ gets closer to a positive definite matrix, the optimal signal discloses more information about $\theta$.
Implementing an optimal public information structure requires the designer’s detailed knowledgeable about the basic game in play since the key matrix $\widehat{C}$ hinges on the entire payoff structure. In practice, however, $Q$ and $R$ may be agents’ private information, as being parts of their payoff functions, and the designer may be agnostic on these parameters. Still, in Proposition 16, we derive some relationships between the number of positive eigenvalues of $\widehat{C}$ and that of $\mathbf{V}$. This enables the designer to, at least, partially identify the optimal degree of public information disclosure from her objective function alone. Since the tightness of the bounds on $k^{*}$ depends on $n$, such partial identification tends to be accurate when the number of agents is small relative to $m$.
Tamura (2018) studies an information design problem with a single player under a quadratic specification of the designer’s objective. His model assumes that the individual’s best response is given as the best estimate about a multi-dimensional state, $\operatorname{E}[\theta\mid t]$, upon receiving a signal $t$. He then shows that the designer’s objective maximization is reduced to the SDP problem,
$\max_{S\in\mathcal{S}^{m}}\quad W\bullet S\quad\text{subject to}\quad\operatorname{Var}\quantity[\theta]\succeq S\succeq O,$ (32)
where $W\in\mathcal{S}^{m}$ is a constant matrix, and obtains a closed form solution to the problem. He also demonstrates that this SDP formulation is useful in studying an LQG network game with $n$ players and obtains an optimal public information structure in a special case of an LQG network game when $m=n$. In this regard, Proposition 16 generalizes his result by allowing for an arbitrary $m$.
We provide a full proof of Proposition 16 in the appendix of this paper for the sake of completeness, while the solution method of Tamura (2018) can be utilized to establish the optimality of the public information structure. Our proof is more direct and simpler than that of Tamura (2018) because we use the properties of normal distributions, which he does not assume in solving (32).
To conclude, we address a natural question of what conditions are needed for the optimal public information structure to be “globally optimal” among all information structures, including non-public ones.
Corollary 17.
The optimal public information structure in Proposition 16 attains $v^{\rm p}$ if and only if there exists $\Lambda\in\mathcal{S}^{n}_{\rm diag}$ with $A_{\Lambda}\succeq O$ such that
$\quantity(A_{\Lambda}Q^{-1}R-B_{\Lambda})Z^{\frac{1}{2}}\mathbf{u}_{k}=\mathbf{0},\quad\forall k=1,\ldots,k^{*};\quad{\rm and}$ (33)
$B_{\Lambda}Z^{\frac{1}{2}}\mathbf{u}_{k}=\mathbf{0},\quad\forall k=k^{*}+1,\ldots,m.$ (34)
This corollary stems from the observation that the optimal public information structure $\pi$ in Proposition 16 is noise-free by construction. Therefore, we can apply Corollary 8 to derive the condition for $\pi$ to be globally optimal in terms of its certifiability. Incidentally, the optimality condition involves both matrix equations in Corollary 7, while each equation only needs to be partially satisfied on the linear subspace spanned by $\{Z^{1/2}\mathbf{u}_{1},\ldots,Z^{1/2}\mathbf{u}_{k^{*}}\}$ and $\{Z^{1/2}\mathbf{u}_{k^{*}+1},\ldots,Z^{1/2}\mathbf{u}_{m}\}$, respectively. [^22] Note that when $k^{*}=0$ so that $\underline{\pi}$ is the optimal public information structure, (33) is vacuously satisfied, while (34) is equivalent to $B_{\Lambda}=O$ since $\{Z^{1/2}\mathbf{u}_{1},\ldots,Z^{1/2}\mathbf{u}_{m}\}$ constitutes a basis of $\mathbb{R}^{m}$. Similarly, when $k^{*}=m$, it can be seen that (33) and (34) hold if and only if $A_{\Lambda}Q^{-1}R=B_{\Lambda}$. Therefore, Corollary 17 generalizes Corollary 7. Nevertheless, since (33) and (34) collectively impose $nm$ number of linear equations on $n$ Lagrangian multipliers, the optimal public information structure is not globally optimal for generic payoff structures. This indicates that the designer typically benefits from manipulating each agent’s higher-order belief about others’ private information, aligning with the findings of Arieli and Babichenko (2019) and Hoshino (2022).
Appendix
Appendix A Mathematical preliminaries
Given $C,A_{1},\ldots,A_{m}\in\mathcal{S}^{n}$ and $b\in\mathbb{R}^{n}$, a canonical SDP problem takes the following form:
$V^{\rm p}=\max_{X\in\mathcal{S}^{n}}C\bullet X\quad\text{subject to}\quad A_{i}\bullet X=b_{i}\quad\forall i\in[m],\quad X\succeq O.$ (A.1)
The dual problem of this problem takes the following minimization form:
$V^{\rm d}=\min_{y\in\mathbb{R}^{n}}b^{\top}y\quad\text{subject to}\quad\sum_{i\in[m]}y_{i}A_{i}\succeq C.$ (A.2)
Each problem is said to be feasible if it has a choice variable satisfying all subjective constraints, and it is strictly feasible if there exists a feasible choice that strictly satisfies the matrix inequality constraint.
The following lemma summarizes some well-known results regarding weak and strong duality, holding between the primal and dual problems; for reference, see, e.g., Section 3 of Vandenberghe and Boyd (1996).
Lemma 5.
If both primal and dual problems are feasible, then both admit solutions and $-\infty<V^{\rm p}\leq V^{\rm d}<\infty$ holds. In particular, if a primal-dual feasible pair $(X,y)$ satisfies
$X\bullet\quantity(\sum_{i\in[m]}y_{i}A_{i}-C)=0,$ (A.3)
then $X$ attains $V^{\rm p}$ and $y$ attains $V^{\rm d}$. If, in addition, either problem is strictly feasible, then $V^{\rm p}=V^{\rm d}$ holds, and any primal-dual feasible pair $(X,y)$ satisfies (A.3) if and only if $X$ attains $V^{\rm p}$ and $y$ attains $V^{\rm d}$.
It should be clear from construction that the solution set to either problem is convex. The next lemma will be used to prove results in Section 5.
Lemma 6.
If $(X,y)$ and $(X^{\prime},y^{\prime})$ are primal-dual optimal pairs, then so is their convex combination $(\lambda X+(1-\lambda)X^{\prime},\lambda y+(1-\lambda)y^{\prime})$ for any $\lambda\in[0,1]$.
For the sake of later reference, we record some known facts about pseudoinverse matrices. The next result characterizes the positive semidefiniteness of block matrices based on generalized Schur complements; see, e.g., Proposition 10.2.5 of Bernstein (2018).
Lemma 7.
For any block matrix $\mathbf{A}=\quantity[\begin{smallmatrix}A&B\\ B^{\top}&C\end{smallmatrix}]\in\mathcal{S}^{n+m}$, the following are equivalent:
  1. i).
    $\mathbf{A}\succeq O$.
  2. ii).
    $A\succeq O$, $B=AA^{+}B$, and $C\succeq B^{\top}A^{+}B$.
  3. iii).
    $C\succeq O$, $B=BC^{+}C$, and $A\succeq BC^{+}B^{\top}$.
The next result provides a well-known conditional formula for multivariate normal distributions. The following version, which can be found in Rao (1973), does not assume the invertibility of a variance matrix.
Lemma 8.
Let $\xi=(\xi_{1},\xi_{2})$ be a multivariate normal random vector, partitioned into $n$-dimensional $\xi_{1}$ and $m$-dimensional $\xi_{2}$, with
$\operatorname{E}\matrixquantity[\xi_{1}\\ \xi_{2}]=\matrixquantity[\mu_{1}\\ \mu_{2}]\in\mathbb{R}^{n+m}\quad\text{and}\quad\operatorname{Var}\matrixquantity[\xi_{1}\\ \xi_{2}]=\matrixquantity[\Sigma_{11}&\Sigma_{12}\\ \Sigma_{21}&\Sigma_{22}]\in\mathcal{S}^{n+m}_{+}.$
Then,
$\operatorname{E}\quantity[\xi_{1}\mid\xi_{2}]=\mu_{1}+\Sigma_{12}\Sigma_{22}^{+}\quantity(\xi_{2}-\mu_{2})\quad{\rm and}\quad\operatorname{Var}\quantity[\xi_{1}\mid\xi_{2}]=\Sigma_{11}-\Sigma_{12}\Sigma_{22}^{+}\Sigma_{21}.$
Appendix B Omitted proofs
Proof of Lemma 2
Given any information structure $\pi$, Proposition 4 of Miyashita and Ui (2024) implies that $(\operatorname{Var}\quantity[\sigma^{\pi}],\operatorname{Cov}\quantity[\sigma^{\pi},\theta])$ is primal feasible.
Conversely, let $(X,Y)\in\mathcal{S}^{n}\times\mathcal{M}^{n,m}$ be primal feasible. Let $\pi$ be a one-dimensional Gaussian information structure such that $\operatorname{E}\quantity[t]=\bar{a}$, $\operatorname{Var}\quantity[t]=X$, and $\operatorname{Cov}\quantity[t,\theta]=Y$, which exists since $\mathbf{M}_{X,Y}\succeq O$. Then, (10) implies
$\sum_{j=1}^{n}q_{ij}\operatorname{Cov}[t_{i},t_{j}]=\sum_{k=1}^{m}r_{ik}\operatorname{Cov}[t_{i},\theta_{k}].$
Assuming $\operatorname{Var}\quantity[t_{i}]>0$ for a while, by multiplying both sides of the above by $(t_{i}-\bar{a}_{i})/\operatorname{Var}[t_{i}]$, we obtain
$\sum_{j=1}^{n}q_{ij}\quantity(\frac{\operatorname{Cov}[t_{i},t_{j}]}{\operatorname{Var}[t_{i}]}\cdot\quantity(t_{i}-\bar{a}_{i}))=\sum_{k=1}^{m}r_{ik}\quantity(\frac{\operatorname{Cov}[t_{i},\theta_{k}]}{\operatorname{Var}[t_{i}]}\cdot\quantity(t_{i}-\bar{a}_{i})),$
from which Lemma 8 yields
$\sum_{j=1}^{n}q_{ij}\quantity(\operatorname{E}[t_{j}\mid t_{i}]-\bar{a}_{j})=\sum_{k=1}^{m}r_{ik}\quantity(\operatorname{E}[\theta_{k}\mid t_{i}]-\bar{\theta}_{k}).$
By the definition of $\bar{a}$, it follows that
$\sum_{j=1}^{n}q_{ij}\operatorname{E}[t_{j}\mid t_{i}]=\sum_{k=1}^{m}r_{ik}\operatorname{E}[\theta_{k}\mid t_{i}].$ (B.1)
Note that (B.1) holds true even when $\operatorname{Var}\quantity[t_{i}]=0$ since $\operatorname{E}[t_{j}\mid t_{i}]=a_{j}$ and $\operatorname{E}[\theta_{k}\mid t_{i}]=\bar{\theta}_{k}$.
Now, let $\sigma$ be a strategy profile such that $\sigma_{i}(t_{i})=t_{i}$ for all $i\in[n]$ and $t_{i}\in\mathbb{R}$. Then, (B.1) means that $\sigma$ constitutes a BNE under $\pi$, i.e., $\sigma=\sigma^{\pi}$. Moreover, by the distribution of $t$, we have $\mathbf{M}^{\pi}=\mathbf{M}_{X,Y}$, as desired. Q.E.D.
Proof of Proposition 2
In order to apply Lemma 5, our first task is reformulating the problem (P) in the SDP canonical form as in (A.1). For $i,j\in[n+m]$, let $E_{ij}$ denote the $(n+m)\times(n+m)$ square matrix whose $(i,j)$-entry is $1$ and all others are 00. Then, for each $i\in[n]$, let $\Phi_{i}\in\mathcal{S}^{n+m}$ be given as
$\Phi_{i}=\quantity(\sum_{j\in[n]}q_{ij}\frac{E_{ij}+E_{ji}}{2})-\quantity(\sum_{k\in[m]}r_{ik}\frac{E_{i,n+k}+E_{n+k,i}}{2}).$ (B.2)
Also, for each $k,k^{\prime}\in[m]$, let $\Psi_{kk^{\prime}}\in\mathcal{S}^{n+m}$ be given as
$\Psi_{kk^{\prime}}=\frac{E_{n+k,n+k^{\prime}}+E_{n+k^{\prime},n+k}}{2}.$ (B.3)
Now, consider any symmetric block matrix $\mathbf{M}$ of the form
$\mathbf{M}=\matrixquantity[X&Y\\ Y^{\top}&Z]\quad\text{where}\quad X\in\mathcal{S}^{n},\quad Y\in\mathcal{M}^{n,m},\quad Z\in\mathcal{S}^{m}.$
Since the upper-left block $X$ is symmetric, $\Phi_{i}\bullet\mathbf{M}=0$ holds if and only if $\sum_{j\in[n]}q_{ij}x_{ij}=\sum_{k\in[m]}r_{ik}y_{ik}$. Additionally, since $Z$ is symmetric, $\Psi_{kk^{\prime}}\bullet\mathbf{M}=\operatorname{Cov}[\theta_{k},\theta_{k^{\prime}}]$ is satisfied for all $k,k^{\prime}\in[m]$ if and only if $Z=\operatorname{Var}[\theta]$. Therefore, we can rewrite the problem (P) as
$v^{\rm p}=$ $\ \max_{\mathbf{M}\in\mathcal{S}^{n+m}}\quad$ $\mathbf{V}\bullet\mathbf{M}$
subject to $\Phi_{i}\bullet\mathbf{M}=0,\quad\forall i\in[n],$
$\Psi_{kk^{\prime}}\bullet\mathbf{M}=\operatorname{Cov}[\theta_{k},\theta_{k^{\prime}}],\quad\forall k,k^{\prime}\in[m],$
$\mathbf{M}\succeq O,$
which takes the canonical form. Notice that this problem is feasible since all constraints are satisfies by letting $X=O$, $Y=O$, and $Z=\operatorname{Var}[\theta]$.
To derive the dual problem, we denote by $\lambda_{i}$ the Lagrangian multiplier for each equality constraint $\Phi_{i}\bullet\mathbf{M}=0$ and by $\gamma_{kk^{\prime}}$ the one for $\Psi_{kk^{\prime}}\bullet\mathbf{M}=0$. Then, let $\Lambda=\operatorname{Diag}(\lambda_{1},\ldots,\lambda_{n})$ and $\Gamma=[\gamma_{kk^{\prime}}]_{m\times m}$. With these notations, following (A.2), the dual problem can be formulated as follows:
$v^{\rm d}=$ $\min_{\Lambda\in\mathcal{S}^{n}_{\rm diag},\,\Gamma\in\mathcal{M}^{m\times m}}\quad$ $\operatorname{Var}[\theta]\bullet\Gamma$
subject to $\sum_{i\in[n]}\lambda_{i}\Phi_{i}+\sum_{k,k^{\prime}\in[m]}\gamma_{kk^{\prime}}\Psi_{kk^{\prime}}\succeq\mathbf{V}.$
By the definitions of $\Phi_{i}$ and $\Psi_{kk^{\prime}}$, it is not hard to see that
$\sum_{i\in[n]}\lambda_{i}\Phi_{i}+\sum_{k,k^{\prime}\in[m]}\gamma_{kk^{\prime}}\Psi_{kk^{\prime}}=\matrixquantity[\quantity(\Lambda Q+Q^{\top}\Lambda)/2&-\Lambda R/2\\ -R^{\top}\Lambda/2&\quantity(\Gamma+\Gamma^{\top})/2].$
Since $v^{\rm d}$ depends on $\Gamma$ only through their symmetric parts, it is without loss of optimality to let $\Gamma\in\mathcal{S}^{n}$. By doing so, the above-constructed dual problem coincides with (4.1), and note that the problem is feasible since $\Lambda=aI$ and $\Gamma=C_{\Lambda}$ satisfies dual feasibility if $a>0$ is large enough.
Let us show that the dual problem is strictly feasible. To this end, let $a,b>0$ be positive numbers, $\Lambda=aI$, and $\Gamma=abI$. Then, it follows that
$\matrixquantity[\quantity(\Lambda Q+Q^{\top}\Lambda)/2&-\Lambda R/2\\ -R^{\top}\Lambda/2&\Gamma]=a\underbrace{\matrixquantity[(Q+Q^{\top})/2&-R/2\\ -R^{\top}/2&bI]}_{(*)},$
which is positive definite if and only if the matrix ()(*)( ∗ ) is positive definite. Moreover, ()(*)( ∗ ) is positive definite if and only if
$\frac{Q+Q^{\top}}{2}-\frac{RR^{\top}}{4b}\succ O.$
By Assumption 1, this matrix inequality holds by taking a sufficiently large $b$. Moreover, having fixed $b$ in this way, we obtain a strictly feasible dual choice by letting $a$ be sufficiently large. Thus, by Lemma 5, it follows that $v^{\rm p}=v^{\rm d}$, and both problems are attainable. Q.E.D.
Proof of Proposition 3
Let $\Lambda\in\mathcal{S}^{n}_{\rm diag}$ be any diagonal matrix such that $\mathbf{M}_{\Lambda}\succeq O$ and $Z\bullet C_{\Lambda}=v^{\rm d}$. Then, by Lemma 5, any primal feasible $(X,Y)\in\mathcal{S}^{n}\times\mathcal{M}^{n,m}$ attains $v^{\rm p}$ if and only if $\mathbf{M}_{X,Y}\bullet\mathbf{M}_{\Lambda}=0$. Moreover, since $\mathbf{M}_{X,Y}$ and $\mathbf{M}_{\Lambda}$ are positive semidefinite, $\mathbf{M}_{X,Y}\bullet\mathbf{M}_{\Lambda}=0$ is equivalent to $\mathbf{M}_{X,Y}\mathbf{M}_{\Lambda}=O$, which can be expanded to the following system of matrix equations:
$XA_{\Lambda}=YB_{\Lambda}^{\top},\quad XB_{\Lambda}=YC_{\Lambda},\quad Y^{\top}A_{\Lambda}=ZB_{\Lambda}^{\top},\quad Y^{\top}B_{\Lambda}=ZC_{\Lambda}.$ (B.4)
In the system, the first and third equations are nothing more than (16). The next lemma shows that the rest are redundant in the presence of these two.
Lemma 9.
For any primal feasible $(X,Y)$ and dual feasible $\Lambda$, if they satisfy the first and third equations in (B.4), then they satisfy the second and forth as well.
Proof.
Since $\mathbf{M}_{\Lambda}\succeq O$, Lemma 7 implies $B_{\Lambda}=A_{\Lambda}A_{\Lambda}^{+}B_{\Lambda}$. By the definition of $C_{\Lambda}$, the first equation in (B.4) implies that
$XB_{\Lambda}-YC_{\Lambda}=\underbrace{\quantity(XA_{\Lambda}-YB_{\Lambda}^{\top})}_{=\ O}A_{\Lambda}^{+}B_{\Lambda}=O,$
which implies the second equation in (B.4). Likewise, the third equation in (B.4) yields
$Y^{\top}B_{\Lambda}-ZC_{\Lambda}=\underbrace{\quantity(Y^{\top}A_{\Lambda}-ZB_{\Lambda})}_{=\ O}A_{\Lambda}^{+}B_{\Lambda}=O,$
which implies the fourth equation in (B.4). ∎
By Lemma 9, we conclude that any primal feasible $(X,Y)$ attains $v^{\rm p}$ if and only if there is dual feasible $\Lambda$ for which (16) is satisfied. It remains to prove the equivalence among (16), (17), and (18). Let $(X,Y)$ be primal feasible. Since $\mathbf{M}_{X,Y}\succeq O$, Lemma 7 implies that
$Y=XX^{+}Y=YZ^{+}Z.$ (B.5)
Now, whenever $A_{\Lambda}X=B_{\Lambda}Y^{\top}$ is satisfied, (B.5) implies that
$A_{\Lambda}Y=A_{\Lambda}XX^{+}Y=B_{\Lambda}Y^{\top}X^{+}Y,$
from which we see that $A_{\Lambda}Y=B_{\Lambda}Z$ if and only if $B_{\Lambda}Z=B_{\Lambda}Y^{\top}X^{+}Y$. This establishes the equivalence between (16) and (17). Likewise, whenever $A_{\Lambda}Y=B_{\Lambda}Z$ is satisfied, (B.5) implies that
$B_{\Lambda}Y=B_{\Lambda}ZZ^{+}Y^{\top}=A_{\Lambda}YZ^{+}Y^{\top},$
from which we see that $B_{\Lambda}Y=A_{\Lambda}X$ if and only if $A_{\Lambda}X=A_{\Lambda}YZ^{+}Y^{\top}$. This establishes the equivalence between (16) and (18). Q.E.D.
Proof of Corollary 4
The first assertion is discussed in the main text. Moreover, the equivalence holds because $\mathbf{V}\bullet\mathbf{M}_{X,Y}=Z\bullet C_{\Lambda}$ if and only if $\mathbf{M}_{X,Y}\bullet\mathbf{M}_{\Lambda}=0$ (cf. Lemma 5), and because $\mathbf{M}_{X,Y}\bullet\mathbf{M}_{\Lambda}=0$ if and only if (16) (cf. Lemma 9 and the discussion preceding to it). Q.E.D.
Proof of Corollary 5
Let $\pi\in\Pi^{\rm g}$ be any optimal information structure with $X=\operatorname{Var}\quantity[\sigma^{\pi}]$ and $Y=\operatorname{Cov}\quantity[\sigma^{\pi},\theta]$ that is certified by $\Lambda\in\mathcal{S}^{n}_{\rm diag}$. Note that $(X,Y)$ and $\Lambda$ jointly satisfy (B.4). Then, Proposition 2 implies $v^{\rm p}=C_{\Lambda}\bullet Z$, while (B.4) implies $C_{\Lambda}\bullet Z=B_{\Lambda}\bullet Y=A_{\Lambda}\bullet X$. Moreover, by using Lemma 8, it follows that
[…]
where the last equality holds by $\mathbf{M}_{X,Y}\succeq O$. Similarly, by the definition of $C_{\Lambda}$, we have
[…]
where the last equality holds by $\mathbf{M}_{X,Y}\succeq O$ and $\mathbf{M}_{\Lambda}\succeq O$. Q.E.D.
Proof of Lemma 3
Assuming $Z\succ O$, (16) implies that $B_{\Lambda}=A_{\Lambda}YZ^{-1}$ and $C_{\Lambda}=Z^{-1}Y^{\top}A_{\Lambda}YZ^{-1}$. Hence, it follows that
$\mathbf{M}_{\Lambda}=\matrixquantity[A_{\Lambda}&-A_{\Lambda}YZ^{-1}\\ -Z^{-1}Y^{\top}A_{\Lambda}&Z^{-1}Y^{\top}A_{\Lambda}YZ^{-1}]=\matrixquantity[I&O\\ O&YZ^{-1}]^{\top}\matrixquantity[A_{\Lambda}&-A_{\Lambda}\\ -A_{\Lambda}&A_{\Lambda}]\matrixquantity[I&O\\ O&YZ^{-1}],$
from which we see that $\mathbf{M}_{\Lambda}\succeq O$ if and only if $A_{\Lambda}\succeq O$. Q.E.D.
Proof of Corollary 6
For the first part of the corollary, we prove the following stronger statement, which does not require $Z\succ O$.
Lemma 10.
If $\pi\in\Pi^{\rm g}$ is optimal and certified by $\Lambda\in\mathcal{S}^{n}_{\rm diag}$, then the following are satisfied:
$B_{\Lambda}\quantity(\theta-\bar{\theta})=A_{\Lambda}XX^{+}\quantity(\sigma^{\pi}-\bar{a})\quad{\rm a.s.}$ (B.6)
$A_{\Lambda}\quantity(\sigma^{\pi}-\bar{a})=B_{\Lambda}ZZ^{+}\quantity(\theta-\bar{\theta})\quad{\rm a.s.}$ (B.7)
Proof.
Let $\pi\in\Pi^{\rm g}$ be any optimal information structure with $X=\operatorname{Var}\quantity[\sigma^{\pi}]$ and $Y=\operatorname{Cov}\quantity[\sigma^{\pi},\theta]$ that is certified by $\Lambda\in\mathcal{S}^{n}_{\rm diag}$. Coupled with Lemma 8, the condition (17) implies that
$\operatorname{E}\quantity[B_{\Lambda}\quantity(\theta-\bar{\theta})\mid\sigma^{\pi}]=B_{\Lambda}Y^{\top}X^{+}\quantity(\sigma^{\pi}-\bar{a})=A_{\Lambda}XX^{+}\quantity(\sigma^{\pi}-\bar{a}),$
$\operatorname{Var}\quantity[B_{\Lambda}\quantity(\theta-\bar{\theta})\mid\sigma^{\pi}]=B_{\Lambda}\quantity(Z-YX^{+}Y^{\top})B_{\Lambda}=O.$
This implies (B.6). Also, by (18), it follows that
$\operatorname{E}\quantity[A_{\Lambda}\quantity(\sigma^{\pi}-\bar{a})\mid\theta]=A_{\Lambda}YZ^{+}\quantity(\theta-\bar{\theta})=B_{\Lambda}ZZ^{+}\quantity(\theta-\bar{\theta}),$
$\operatorname{Var}\quantity[A_{\Lambda}\quantity(\sigma^{\pi}-\bar{a})\mid\theta]=A_{\Lambda}\quantity(X-Y^{\top}Z^{+}Y)A_{\Lambda}=O.$
This implies (B.7). ∎
In particular, if $Z\succ O$, then (B.7) implies that $A_{\Lambda}\quantity(\sigma^{\pi}-\bar{a})=B_{\Lambda}\quantity(\theta-\bar{\theta})$ a.s.
Conversely, suppose that $\pi$ and dual feasible $\Lambda$ jointly satisfy (19). Multiplying both sides of (19) by $\quantity(\sigma^{\pi}-\bar{a})^{\top}$ from right and taking unconditional expectations, it follows that
$A_{\Lambda}\quantity(\sigma^{\pi}-\bar{a})\quantity(\sigma^{\pi}-\bar{a})^{\top}=B_{\Lambda}\quantity(\theta-\bar{\theta})\quantity(\sigma^{\pi}-\bar{a})^{\top}\quad\Longrightarrow\quad A_{\Lambda}\operatorname{Var}\quantity[\sigma^{\pi}]=B_{\Lambda}\operatorname{Cov}\quantity[\theta,\sigma^{\pi}].$
Analogously, multiplying both sides of (19) by $\quantity(\theta-\bar{\theta})^{\top}$ from right and taking unconditional expectations, we get
$A_{\Lambda}\quantity(\sigma^{\pi}-\bar{a})\quantity(\theta-\bar{\theta})^{\top}=B_{\Lambda}\quantity(\theta-\bar{\theta})\quantity(\theta-\bar{\theta})^{\top}\quad\Longrightarrow\quad A_{\Lambda}\operatorname{Cov}\quantity[\sigma^{\pi},\theta]=B_{\Lambda}Z.$
Hence, $\pi$ and $\Lambda$ satisfy (16), from which Proposition 3 concludes that $\pi$ is optimal. Q.E.D.
Proof of Corollary 7
As we discuss in the main text, $\underline{\pi}$ is optimal if and only if it is certified by $\Lambda$ such that $\mathbf{M}_{\Lambda}\succeq O$ and $B_{\Lambda}=O$, while Lemma 3 implies $\mathbf{M}_{\Lambda}\succeq O$ is equivalent to $A_{\Lambda}\succeq O$ when $Z\succ O$. For the optimality of $\overline{\pi}$, taking $X=Q^{-1}RZ(Q^{-1}R)^{\top}$ and $Y=Q^{-1}RZ$, the first equality in (18) is automatically satisfied irrespective of $\Lambda$, and the second equality reduces to $A_{\Lambda}Q^{-1}RZ=B_{\Lambda}Z$. When $Z\succ O$, this holds true if and only if $A_{\Lambda}Q^{-1}R=B_{\Lambda}$. Q.E.D.
Proof of Corollary 8
By (B.7) in the claim obtained in the proof of Corollary 6, if there exists a dual optimal $\Lambda$ such that $A_{\Lambda}\succ O$, then any optimal $\pi\in\Pi^{\rm g}$ is noise-free. Consider any primal solutions $(X,Y)$ and $(\tilde{X},\tilde{Y})$. Then, by Proposition 3, the following equalities are satisfied:
$A_{\Lambda}X=B_{\Lambda}Y^{\top},\quad A_{\Lambda}Y=B_{\Lambda}Z,\quad A_{\Lambda}\tilde{X}=B_{\Lambda}\tilde{Y}^{\top},\quad A_{\Lambda}\tilde{Y}=B_{\Lambda}Z.$
By the second and forth equations, we have $A_{\Lambda}(Y-\tilde{Y})=O$, from which $Y=\tilde{Y}$ since $A_{\Lambda}$ is non-singular. Then, by the first and third equations, we have $A_{\Lambda}(X-\tilde{X})=O$, from which $X=\tilde{X}$ again because $A_{\Lambda}$ is non-singular. Q.E.D.
Proof of Corollary 9
[…]
Proof of Corollary 10
Suppose $m=1$. As for the optimality of no disclosure $\underline{\pi}$, notice that $B_{\Lambda}=O$ simplifies to $\lambda_{i}r_{i}+w_{i}=0$ for each $i\in[n]$, thereby dictating $\Lambda=-D$, where $D=\operatorname{Diag}\quantity(w_{1}/r_{1},\ldots,w_{n}/r_{n})$. Hence, by Corollary 7, it follows that $\underline{\pi}$ is optimal if and only if $A_{-D}\succeq O$, or equivalently, $O\succeq V+(DQ+Q^{\top}D)/2$.
Now, suppose that $O\not\succeq V+(DQ+Q^{\top}D)/2$. By the previous conclusion, it follows that $\underline{\pi}$ is suboptimal, meaning $v^{\rm p}>0$. Then, by Corollary 5, we need to have $B_{\Lambda}\neq O$ for any $\Lambda\in\mathcal{S}^{n}_{\rm diag}$ that certifies an optimal information structure. By $m=1$, this implies $B_{\Lambda}$ has full column rank, from which Corollary 9 implies that any optimal information structure is aggregatively disclosing. Q.E.D.
Proof of Proposition 11
Let $(X,Y)$ be any solution to (P), which is not necessarily $G$-invariant.
Lemma 11.
For any $\tau\in\operatorname{Aut}\quantity(G)$, $(\tau(X),\tau(Y))$ is primal feasible and solves (P).
Proof.
Take any $\tau\in\operatorname{Aut}\quantity(G)$ and write $P=P_{\tau}$ for simplicity. Note that $P$ is an orthogonal matrix, i.e., $P^{\top}=P^{-1}$. Since $Q=Q_{G}$ is $G$-invariant, we have
$\tau(X)Q^{\top}=PXP^{\top}\quantity(PQP^{\top})^{\top}=PXQ^{\top}P^{\top}.$ (B.8)
Since $(X,Y)$ is primal feasible, and since $R=I$, we have $\operatorname{diag}\quantity(XQ^{\top})=\operatorname{diag}\quantity(Y)$. By the construction of $P$, this implies
$\operatorname{diag}\quantity(PXQ^{\top}P^{\top})=\operatorname{diag}\quantity(PYP^{\top})=\operatorname{diag}\quantity(\tau(Y)).$ (B.9)
By (B.8) and (B.9), it follows that $\operatorname{diag}(\tau(X)Q^{\top})=\operatorname{diag}\quantity(\tau(Y))$. In addition, since $Z$ is $G$-invariant,
$\mathbf{M}_{\tau(X),\tau(Y)}=\matrixquantity[\tau(X)&\tau(Y)\\ \tau(Y)^{\top}&Z]=\matrixquantity[PXP^{\top}&PYP^{\top}\\ PY^{\top}P&PZP^{\top}]=\matrixquantity[P&O\\ O&P]\mathbf{M}_{X,Y}\matrixquantity[P&O\\ O&P]^{\top}\succeq O.$
Therefore, we have shown that $(\tau(X),\tau(Y))$ is primal feasible. Furthermore, since $V$ and $W$ are $G$-invariant, by using the orthogonality of $P$, it follows that
[…]
from which we confirm that $(\tau(X),\tau(Y))$ provides the same value as $(X,Y)$ to the designer. Since $(X,Y)$ attains $v^{\rm p}$, so does $(\tau(X),\tau(Y))$. ∎
Now, consider the average of $\tau(X)$ and $\tau(Y)$ over all $\tau\in\operatorname{Aut}(G)$,
$\bar{X}=\frac{\sum_{\tau\in\operatorname{Aut}\quantity(G)}\tau(X)}{\quantity|\operatorname{Aut}\quantity(G)|}\quad{\rm and}\quad\bar{Y}=\frac{\sum_{\tau\in\operatorname{Aut}\quantity(G)}\tau(Y)}{\quantity|\operatorname{Aut}\quantity(G)|}.$
By construction, $\bar{X}$ and $\bar{Y}$ are $G$-invariant matrices. Moreover, since $(\bar{X},\bar{Y})$ is a convex combination of $(\tau(X),\tau(Y))$, each of which solves (P) by Lemma 11, Lemma 6 implies that $(\bar{X},\bar{Y})$ solves (P) as well.
Let $(\Lambda,\Gamma)$ be any solution to (4.1). By similar arguments, one can show that $(\tau\quantity(\Lambda),\tau\quantity(\Gamma))$ is a solution to (4.1), and thus $(\bar{\Lambda},\bar{\Gamma})$, given by
$\bar{\Lambda}=\frac{\sum_{\tau\in\operatorname{Aut}\quantity(G)}\tau(\Lambda)}{\quantity|\operatorname{Aut}\quantity(G)|}\quad{\rm and}\quad\bar{\Gamma}=\frac{\sum_{\tau\in\operatorname{Aut}\quantity(G)}\tau(\Gamma)}{\quantity|\operatorname{Aut}\quantity(G)|},$
is also a solution. Hence, by Corollary 4, it follows that the $G$-invariant diagonal matrix $\bar{\Lambda}$ certifies the $G$-invariant optimal information structure. Q.E.D.
Proof of Corollary 12
Consider any payoff structure $(Q_{G},I,V,W)$ that maintains Assumption 3, where $\operatorname{Aut}\quantity(G)$ is transitive. To prove the contrapositive statement, suppose that there exists an optimal information structure $\pi$ that is not noise-free. By Proposition 11 (and Corollary 4), we know that $\pi$ is certified by some scalar matrix $\Lambda=\lambda I$. Then, $A_{\Lambda}$ and $B_{\Lambda}$ are calculated for $(Q_{G},I,V,W)$ as follows:
$A_{\Lambda}=\frac{\lambda}{2}\cdot\quantity(Q_{G}+Q_{G}^{\top})-V\quad\text{and}\quad B_{\Lambda}=\frac{\lambda I+W}{2},$
where $A_{\Lambda}\succeq O$. Since $\pi$ is not noise-free, Corollary 8 implies $A_{\Lambda}$ is singular, which in turn implies $A_{\Lambda}\operatorname{Cov}\quantity[\pi,\sigma^{\pi}]$ is singular. Thus, by (16), $B_{\Lambda}Z$ is singular. Moreover, since $Z$ is non-singular by assumption, it follows that $B_{\Lambda}$ is singular. Hence, by the construction of $A_{\Lambda}$ and $B_{\Lambda}$, we conclude that both $\lambda I+W$ and $\lambda/2\cdot(Q_{G}+Q_{G}^{\top})-V$ are singular. In addition, $\lambda/2\cdot(Q_{G}+Q_{G}^{\top})\succeq V$ since $A_{\Lambda}\succeq O$.
Therefore, we have shown that whenever the assumption of Corollary (12) is satisfied, any optimal information structure $\pi$ is noise-free, and $A_{\Lambda}\succ O$ for any $\Lambda$ certifying $\pi$. Moreover, in this case, Corollary 8 implies the solution to (P) is unique. Q.E.D.
Proof of Corollary 13
Consider any payoff structure $(Q_{G},I,V,wI)$ that maintains Assumption 3, where $w\in\mathbb{R}$ and $\operatorname{Aut}\quantity(G)$ is transitive.
We first show that for any primal feasible $(X,Y)$,
$V\bullet X+wI\bullet Y=\tilde{V}\bullet X,\quad\text{where}\quad\tilde{V}=V+\frac{w}{2}\cdot\quantity(Q_{G}+Q_{G}^{\top}).$
[…]
As a result, any $(X,Y)$ is optimal for the designer’s objective $(V,wI)$ if and only if it is optimal for $(\tilde{V},O)$.
By Proposition 11 (and Corollary 4), any optimal information structure $\pi$ is certified by some scalar matrix $\Lambda=\lambda I$. Then, $A_{\Lambda}$ and $B_{\Lambda}$ are calculated for the payoff structure $(Q_{G},I,\tilde{V},O)$ as follows:
$A_{\Lambda}=\frac{\lambda-w}{2}\cdot\quantity(Q_{G}+Q_{G}^{\top})-V\quad\text{and}\quad B_{\Lambda}=\frac{\lambda I}{2}.$
By Lemma 7, note that $\Lambda$ is dual feasible if and only if $A_{\Lambda}\succeq O$ and $B_{\Lambda}=A_{\Lambda}A_{\Lambda}^{+}B_{\Lambda}$. Then, Proposition 3 implies that $\underline{\pi}$ is optimal if and only if there exists $\lambda\in\mathbb{R}$ such that $A_{\Lambda}\succeq O$ and $B_{\Lambda}Z=O$, while the latter condition holds if and only if $\lambda=0$. Moreover, setting $\lambda=0$, $B_{\Lambda}=A_{\Lambda}A_{\Lambda}^{+}B_{\Lambda}$ is satisfied. Then, we have $A_{\Lambda}\succeq O$ if and only if $O\succeq\tilde{V}$, and hence $\underline{\pi}$ is optimal if and only if $O\succeq\tilde{V}$.
Now, suppose that $O\not\succeq\tilde{V}$. By the previous conclusion, $\underline{\pi}$ is suboptimal so that $v^{\rm p}>0$. Then, Corollary 5 implies $B_{\Lambda}\neq O$, from which we have $\lambda\neq 0$. This means that $B_{\Lambda}$ has full rank, thus Corollary 9 implies that $\pi$ is aggregatively disclosing.
Lastly, suppose that $Z\succ O$, and let $\lambda\in\mathbb{R}$ be such that $\lambda I+W$ is singular. Then, since $W=wI$, we have $\lambda=-w$. As a result, the matrix $\lambda/2\cdot(Q_{G}+Q_{G}^{\top})-V$ is singular if and only if so is $\tilde{V}$. Applying Corollary 12, therefore, an optimal information structure is noise-free, and the solution to (P) is unique whenever $\tilde{V}$ is non-singular. Q.E.D.
Proof of Corollary 14
Given a common Lagrangian multiplier $\lambda\in\mathbb{R}$, let us calculate
$A_{\Lambda}=(\lambda-w)Q-V\quad{\rm and}\quad B_{\Lambda}=\frac{\lambda I}{2}.$
Note that each on-diagonal entry of $A_{\Lambda}$ is $\lambda-\alpha$, and each off-diagonal entry is $q\lambda-\beta$. Consequently, we have $A_{\Lambda}\succeq O$ if and only if
$-\frac{\lambda-\alpha}{n-1}\leq q\lambda-\beta\leq\lambda-\alpha.$ (B.10)
Then, (i) readily follows from Corollary 13 since $O\succeq\tilde{V}$ if and only if (21) holds.
To prove (ii), note that $\rho_{x}=1$ holds under full disclosure. So, the designer’s value of $\overline{\pi}$ is given as $\tilde{V}\bullet X=n\quantity(\alpha+(n-1)\beta)x$. On the other hand, by Corollary 5, the designer’s optimal value equals to $A_{\Lambda}\bullet X=n\quantity(\lambda-\alpha+(n-1)(q\lambda-\beta))x$, where $\Lambda=\lambda I$ is dual optimal. Hence, a necessary condition for the optimality of $\overline{\pi}$ is
$\alpha+(n-1)\beta=\lambda-\alpha+(n-1)(q\lambda-\beta),$
which pins down $\lambda$ as
$\lambda=\frac{2\quantity(\alpha+(n-1)\beta)}{1+(n-1)q}.$
By strong duality (Proposition 2), $\overline{\pi}$ is optimal if and only if the above $\lambda$ gives rise to a dual feasible matrix $\Lambda=\lambda I$, i.e., (B.10) holds. Using the expression for $\lambda$, since $q>-1/(n-1)$, (B.10) is equivalent to
$-\frac{\alpha+(n-1)(2\beta-q\alpha)}{n-1}\leq(n-1)q\beta+2q\alpha-\beta\leq\alpha+(n-1)(2\beta-q\alpha),$
which simplifies to (24). Moreover, $-1/n<b(q)<1$ can be verified from the assumption that $-1/(n-1)<q<1$.
Lastly, for (iii), suppose that neither $\underline{\pi}$ nor $\overline{\pi}$ is optimal. Let $(x,y,\rho_{x},\rho_{y})$ be a candidate profile that gives rise to a primal solution $(X,Y)$ of the form (22). By the first equation in (16), we have
$(\lambda-\alpha)x+(n-1)(\lambda q-\beta)\rho_{x}x=\lambda y/2,$ (B.11)
$(\lambda q-\beta)x+(\lambda-\alpha)\rho_{x}x+(n-2)(\lambda q-\beta)\rho_{x}x=\lambda y/2,$
from which
$\underbrace{\quantity((\lambda-\alpha)-(\lambda q-\beta))}_{(*)}(1-\rho_{x})x=0.$
Since $\underline{\pi}$ is suboptimal, we have $x>0$. Also, since $\overline{\pi}$ is suboptimal, yet an optimal information structure is aggregatively disclosing, we have $\rho_{x}<1$. This implies that ()(*)( ∗ ) must be equal to 00, which pins down $\lambda$ as
$\lambda=\frac{\alpha-\beta}{1-q}.$ (B.12)
By the previous conclusions, since $\underline{\pi}$ and $\overline{\pi}$ are suboptimal, we see that $\lambda>0$. In addition, it follows that
$\lambda-\alpha=\lambda q-\beta=\frac{q\alpha-\beta}{1-q}.$ (B.13)
Noticing that $q>b(q)$, the suboptimality of $\overline{\pi}$ implies that $q\alpha>\beta$. Hence, we see that $A_{\Lambda}$ is a constant matrix whose entries are all positive, and thus, $A_{\Lambda}\succeq O$. Now, since $\rho_{y}=1$, the second equation in (16) yields
$(\lambda-\alpha)y+(n-1)(\lambda q-\beta)y=\lambda z/2.$
By substituting (B.12) and (B.13) into the above equation, we obtain the value of $y$ as in the statement. Then, we can solve for $x$ and $\rho_{x}$ from (23) amd (B.11). Finally, the variance reduction is computed by using the formula $z-\operatorname{Var}[\hat{\theta}_{i}\mid\sigma^{\pi}_{i}]=y^{2}/x$. Q.E.D.
Proof of Corollary 15
By Corollary 13, an optimal information structure is either no disclosure or aggregative disclosure, depending on whether (21) holds. Moreover, under Assumption 4, we have $\det(\lambda I+W)=0$ if and only if $\lambda=-w$. By specifying $\lambda$ in this way, we have $\lambda Q-V=-(V+wQ)$, and this matrix is positive semidefinite if and only if (21) holds. Moreover, it holds that
$\det\quantity(V+wQ)=(\alpha-\beta)^{n-1}(\alpha+(n-1)\beta),$
which implies that $V+wQ$ is singular if and only if $\beta=\alpha$ or $\beta=-\alpha/(n+1)$. Hence, by Corollary 12, if $(\alpha,\beta)$ is not on the boundary of (21), then (4.1) is solved by a strictly feasible choice. Also, any optimal information structure is noise-free, and the solution to (P) is unique.
Our remaining task is to prove (ii). By substituting (25) into the condition obtained in Corollary 7, we see that $\overline{\pi}$ is optimal if and only if there exists $\lambda\in\mathbb{R}$ such that
$(\lambda-w)Q-V\succeq O\quad{\rm and}\quad\quantity((\lambda-w)Q-V)Q^{-1}=\lambda I/2.$ (B.14)
In (B.14), the second condition reduces to $\lambda/2\cdot Q=V+wQ$, which is equivalent to that $\lambda=2\alpha$ and $q\lambda=2\beta$. Such $\lambda$ exists if and only if $\beta=q\alpha$. Moreover, since $\lambda-\alpha=\alpha$ and $q\lambda-\beta=\beta$, the first condition in (B.14) simplifies to $-\alpha/(n-1)\leq q\alpha\leq\alpha$, which is equivalent to $\alpha\geq 0$ by the assumption on $q$. Q.E.D.
Proof of Lemma 4
As discussed in the main text, a unique BNE is characterized as $\sigma^{\pi}=Q^{-1}R\operatorname{E}\quantity[\theta\mid t_{0}]$ for any public information structure $\pi\in\Pi^{\rm pb}$, from which $\mathbf{M}^{\pi}=\mathbf{M}_{\operatorname{Var}\quantity[\operatorname{E}\quantity[\theta\mid t_{0}]]}$ is confirmed. Moreover, $Z\succeq\operatorname{Var}\quantity[\operatorname{E}\quantity[\theta\mid t_{0}]]\succeq O$ holds by the law of total variance.
Conversely, take any $S\in\mathcal{S}^{m}$ such that $Z\succeq S\succeq O$. Then, we let $\pi\in\Pi^{\rm pb}\cap\Pi^{\rm g}$ be a public Gaussian information structure, sending an $m$-dimensional public signal $t_{0}$, such that $t_{0}$ and $\theta$ are jointly normally distributed with $\operatorname{Var}\quantity[t_{0}]=\operatorname{Cov}\quantity[t_{0},\theta]=S$. Such a random vector exists because we have
$\operatorname{Var}\matrixquantity[S&S\\ S&Z]\succeq O,$
as confirmed by $Z-SS^{+}S=Z-S\succeq O$. Moreover, using Lemma 8, the unique BNE under $\pi$ is computed as
$\sigma^{\pi}\quantity(t_{0})=Q^{-1}R\operatorname{E}\quantity[\theta\mid t_{0}]=Q^{-1}RSS^{+}t_{0}-\underbrace{Q^{-1}R\quantity(\bar{\theta}+SS^{+}\operatorname{E}\quantity[t_{0}])}_{\rm constant}.$
Hence, it follows that
$\operatorname{Var}\quantity[\sigma^{\pi}(t_{0})]=Q^{-1}RSS^{+}\operatorname{Var}\quantity[t_{0}]S^{+}S\quantity(Q^{-1}R)^{\top}=Q^{-1}RS\quantity(Q^{-1}R)^{\top},$
$\operatorname{Cov}\quantity[\sigma^{\pi}(t_{0}),\theta]=Q^{-1}RSS^{+}\operatorname{Cov}\quantity[t_{0},\theta]=Q^{-1}RSS^{+}.$
Thus we have $\mathbf{M}^{\pi}=\mathbf{M}_{S}$.
Now, substituting $\mathbf{M}^{\pi}=\mathbf{M}_{S}$ into (13), it follows that
$\operatorname{E}\quantity[v(\sigma^{\pi},\theta)]$ $=\mathbf{V}\bullet\mathbf{M}_{S}$
$=V\bullet Q^{-1}RS(Q^{-1}R)^{\top}+\frac{1}{2}W\bullet Q^{-1}RS+\frac{1}{2}W^{\top}\bullet S(Q^{-1}R)^{\top}$
[…]
$=\overline{C}\bullet S,$
which concludes the proof of Lemma 4. Q.E.D.
Proof of Proposition 16
Consider the eigendecomposition $\widehat{C}=UDU^{\top}$, as introduced in the main text, where we denote $D=\operatorname{Diag}(\gamma_{1},\ldots,\gamma_{m})$. We first show that the sum of positive eigenvalues of $\widehat{C}$ serves as an upper bound on the value of any feasible choice.
Lemma 12.
For any $S\in\mathcal{S}^{m}$, if $I\succeq S\succeq O$, then all diagonal entries of $U^{\top}SU$ are confined in $[0,1]$. Consequently,
$\widehat{C}\bullet S\leq\sum_{k=1}^{k^{*}}\gamma_{k}.$
Proof.
Denote by $s_{1},\ldots,s_{m}$ the diagonal entries of $U^{\top}SU$. Since $S\succeq O$, we have $U^{\top}SU\succeq O$, thereby all of $s_{1},\ldots,s_{m}$ being nonnegative. Moreover, since $I\succeq S$, and since $U$ is orthogonal, it follows that $I-U^{\top}SU=U^{\top}(I-S)U\succeq O$. This implies that all diagonal entries of $I-U^{\top}SU$ are nonnegative, and thus all $s_{1},\ldots,s_{m}$ are weakly less that $1$. Hence, it follows that
[…]
where the inequality holds since $\gamma_{k}>0$ if and only if $k\leq k^{*}$ and $s_{k}\in[0,1]$ for all $k$. ∎
We show that the designer can achieve the upper bound derived in Lemma 12 by sending the public signal $t_{0}=U_{k^{*}}^{\top}Z^{-1/2}\theta$. Notice that $\operatorname{Var}\quantity[t_{0}]=U_{k^{*}}^{\top}U_{k^{*}}$ and $\operatorname{Cov}\quantity[t_{0},\theta]=U_{k^{*}}^{\top}Z^{1/2}$ hold. Then, by Lemma 8, we have
$\operatorname{Var}\quantity[\operatorname{E}\quantity[\theta\mid t_{0}]]=Z^{\frac{1}{2}}U_{k^{*}}\quantity(U_{k^{*}}^{\top}U_{k^{*}})^{-1}U_{k^{*}}^{\top}Z^{\frac{1}{2}}=Z^{\frac{1}{2}}U_{k^{*}}U_{k^{*}}^{\top}Z^{\frac{1}{2}},$
where the second equality holds since the column vectors of $U_{k^{*}}$ constitutes an orthonormal system, thereby $U_{k^{*}}^{\top}U_{k^{*}}$ being an identify matrix. Hence, by Lemma 4, the designer’s expected value can be calculated as follows:
[…]
Note that the last equality holds because $U^{\top}U_{k^{*}}=[\delta_{ij}]_{m\times k^{*}}$, where $\delta_{ij}$ is the Kronecker delta.
Let us show that $p(\mathbf{V})-n\leq p(\widehat{C})\leq p(\mathbf{V})$. Since $\widehat{C}$ and $\overline{C}$ are congruent, it follows that $p(\widehat{C})=p(\overline{C})$ by Sylvester’s law of inertia. Moreover, by arranging (14), we have
$\overline{C}=\mathbf{Q}^{\top}\mathbf{V}\mathbf{Q}\quad{\rm where}\quad\mathbf{Q}\equiv\matrixquantity[Q^{-1}R\\ I]\in\mathcal{M}^{n+m,m}.$
[…]
Now, suppose that the $k$-th largest eigenvalue of $\overline{C}$ is strictly positive. By the Courant–Fischer theorem, there exist $\epsilon>0$ and a $k$-dimensional linear subspace $\mathcal{U}\subseteq\mathbb{R}^{m}$ such that
$\min_{\mathbf{u}\in\mathcal{U}\setminus\{\mathbf{0}\}}\frac{\mathbf{u}^{\top}\overline{C}\mathbf{u}}{\mathbf{u}^{\top}\mathbf{u}}\geq\epsilon.$
Since $\ker\quantity(\mathbf{Q})=\{\mathbf{0}\}$, it follows that
$\frac{\quantity(\mathbf{Q}\mathbf{u})^{\top}\mathbf{V}\quantity(\mathbf{Q}\mathbf{u})}{\quantity(\mathbf{Q}\mathbf{u})^{\top}\quantity(\mathbf{Q}\mathbf{u})}\cdot\frac{\quantity(\mathbf{Q}\mathbf{u})^{\top}\quantity(\mathbf{Q}\mathbf{u})}{\mathbf{u}^{\top}\mathbf{u}}\geq\epsilon,\quad\forall\mathbf{u}\in\mathcal{U}\setminus\{\mathbf{0}\},$
from which we have
$\min_{\mathbf{u}\in\mathcal{U}\setminus\{\mathbf{0}\}}\frac{\quantity(\mathbf{Q}\mathbf{u})^{\top}\mathbf{V}\quantity(\mathbf{Q}\mathbf{u})}{\quantity(\mathbf{Q}\mathbf{u})^{\top}\quantity(\mathbf{Q}\mathbf{u})}\geq\frac{\epsilon}{\delta},\quad{\rm where}\quad\delta=\max\quantity{\frac{\quantity(\mathbf{Qu})^{\top}\mathbf{Qu}}{\mathbf{u}^{\top}\mathbf{u}}:\mathbf{u}\in\mathbb{R}^{m}\setminus\{\mathbf{0}\}}.$
Note that $\delta>0$ since $\mathbf{Q}^{\top}\mathbf{Q}$ is positive definite. Then, since $\{\mathbf{Qu}\in\mathbb{R}^{n+m}:\mathbf{u}\in\mathcal{U}\}$ is a $k$-dimensional linear subspace of $\mathbb{R}^{n+m}$, again by the Courant–Fischer theorem, it follows that the $k$-th largest eigenvalue of $\mathbf{V}$ is strictly positive. Thus $p(\overline{C})\leq p(\mathbf{V})$.
Let $p^{\prime}(\cdot)$ denote the number of non-negative eigenvalues of a given square matrix. Then, we can show that $p^{\prime}(\overline{C})\leq p^{\prime}(\mathbf{V})$ by slightly changing the preceding proof. Moreover, applying the same result to $-\overline{C}$ and $-\mathbf{V}$, it follows that $p^{\prime}(-\overline{C})\leq p^{\prime}(-\mathbf{V})$. Then, noticing that $p^{\prime}(-\overline{C})=m-p(\overline{C})$ and $p^{\prime}(-\mathbf{V})=(n+m)-p(\mathbf{V})$, we obtain
$m-p(\overline{C})\leq(n+m)-p(\mathbf{V}),$
from which $p(\mathbf{V})-n\leq p(\overline{C})$. Q.E.D.
Proof of Corollary 17
Let $\pi$ be the optimal public information structure as in Proposition 16. By construction, $\pi$ is noise-free, and
$\operatorname{Cov}\quantity[\sigma^{\pi},\theta]=Q^{-1}RZ^{\frac{1}{2}}U_{k^{*}}U_{k^{*}}^{\top}Z^{\frac{1}{2}}.$
Hence, by Corollary 8, $\pi$ is optimal in $\Pi^{*}$ if and only if there exists $\Lambda\in\mathcal{S}^{n}_{\rm diag}$ with $A_{\Lambda}\succeq O$ such that
$A_{\Lambda}Q^{-1}RZ^{\frac{1}{2}}U_{k^{*}}U_{k^{*}}^{\top}Z^{\frac{1}{2}}=B_{\Lambda}Z.$
By construction, $U_{k^{*}}U_{k^{*}}^{\top}=\sum_{k=1}^{k^{*}}\mathbf{u}_{k}\mathbf{u}_{k}^{\top}$ and $I=UU^{\top}=\sum_{k=1}^{m}\mathbf{u}_{k}\mathbf{u}_{k}^{\top}$. Hence, the above condition can be equivalently rewritten as follows:
$A_{\Lambda}Q^{-1}RZ^{\frac{1}{2}}U_{k^{*}}U_{k^{*}}^{\top}=B_{\Lambda}Z^{\frac{1}{2}}$
$\quad\iff\quad A_{\Lambda}Q^{-1}RZ^{\frac{1}{2}}\quantity(\sum_{k=1}^{k^{*}}\mathbf{u}_{k}\mathbf{u}_{k}^{\top})=B_{\Lambda}Z^{\frac{1}{2}}\quantity(\sum_{k=1}^{m}\mathbf{u}_{k}\mathbf{u}_{k}^{\top})$
$\quad\iff\quad\quantity(A_{\Lambda}Q^{-1}RZ^{\frac{1}{2}}-B_{\Lambda})Z^{\frac{1}{2}}\sum_{k=1}^{k^{*}}\mathbf{u}_{k}\mathbf{u}_{k}^{\top}-B_{\Lambda}Z^{\frac{1}{2}}\sum_{k=k^{*}+1}^{m}\mathbf{u}_{k}\mathbf{u}_{k}^{\top}=O.$ (B.15)
It is evident that (B.15) holds if both (33) and (34) are satisfied. Conversely, suppose that (B.15) is satisfied. Then, since $\{\mathbf{u}_{1},\ldots,\mathbf{u}_{m}\}$ constitutes an orthonormal basis in $\mathbb{R}^{m}$, multiplying (B.15) by $\mathbf{u}_{k}$ from right, we obtain (33) when $k\leq k^{*}$, and (34) when $k>k^{*}$. This implies that (B.15) is equivalent to (33) and (34), which concludes the proof. Q.E.D.